### v1

# Completely regular growth solutions to linear differential equations with exponential polynomials coefficients

Abstract
Consider the linear differential equation
$f^{(n)}+A_{n-1}f^{(n-1)}+\cdots+A_{0}f=0$ (+)( + )
where the coefficients $A_{j},j=0,\ldots,n-1,$ are exponential polynomials. It is known that every solution is entire. This paper will show that all transcendental solutions of finite growth order are of completely regular growth. This problem was raised in Heittokangas et al.[8, p. 33], which involves an extensive question about Gol’dberg-Ostrovskiǐ’s Problem [5, p. 300]. Moreover, we define functions in a generalized class concluding exponential polynomial functions, which are also of completely regular growth.
Keywords: completely regular growth, exponential polynomials, Stokes rays, matrix asymptotic solutions
2020 MSC: 30D15, 30D35
1 Introduction
An exponential polynomial is an entire function of the form
$G=P_{1}e^{Q_{1}}+\cdots+P_{l}e^{Q_{l}}$ (1.1)
where $P_{j},Q_{j},j=1,\ldots,l,$ are polynomials, $l\in\mathbb{N}$. The Phragmén-Lindelöf indicator function of an entire function $f(z)$ of finite order $\rho(f)>0$ is given by
$h_{f}(\theta)=\limsup_{r\to\infty}r^{-\rho(f)}\log|f(re^{i\theta})|$
for $\theta\in[0,2\pi)$. We say that $f$ is of completely regular growth (briefly denoted by c.r.g., see [9, pp. 139-140] or [11, pp. 6-8]) if there exists a sequence of Euclidean discs $D(z_{k},r_{k})$ satisfying
$\sum_{|z_{k}|\leq r}r_{k}=o(r)$ (1.2)
such that
$\log|f(re^{i\theta})|=(h_{f}(\theta)+o(1))r^{\rho(f)},\quad re^{i\theta}\not\in\bigcup_{k}D(z_{k},r_{k}),$ (1.3)
as $r\to\infty$ uniformly in $\theta$.
Detailed studies of functions of c.r.g. can be found in [9] and [11]. Transcendental exponential polynomial functions are of completely regular growth, see [6, Lemma 1.3]. Petrenko [10, pp. 104-112] and Steinmetz [12] independently proved that the solutions of linear differential equations with polynomial coefficients are of c.r.g.. Gol’dberg and Ostrovskiǐ stated a classic problem in the book [5, p. 300]:
(Gol’dberg-Ostrovskiǐ) Suppose that $f$ is a finite order entire transcendental solution of
$f^{(n)}+a_{n-1}f^{(n-1)}+\cdots+a_{1}f^{\prime}+a_{0}f=0,$
whose coefficients are entire functions of completely regular growth. Is $f$ also of completely regular growth?
Recently Bergweiler [1] has constructed a counter-example about G.-O.’s problem. Therefore, Heittokangas et al.[8, p. 33] propose a relevant question by strengthening the assumption.
If the coefficients $a_{0}(z),\ldots,a_{n-1}(z)$ of
$f^{(n)}+a_{n-1}f^{(n-1)}+\cdots+a_{1}f^{\prime}+a_{0}f=0,$ (1.4)
are exponential polynomials, and if it possesses a transcendental entire solution $f$ of finite order of growth, then is it true that $f$ is of c.r.g.?
Our main work is to positively answer the problem above. We give an estimation of $\log|f|$ on every small sector except for some areas around a finite number of rays from the origin, resulting $f$ of c.r.g., see Theorem 3.2 as well as Theorem 3.6.
Here is the structure of the paper. Section 2.1 is the brief introduction of c.r.g. on the rays with a useful lemma; In Section 2.2, we present Steinmetz’s result for $\log|f|$ regarding solutions $f$ of linear differential equations with rational function coefficients in the whole complex plane, and then we use techniques from Wasow’s book [14] to modify it to expressions for $\log|f|$ regarding solutions $f$ of linear differential equations with asymptotically expressed coefficients in some sector. Contents of Section 3 cover the main results and corresponding proofs of Theorems in Section 3 are stated in Section 4; In Section 5 we discuss some generalizations. There we define a new class of functions, generalizing the class of exponential polynomials in the class of functions of completely regular growth.
2 Basic concepts and lemmas
2.1 Compactification of the angular set for functions of completely regular growth on the rays
We have given the definition of functions of completely regular growth above. In fact, completely regular growth is not only a global property but a local one. A set which can be covered by a sequence of discs $D(z_{k},r_{k})$ satisfying (1.2) will be called a $C_{0}$-set. The relative measure of a set $E$ of positive numbers is defined to be the limit
$m^{*}(E)=\lim\limits_{r\rightarrow\infty}\frac{mes(E_{r})}{r},$
where $E_{r}$ denotes the intersection of the set $E$ with the interval $(0,r)$, and mes$(E_{r})$ is the measure of $E_{r}$. It is obvious that the set of those values of $r$ for which the circle $|z|=r$ intersects a $C_{0}$-set is of zero relative measure. Refering to the definitions in [9, p. 139], a function $f$ that is holomorphic in $\{re^{i\theta}:\theta_{1}<\theta<\theta_{2},\theta_{1},\theta_{2}\in\mathbb{R},r>0\}$ and of order $\rho$ will be called a function of completely regular growth on the ray $\arg z=\theta$ for some $\theta\in(\theta_{1},\theta_{2})$ if the limit
$h_{f}(\theta)=\displaystyle\lim_{r\to\infty\atop r\notin E_{\theta}}\frac{\log|f(re^{i\theta})|}{r^{\rho}}$
exists, where $E_{\theta}$ is a set of zero relative measure depending on $\theta$. Denote $R_{\theta}=\{re^{i\theta},r\notin E_{\theta},r>0\}$ for some angle $\theta$, a subset $\mathbb{M}\subset(\theta_{1},\theta_{2})$ and
$R_{\mathbb{M}}=\{re^{i\theta}:r>0,r\notin\displaystyle\bigcup_{\theta\in\mathbb{M}\subset(\theta_{1},\theta_{2})}E_{\theta}\}\subset\displaystyle\bigcup_{\theta\in\mathbb{M}\subset(\theta_{1},\theta_{2})}R_{\theta}.$
We will say a function of completely regular growth on the set of rays $R_{\mathbb{M}}$ if
$H_{f}(r,\theta)=\frac{\log|f(re^{i\theta})|}{r^{\rho}},r\notin\displaystyle\bigcup_{\theta\in\mathbb{M}}E_{\theta},$
converges uniformly to $h_{f}(\theta)$ for all $\theta,\theta\in\mathbb{M}.$ Additionally, Levin [9, p. 140] has proved that $f(z)$ is a function of completely regular growth on every ray which is a limiting ray of the set of rays along which $f(z)$ is of completely regular growth.
Lemma 2.1
If a holomorphic function of growth order $\rho$ is of completely regular growth on each ray $R_{\theta}$ for $\theta\in\mathbb{M}$, then it is of completely regular growth on the set of rays $R_{\overline{\mathbb{M}}}$.
It follows from this, in particular, that an entire function of completely regular growth on rays that form an everywhere dense set is a function of completely regular growth on the whole plane.
2.2 The asymptotic solutions of differential equations in a sector
Considering the general differential equation
$f^{(n)}(z)+a_{n-1}(z)f^{(n-1)}(z)+\cdots+a_{1}(z)f^{\prime}(z)+a_{0}(z)f(z)=0$ (2.1)
with entire coefficients $a_{0}(z),\ldots,a_{n-1}(z)$, Petrenko [10, pp. 104–112] has shown that transcendental solutions to linear differential equations (2.1) with polynomial coefficients are of completely regular growth. Steinmetz[12, Theorem 1] has proved a more generalized theorem in the following. The differential equation
$f^{(n)}(z)+R_{n-1}(z)f^{(n-1)}(z)+\cdots+R_{1}(z)f^{\prime}(z)+R_{0}(z)f(z)=0$ (2.2)
allows rational coefficients $R_{0}(z),\ldots,R_{n-1}(z)$. And it is always assumed the solutions of (2.2) are meromorphic functions.
Lemma 2.2 (Steinmetz)
Let $f(z)$ be a non-rational meromorphic solution of (2.2) and let $\arg z=\theta$ be an arbitrary direction. Then there is $p\in\mathbb{N}$, $h>0$, and a polynomial $Q$ in $z^{1/p}$ such that
$\log|f(z)|=\Re Q(z^{1/p})+O(\log|z|)$ (2.3)
as $z\rightarrow\infty$ in $\theta\leq\arg z\leq\theta+h,$ possibly outside an exceptional set consisting of
  1. 1.
    (countably many) disks $|z-z_{\mu}|<|z_{\mu}|^{1-\varepsilon},$ where $\varepsilon$ is positive and the counting function of the sequence {$z_{\mu}$} is $O(\log r)$;
  2. 2.
    a logarithmic semi-strip $0\leq\arg z-\theta<\dfrac{\log^{+}|z|}{|z|^{\frac{1}{p}}}$.
The latter occurs only for some finite numbers of rays.
This kind of exceptional rays are called Stokes rays of $f(z).$ Concerning the definition of Stokes ray, one can refer to Steinmetz[12], and the general definition in Wasow[14, Sec. 15].
In proving our results, our goal is not to assume anymore that the coefficients of the equation (2.2) are rational functions, but to assume that the coefficients have an asymptotic expansion at $\infty$ in a sector $S$. From this point of view, it is necessary to define the asymptotic symbol $\sim$. We write
$f(x)\sim C_{0}+\frac{C_{1}}{x}+\frac{C_{2}}{x^{2}}+\cdots,$
when
$\lim_{x\to\infty}\left(f(x)-C_{0}-\frac{C_{1}}{x}-\frac{C_{2}}{x^{2}}-\cdots-\frac{C_{m}}{x^{m}}\right)x^{m}=0,$
for every positive integer $m\geq 0$, or using the symbol $O(\cdot)$ to denote,
$f(x)=C_{0}+\frac{C_{1}}{x}+\frac{C_{2}}{x^{2}}+\cdots+\frac{C_{m}}{x^{m}}+O\left(\frac{1}{x^{m+1}}\right).$
A linear system which has a pole at the origin written in the matrix expression is
$z^{h}\frac{d\bm{Y}}{dz}=\bm{B}(z)\bm{Y},\enspace h\in\mathbb{N},$ (2.4)
where $\bm{Y},\bm{B}$ are $n$-by-$n$ matrixes, $n\in\mathbb{N^{*}}$, and the matrix $\bm{B}(z)$ is holomorphic at $z=0$. When $h>1,$ the singular point is irregular. We want to make a transformation $x=1/z$ taking into the equation (2.4), affording
$x^{-d}\frac{d\bm{Y}}{dx}=\bm{A}(x)\bm{Y},$ (2.5)
with $d=h-2$ and $\bm{A}(x)=-\bm{B}(1/z)$. The integer $d+1$ is called the rank of the singular point. Then for $n$-by-$n$ constant matrixes $\bm{A}_{r}$, the matrix $\bm{A}(x)$ could have its expansion with form
$\bm{A}(x)=\sum_{j=0}^{\infty}\bm{A}_{j}x^{-j},\quad|x|>x_{0},$ (2.6)
which is holomorphic at $x=\infty$, or an expansion of this form asymptotic to $\bm{A}(x)$ in some sector $S$, i.e.
$\bm{A}(x)\sim\sum_{j=0}^{\infty}\bm{A}_{j}x^{-j},\enspace x\in S.$ (2.7)
We pay more interest to the latter case (2.7) and wonder if there exists an asymptotic expression of the solution $\bm{Y}$ in the same area. The matrix solution in the sector $S$ is summarized in the following theorem, see [14, p.111]:
Lemma 2.3
Let $\bm{A}(x)$ be an $n$-by-$n$ matrix function, which has the asymptotic form (2.7) in $S$. Then, the differential equation (2.5) possesses a fundamental matrix solution of the form
$\bm{Y}(x)=\hat{\bm{Y}}(x)x^{\bm{G}}e^{\bm{Q}(x)}$ (2.8)
corresponding to every sufficiently narrow subsector of $S$. Here $\hat{\bm{Y}}(x)$ permits an asymptotic series in power of $x^{-1/p},p\in\mathbb{N^{*}}$, in this subsector as $x\rightarrow\infty;$ $\bm{G}$ is a constant matrix, $\bm{Q}(x)$ is a diagonal matrix whose diagonal elements are polynomials in $x^{1/p}$.
Suppose a differential equation
$f^{(n)}(x)+A_{n-1}(x)f^{(n-1)}(x)+\cdots+A_{1}(x)f^{\prime}(x)+A_{0}(x)f(x)=0$ (2.9)
possesses coefficients $A_{k}(x)$ with the asymptotic expansions
$A_{k}(x)\sim x^{d_{k}}\sum_{j=0}^{\infty}A_{kj}x^{-j},\enspace x\in S,A_{kj}\in\mathbb{C},d_{k}\in\mathbb{N},k=0,\cdots,n-1.$
If we transform the (2.9) into a linear system, setting $\bm{y}=\{f,\ldots,f^{(n-1)}\}^{T},$ where $T$ represents the transpose of matrix, and $\tilde{\bm{A}}$ as the companion matrix
$\tilde{\bm{A}}=\begin{pmatrix}0&1&&&\\ &0&1&&\\ &&\ddots&\ddots&\\ &&&0&1\\ -A_{0}&-A_{1}&\cdots&-A_{n-2}&-A_{n-1}\end{pmatrix},$
it follows that (2.9) satifies the linear system
$x^{-d}\frac{d\bm{y}}{dx}=\bm{A}(x)\bm{y},\enspace x\in S,$
with $x^{d}\bm{A}=\tilde{\bm{A}}$ and integer $d=\max\limits_{k=0,\ldots,n-1}\{d_{k}\}$. Therefore, we transform equation (2.9) to a linear system (2.5) with $\bm{A}$ in the form of (2.7) after set $\bm{Y}=\{\bm{y}_{1},\ldots,\bm{y}_{n}\}$, $\bm{y}_{i}=\{f_{i},\ldots,f^{(n-1)}_{i}\}^{T},$ where $f_{i}$ are $n$ linearly independent solutions of (2.9), $i=1,\ldots,n$. The elements of $\bm{A}_{j}$ in (2.7) consisted of some of the $A_{kj},0,1$, and $k=0,\ldots,n-1,j\in\mathbb{N}.$
By Lemma 2.3 and fundamental calculations of elements of Matrixes, the $n$ linear independent solutions $f_{i}$ in $x$ in $S$ are
$f_{i}(x)=e^{\varLambda_{i}(x^{\frac{1}{p}})}x^{c_{i}}\varOmega_{i}(x,\log x),\quad 1\leq i\leq n.$ (2.10)
Here $\varLambda_{i}$ is a polynomial in $x^{1/p}$, $p\in\mathbb{N^{*}}$, $c_{i}\in\mathbb{C}$, and $\varOmega_{i}$ is a polynomial in $\log x$ with coefficients $\beta_{i,j}(x)$ possessing asymptotic forms
$\beta_{i,j}(x)\sim\sum_{s=0}^{\infty}\alpha_{i,j,s}x^{-s/p},\quad\alpha_{s}\in\mathbb{C},1\leq i\leq n,0\leq j\leq n-1.$
Repeating the proof of Lemma 2.2 given by Steinmetz [12], we get the following lemma:
Lemma 2.4
Let $f(x)$ be any solution of (2.9) in the sector $S$ and let $\arg z=\theta$ be an arbitrary direction in the interior of $S$. Then there is $p\in\mathbb{N},h>0,$ a polynomial $Q_{\theta}$ in $z^{1/p}$, such that
$\log|f(x)|=\Re Q_{\theta}(x^{1/p})+O(\log|x|)$ (2.11)
as $x\rightarrow\infty$ in $\theta\leq\arg x\leq\theta+h$ possibly outside an exceptional set consisting of
  1. 1.
    (countably many) disks $|x-x_{\mu}|<|x_{\mu}|^{1-\varepsilon},$ where $\varepsilon$ is positive and the counting function of the sequence {$x_{\mu}$} is $O(\log r)$;
  2. 2.
    a logarithmic semi-strip $0\leq\arg x-\theta<\dfrac{\log^{+}|x|}{|x|^{\frac{1}{p}}}$.
The latter occurs only if $\arg z=\theta$ is a Stokes ray in $S$.
Note that it is always possible to take the small sector in $S$ for $\theta\leq\arg x\leq\theta+h,h>0,$ when the ray $\arg z=\theta$ is in the interior of $S$. The polynomial $Q_{\theta}$ depending on $\theta$, but is piecewise constant and changes only when $\arg z=\theta$ is a Stokes ray.
3 Main results
3.1 LDE with exponential sums coefficients
An exponential sum is an entire function of the form
$F_{1}(z)e^{\lambda_{1}z}+\cdots+F_{m}(z)e^{\lambda_{m}z},$ (3.1)
where $\lambda_{1},\ldots,\lambda_{m}\in\mathbb{C}$ are distinct constants called leading coefficients (or frequencies) of the exponential sum (3.1), and the coefficients $F_{1}(z),\ldots,F_{m}(z)$ are polynomials.
Firstly, we consider the linear differential equation
$f^{(n)}(z)+A_{n-1}(z)f^{(n-1)}(z)+\cdots+A_{1}(z)f^{\prime}(z)+A_{0}(z)f(z)=0$ (3.2)
with exponential sums coefficients $A_{i}(z)$ for $i=0,1,\ldots,n-1$ of the form (3.1). The equation (3.2) can be transformed into a normal form
$\sum_{j=0}^{N_{0}}e^{\lambda_{j}z}\left(a_{n,j}(z)f^{(n)}(z)+a_{n-1,j}(z)f^{(n-1)}(z)+\cdots+a_{0,j}(z)f(z)\right)=0,\quad$ (3.3)
where $a_{n,j}(z)\in\{0,1\}$, $a_{n-1,j}(z),\ldots,a_{0,j}(z)$ are polynomials, and $\{\lambda_{j}\}$ is the set of all the mutually different leading coefficients of $A_{0}(z),\ldots,A_{n-1}(z)$.
$a_{n,j}(z)f^{(n)}(z)+a_{n-1,j}(z)f^{(n-1)}(z)+\cdots+a_{0,j}(z)f(z)=0$
are called coefficient differential equations about $\lambda_{j}$ (or at $\overline{\lambda}_{j}$) individually for $j=0,\ldots,N_{0}$. Set $W=\left\{\overline{\lambda}_{j}\right\}$, the convex hull $co(W)$ of $W$ is defined as the intersection of all closed convex sets containing $W$, also to be said as the convex hull to the differential equation (3.2). Suppose the vertexes of $co(W)$ consist a set $\tilde{W}=\{\overline{\omega}_{j}\}(j=0,\ldots,N_{0}^{\prime})$ with elements arranged anticlockwise, $N_{0}^{\prime}\leq N_{0}$, and denote $s_{j}$ the segment between $\overline{\omega}_{j}$ and $\overline{\omega}_{j+1}$. The critical ray $\arg z=\eta_{j}$ is defined as the ray originated at 0 with the direction of the outer normal to $s_{j}$. For small enough $\varepsilon>0,$ we define
$S_{j}(\varepsilon)=\{z|\eta_{j-1}+\varepsilon\leq\arg z\leq\eta_{j}-\varepsilon\},\enspace T_{j}(\varepsilon)=\{z|\eta_{j}-\varepsilon<\arg z<\eta_{j}+\varepsilon\}$
and
$\overline{S}_{j}=\{z|\eta_{j-1}\leq\arg z\leq\eta_{j}\}.$
See Figure 1.
[…]
Figure 1: $\tilde{W},\arg z=\eta_{j},S_{j}(\varepsilon),T_{j}(\varepsilon),\overline{S}_{j}$
Lemma 3.1
If $\overline{\lambda}\in{W},$ then for every $z\in S_{j}(\varepsilon)$,
$\Re((\omega_{j}-\lambda)z)\geq|z|\cdot|\omega_{j}-\lambda|\cdot|\sin\varepsilon|.$
If a given exponential point $\overline{\lambda}_{k}\notin s_{j}$, then for a given number $\psi_{\varepsilon}>0$ with $\psi_{\varepsilon}\neq 0$, such that for every $z\in\overline{S}_{j}\cap T_{j}(\theta)$ the inequality
$\Re((\omega_{j}-\lambda_{k})z)\geq|z|\cdot|\omega_{j}-\lambda_{k}|\cdot|\sin\psi_{\varepsilon}|$
is valid.
If a given exponential point $\overline{\lambda}_{k}\notin s_{j}$, then for a given number $\psi_{\varepsilon}>0$ with $\psi_{\varepsilon}\neq 0$, such that for every $z\in\overline{S}_{j+1}\cap T_{j}(\theta)$ the inequality
$\Re((\omega_{j}-\lambda_{k})z)\geq|z|\cdot|\omega_{j+1}-\lambda_{k}|\cdot|\sin\psi_{\varepsilon}|$
is valid.
See Dickson [2, Lemma, 1].
Our conclusions are as follows.
Theorem 3.2
Every finite order trancendental entire solution $f$ of (3.2) satisfies
$\log|f(z)|=\Re G_{j\theta}(z^{1/p_{j}})+O(\log|z|),$
for some polynomial $G_{j\theta}$ in $z^{1/p_{j}}$, $p_{j}\in\mathbb{N^{*}},j=1,\ldots,N_{0}^{\prime}$, outside a $r$-value set $E$ of finite linear measure, in $\theta\leq\arg z\leq\theta+h,\theta\in(0,2\pi]$, sufficiently small $h>0$, and enough large $|z|$, besides two kinds of areas:
  1. 1.
    $z$ in $T_{j}(\varepsilon)$, $\varepsilon>0$;
  2. 2.
    the logarithmic semi-strips
    $0\leq\arg z-\xi_{jk}<\frac{\log^{+}|z|}{{|z|}^{\frac{1}{p_{j}}}},k,p\in\mathbb{N^{*}}$
    in $S_{j}(\varepsilon)$, where $\arg z=\xi_{jk}$ are the Stokes rays to the coefficient differential equations at $\overline{\omega}_{j}$ in $S_{j}$,
$j=1,\ldots,N_{0}^{\prime}$ with $N_{0}^{\prime}$ the number of sides to the convex hull $co(W)$.
Thus, all finite order transcendental entire solutions are of completely regular growth.
Note. The polynomial $G_{j,\theta}$ will change according to the rotation of $\arg z=\theta$ from one side of $S_{j}$ to the other inside $S_{j}$, but this change occurs only when $\arg z=\theta$ is a Stokes ray in $S_{j},j=1,\ldots,N_{0}^{\prime}$.
Example 3.3
The function $f(z)=e^{-\frac{4}{3}z}(1-7e^{z})$ solves the differential equation
$f^{\prime\prime\prime}+3e^{z}f^{\prime\prime}+\left(-\frac{4}{3}-2e^{z}\right)f^{\prime}-\left(e^{z}-\frac{16}{27}\right)f=0.$
The equation can be transformed into its normal form
$\left(f^{\prime\prime\prime}-\frac{4}{3}f^{\prime}+\frac{16}{27}f\right)+e^{z}\left(3f^{\prime\prime}-2f^{\prime}-f\right)=0.$
Consider the differential polynomials in $g$,
$G_{1}(g)=\frac{g^{\prime\prime\prime}}{g}-\frac{4}{3}\frac{g^{\prime}}{g}+\frac{16}{27},\enspace\text{G}_{2}(g)=3\frac{g^{\prime\prime}}{g}-2\frac{g^{\prime}}{g}-1,$
and the equations $G_{1}=0,G_{2}=0.$ When we take $f(z)=e^{-\frac{4}{3}z}(1-7e^{z})$ into $G_{1}$ and $G_{2}$ individually, we find
$G_{1}(f(z))=1-\frac{1}{1-7e^{z}}\rightarrow 0,\enspace z\in S_{0}(\varepsilon)=\{z|\pi/2+\varepsilon\leq\arg z\leq 3\pi/2-\varepsilon\},$
and
$G_{2}(f(z))=\frac{7}{1-7e^{z}}\rightarrow 0,\enspace z\in S_{1}(\varepsilon)=\{z|-\pi/2+\varepsilon\leq\arg z\leq\pi/2-\varepsilon\}.$
They are both $o\left(\frac{1}{z^{N}}\right)$ for $N\in\mathbb{N}$.
Example 3.4
The differential equation
$04ze^{iz}f^{\prime\prime\prime}+(6e^{iz}+4ize^{iz}-z)f^{\prime\prime}+\left((2i-1)e^{iz}-\frac{1}{2}\right)f^{\prime}-\left(ie^{iz}-\frac{1}{4}\right)f=$ 0 (3.4)
is given. The normal form of (3.4) is
$e^{iz}\{4zf^{\prime\prime\prime}+(6+4iz)f^{\prime\prime}+(2i-1)f^{\prime}-if\}+\left(-zf^{\prime\prime}-\frac{1}{2}f^{\prime}+\frac{1}{4}f\right)=0.$
The solution $f=\cosh\sqrt{z}$ of (3.4) solves both equations
$04zf^{\prime\prime\prime}+(6+4iz)f^{\prime\prime}+(2i-1)f^{\prime}-if=0,\quad-zf^{\prime\prime}-\frac{1}{2}f^{\prime}+\frac{1}{4}f=$ 0 .
3.2 LDE with exponential polynomials coefficients
For an exponential polynomial (1.1), at least one of the degrees of the polynomials $Q_{j},j=1,\ldots,l,$ is bigger than one, or it is an exponential sum or a polynomial. Set $s=\max\{\deg(Q_{j})\}\geq 1.$ An exponential polynomial can be rewritten in the normalized form
$G(z)=\hat{G_{1}}(z)+\sum_{j=2}^{k}\hat{G}_{j}(z)\exp\left({q_{j}z^{s}}\right)$ (3.5)
where $k\in\mathbb{N},$ the $q_{j}$ are pairwise distinct non-zero constants for $2\leq j\leq k$, and $q_{1}=0;$ the coefficients $\hat{G_{j}}(z)$ are exponential polynomials of growth order $\rho(\hat{G_{j}})\leq s-1$ such that $\hat{G_{j}}(z)\not\equiv 0,$ for $2\leq j\leq{k},$ and ${k}\leq l.$
Now, we step forward to deal with the equation with exponential polynomials coefficients. A differential equation
$f^{(n)}(z)+G_{n-1}(z)f^{(n-1)}(z)+\cdots+G_{1}(z)f^{\prime}(z)+G_{0}(z)f(z)=0$ (3.6)
with the exponential polynomials coefficients
$G_{m}(z)=\sum_{t=1}^{l_{m}}P_{m,t}(z)\exp(Q_{m,t}(z)),\enspace 0\leq m\leq n-1,\enspace l_{m}\in\mathbb{N},$ (3.7)
where $P_{m,t}$ and $Q_{m,t}$ are polynomials, can be transformed into the equation
$\sum_{j=1}^{k}\exp\left(q_{j}z^{s}\right)\left\{\sum_{m=0}^{n}\hat{G}_{m,j}(z)f^{(m)}(z)\right\}=0,$ (3.8)
where $k$ is the maximal number of pairwise different complex number $q_{j},$ appearing in all coefficients of $z^{s}$ in $Q_{m,t}$, $q_{1}=0$ and $s=\max\{\deg(Q_{m,t})\}$. Thereby, the degree of every polynomial $\hat{G}_{m,j}(z)$ is smaller or equal to $s-1.$
The complex numbers ${q_{j}}$ are called leading coefficients to the degree $s$. The differential equation
$\sum_{m=0}^{n}\hat{G}_{m,j}(z)f^{(m)}(z)=0$ (3.9)
is called a coefficient differential equation at the conjugated leading coefficient $\overline{q}_{j}$ of (3.6).
To describe logically, we denote the convex hull to equation (3.8) as $co(W)$, conjugated leading coefficients as $\overline{q}_{j_{1}},j_{1}=1,\ldots,k$; denote the $r_{1}$ many vertexes of $co(W)$ as $\overline{\omega}_{j_{1}},1\leq j_{1}\leq r_{1}\leq k,$ in counterclockwise order and critical rays $\arg z=\eta_{j_{1}},$ defined as the ray originated at 0 with the direction of the outer normal to the segment between $\overline{\omega}_{j_{1}}$ and $\overline{\omega}_{j_{1}+1}$; denote sectors $S_{j_{1}}(\varepsilon),T_{j_{1}}(\varepsilon)$ as
$S_{j_{1}}(\varepsilon)=\{z|\eta_{j_{1}-1}+\varepsilon\leq\arg z\leq\eta_{j_{1}}-\varepsilon\},\enspace T_{j_{1}}(\varepsilon)=\{z|\eta_{j_{1}}-\varepsilon<\arg z<\eta_{j_{1}}+\varepsilon\},$
$1\leq j_{1}\leq r_{1}\leq k$.
[…]
Figure 2: leading coefficients $\{q\}$ of corresponding exponential polynomial terms
In Figure 2, $n(\cdot)$ is a number controlled by the variables ``"``"`$\cdot$"` ` ⋅ ". To write inductively, we set $n(j_{1},\ldots,j_{t-1})=k_{j_{1},\ldots,j_{t-1}}$. We have already separated the dominant exponential polynomial terms $e^{q_{j_{1}}z^{s}}$ in (3.8) and their coefficient differential equation (3.9) at $q_{j_{1}}$. Within each coefficient differential equation (3.9) at $q_{j_{1}}$, we can do this transformation again and sparate the dominant exponential polynomial terms $e^{q_{j_{1},j_{2}}z^{s-1}}$ and their coefficient differential equations at $q_{j_{1},j_{2}}$, where $q_{j_{1},j_{2}}$ are leading coefficients to the degree $s-1$. We denote the convex hull to equation (3.9) as $co(W_{j_{1}}),j_{1}=1,\ldots,k$; denote the $r_{j_{1},j_{2}}$ many conjugated leading coefficients on the vertexes of convex hull $co(W_{j_{1}})$ as $\overline{\omega}_{j_{1},j_{2}},j_{1}=1,\ldots,k,j_{2}=1,\ldots,k_{j_{1}}$ in counterclockwise order, where $k_{j_{1}}$ is the maximal number of pairwise different complex number $q_{j_{1},j_{2}}$; denote the $r_{j_{1},j_{2}}$ many critical rays $\arg z=\eta_{j_{1},j_{2}},$ defined as the ray originated at 0 with the direction of the outer normal to the segment between $\overline{\omega}_{j_{1},j_{2}}$ and $\overline{\omega}_{j_{1},j_{2}+1}$, corresponding sectors as
$S_{j_{1},j_{2}}(\varepsilon)=\{z|\eta_{j_{1},j_{2}-1}+\varepsilon\leq\arg z\leq\eta_{j_{1},j_{2}}-\varepsilon\},\enspace T_{j_{1},j_{2}}(\varepsilon)=\{z|\eta_{j_{1},j_{2}}-\varepsilon<\arg z<\eta_{j_{1},j_{2}}+\varepsilon\},$
$1\leq j_{1}\leq r_{1}\leq k,1\leq j_{2}\leq r_{j_{1}}\leq k_{j_{1}}$.
Repeating the division of original equation by the dominant exponential polynomial terms $e^{q_{j_{1}}z^{s}},e^{q_{j_{1},j_{2}}z^{s-1}},\ldots,e^{q_{j_{1},\ldots,j_{s}}z}$ for $s$ times inductively, if we retain the symbols and get a coefficient differential equation with only polynomials coefficients, we call it a fundamental coefficient differential equation at $\overline{\omega}_{j_{1},\ldots,j_{s}},1\leq j_{t}\leq r_{j_{1},\ldots,j_{t-1}}\leq k_{j_{1},\ldots,j_{t-1}},1\leq t\leq s,r_{j_{0}}=r_{1},k_{j_{0}}=k$. $k_{j_{1},\ldots,j_{t}}$ is the maximal number of pairwise different complex number $q_{j_{1},\ldots,j_{t}},$ see Figure 2, the $r_{j_{1},\ldots,j_{t}}$ many vertexes of $co(W_{j_{1},\ldots,j_{t-1}})$ are denoted by $\overline{\omega}_{j_{1},\ldots,j_{t}}$ in counterclockwise order and critical rays $\arg z=\eta_{j_{1},\ldots,j_{t}},$ defined as the ray originated at 0 with the direction of the outer normal to the segment between $\overline{\omega}_{j_{1},\ldots,j_{t}}$ and $\overline{\omega}_{j_{1},\ldots,j_{t}+1}$, corresponding sectors as
$\begin{split}S_{j_{1},\cdots,j_{t}}(\varepsilon)&=\{z|\eta_{j_{1},,\cdots,j_{t}-1}+\varepsilon\leq\arg z\leq\eta_{j_{1},\cdots,j_{t}}-\varepsilon\},\\ \ T_{j_{1},\cdots,j_{t}}(\varepsilon)&=\{z|\eta_{j_{1},\cdots,j_{t}}-\varepsilon<\arg z<\eta_{j_{1},\cdots,j_{t}}+\varepsilon\},\end{split}$
where $1\leq j_{t}\leq r_{j_{1},\ldots,j_{t-1}}\leq k_{j_{1},\ldots,j_{t-1}},1\leq t\leq s.$ See Figure 3.
$\overline{\omega}_{j_{1},\ldots,j_{t}+1}$$\overline{\omega}_{j_{1},\ldots,j_{t}}$$\overline{\omega}_{j_{1},\ldots,j_{t}-1}$$\overline{\omega}_{j_{1},\ldots,j_{t-1},1}$$\overline{\omega}_{j_{1},\ldots,j_{t-1},r_{j_{1},\ldots,j_{t-1}}}$${co(W_{j_{1},\ldots,j_{t-1}})}$$\arg z=\eta_{j_{1},\ldots,j_{t}}$$\varepsilon$$\varepsilon$$\varepsilon$$S_{j_{1},\ldots,j_{t}}(\varepsilon)$$T_{j_{1},\ldots,j_{t}}(\varepsilon)$$\arg z=\eta_{j_{1},\ldots,j_{t}-1}$$\eta_{j_{1},\ldots,j_{t-1},1}$$\eta_{j_{1},\ldots,j_{t}+1}$$\eta_{j_{1},\ldots,j_{t-1},r_{j_{1},\ldots,j_{t-1}}}$
Figure 3: ${co(W_{j_{1},\ldots,j_{t-1}})},\arg z=\eta_{j_{1},\ldots,j_{t}},S_{j_{1},\ldots,j_{t}}(\varepsilon),T_{j_{1},\ldots,j_{t}}(\varepsilon)$
Additionally, we need some relations about the distribution of $z$ and $z^{s}$, where $s\in\mathbb{C}$. Let $10\leq\alpha<$ 1 and
$\Pi:=\{z\in\mathbb{C}||\Im(z)|<{|z|}^{\alpha}\}.$
If $\Phi$ is formed from $\Pi$ by rotating at a fixed angle around the origin, then $\Phi$ is called the parabolic strip of aperture $\alpha$ along the axis that by rotation around the same angle arising from the positive real axis.
The Lemma is given originally by Droletz [3] in his dissertation, which is compiled below.
Lemma 3.5
Let $s\in\mathbb{N},$ and $\Phi$ be a parabolic strip of aperture $\alpha$ with $z^{s}\in\Phi$. Then there are parabolic strips $\Phi^{(1)},\Phi^{(2)},\ldots,\Phi^{(s)}$, of aperture $\alpha$, such that for a sufficiently large $|z|$ the value $z$ lies in the union
$\bigcup_{k=1}^{s}\Phi^{(k)}.$
Moreover, if the argument to the axis of symmetry of $\Phi$ is $s\eta,$ then the argument to the axis of $\Phi^{(k)}$ are $\eta+\frac{2k\pi}{s}$ individually, for $k=1,\ldots,s.$
Now a similar conclusion to Theorem 3.2 is given for exponential polynomials coefficients.
Theorem 3.6
Every finite order trancendental entire solution $f$ of (3.6) satisfies
$\log|f(z)|=\Re G_{\theta}(z^{1/N_{\theta}})+O(\log|z|),$ (3.10)
for some polynomials $G_{\theta}$ in $z^{1/N_{\theta}}$, $N_{\theta}\in\mathbb{N^{*}},$ outside a $r$-value set $E$ of finite linear measure, in $\theta\leq\arg z\leq\theta+h,\theta\in(0,2\pi]$, sufficiently small $h>0$, and enough large $|z|$, besides two kinds of areas:
  1. 1.
    $z$ in
    $T_{j_{1},\ldots,j_{t}}^{(i)}(\varepsilon)=\left\{z:\frac{\eta_{j_{1},\cdots,j_{t}}+2i\pi-\varepsilon}{s+1-t}<\arg z<\frac{\eta_{j_{1},\cdots,j_{j}}+2i\pi+\varepsilon}{s+1-t}\right\},$
    where $1\leq i\leq s+1-t,1\leq t\leq s;$ $\varepsilon>0$, $1\leq j_{t}\leq r_{j_{1},\ldots,j_{t-1}}\leq k_{j_{1},\ldots,j_{t-1}},1\leq t\leq s$; $\eta_{j_{1},\ldots,j_{t}}$ are arguments of critical rays defined above;
  2. 2.
    the logarithmic semi-strips
    $0\leq\arg z-\xi_{j_{1},\cdots,j_{s},M_{{j_{1},\cdots,j_{s}}}}<\frac{\log^{+}|z|}{{|z|}^{\frac{1}{p_{j_{1},\cdots,j_{s}}}}},p_{j_{1},\cdots,j_{s}}\in\mathbb{N^{*}},M_{j_{1},\cdots,j_{s}}\in\mathbb{N}$
    in
    $\mathcal{S}_{j_{1},\cdots,j_{s}}(\varepsilon)=\bigcap_{t=1,2,\cdots,s}\left(\bigcup_{i=1,\cdots,s+1-t}S_{j_{1},\cdots,j_{t}}^{(i)}(\varepsilon)\right),$
    where $\arg z=\xi_{j_{1},\cdots,j_{s},M_{{j_{1},\cdots,j_{s}}}}$ are the Stokes rays to the coefficient differential equations at $\overline{\omega}_{j_{1},\ldots,j_{s}}$ in $\mathcal{S}_{j_{1},\cdots,j_{s}}(\varepsilon).$
Thus, all finite order transcendental entire solutions are of completely regular growth.
Note. The polynomial $G_{\theta}$ is piecewise constant and changes when $\arg z=\theta$ is a Stokes ray, but these changes occur only for finite many times.
Example 3.7
The function $f(z)=e^{z^{2}}$ solves the equation
$\begin{split}(2z-e^{z}+\cos(z^{2}))f^{\prime\prime}+(e^{z}+e^{2z}-2-4z^{2}-2z\sin(z^{2})-2z\cos(z^{2}))f^{\prime}+&\\ (2e^{z}+4z^{2}e^{z}-2ze^{z}-2ze^{2z}+4z^{2}\sin(z^{2})-2\cos(z^{2}))f=0.\end{split}$ (3.11)
The normal form of (3.11) is
$\begin{split}&e^{iz^{2}}\left\{\frac{1}{2}f^{\prime\prime}+(iz-z)f^{\prime}-(2iz^{2}+1)f\right\}+\\ &e^{-iz^{2}}\left\{\frac{1}{2}f^{\prime\prime}-(iz+z)f^{\prime}+(2iz^{2}-1)f\right\}+\\ &e^{2z}\left\{f^{\prime}-2zf\right\}+e^{z}\left\{-f^{\prime\prime}+f^{\prime}+(2-2z+4z^{2})f\right\}+\left\{2zf^{\prime\prime}-(2+4z^{2})f^{\prime}\right\}=0.\end{split}$
Set the differential polynomial
$\begin{split}&G_{1}(f)=\frac{1}{2}\frac{f^{\prime\prime}}{f}+(iz-z)\frac{f^{\prime}}{f}-(2iz^{2}+1),\\ &G_{2}(f)=\frac{1}{2}\frac{f^{\prime\prime}}{f}-(iz+z)\frac{f^{\prime}}{f}+(2iz^{2}-1),\\ &G_{3}(f)=\frac{f^{\prime}}{f}-2z,G_{4}(f)=-\frac{f^{\prime\prime}}{f}+\frac{f^{\prime}}{f}+2-2z+4z^{2},G_{5}(f)=2z\frac{f^{\prime\prime}}{f}-(2+4z^{2})\frac{f^{\prime}}{f},\end{split}$
and we afford $G_{i}(e^{z^{2}})=0,i=1,\ldots,5.$
4 Proofs
4.1 Proof of Theorem 3.2
Proof. Dividing $f$ of both sides of (3.3), we get
$\sum_{j=0}^{N_{0}}e^{\lambda_{j}z}\left(a_{n,j}f^{(n)}/f+a_{n-1,j}(z)f^{(n-1)}/f+\cdots+a_{0,j}(z)\right)=0.$ (4.1)
For $k,j\in 0,\ldots,N_{0}$ we set
$\kappa_{k,j}(z):=\left|\frac{e^{\lambda_{k}z}\left(\sum\limits_{t=0}^{n}a_{t,k}\dfrac{f^{(t)}}{f}\right)}{e^{\omega_{j}z}\left(\sum\limits_{t=0}^{n}a_{t,j}\dfrac{f^{(t)}}{f}\right)}\right|=e^{\Re{(\lambda_{k}-\omega_{j})}z}\left|\frac{\left(\sum\limits_{t=0}^{n}a_{t,k}\dfrac{f^{(t)}}{f}\right)}{\left(\sum\limits_{t=0}^{n}a_{t,j}\dfrac{f^{(t)}}{f}\right)}\right|.$ (4.2)
From Lemma 3.1, for $z\in S_{j}(\varepsilon)$, $\lambda_{k}\neq\omega_{j}$ ($\omega_{j}$ is a vertex of the minimal convex set containing all the $\lambda_{k},k=0,\ldots,N_{0}$), it follows that
$\Re((\lambda_{k}-\omega_{j})z)\leq-|z|\cdot|\lambda_{k}-\omega_{j}|\cdot|\sin\varepsilon|<0.$
By Gundersen [4, Corollary 3], if we set $\rho=\rho(f)$ as growth order of $f$, there exists $\varepsilon_{0}>0$ and a set $E$ with finite linear measure, such that
$\left|\frac{f^{(k)}}{f}\right|<{|z|^{(k-1)(\rho+\varepsilon_{0})}},\quad z\notin E.$
Then, we claim that the equality below is true:
$\sum\limits_{t=0}^{n}a_{t,j}\dfrac{f^{(t)}}{f}=o\left(\frac{1}{z^{N_{1}}}\right),\enspace\text{for any}\enspace N_{1}\in\mathbb{N},\enspace z\in S_{j}(\varepsilon),\enspace|z|\notin E.$ (4.3)
Otherwise, there is a sequence $\{z_{i}\},z_{i}\in S_{j}(\varepsilon),|z_{i}|\notin E$, with $\lim\limits_{i\rightarrow\infty}z_{i}=\infty,$ such that
$\left|\sum\limits_{t=0}^{n}a_{t,j}(z_{i})\frac{f^{(t)}(z_{i})}{f(z_{i})}\right|>\frac{C}{{|z_{i}|}^{N_{1}}},$ (4.4)
for any $C,N_{1}\in\mathbb{N}$. It holds that $\kappa_{k,j}(z_{i})=o(1),$ for the reason that $e^{\Re{(\lambda_{k}-\omega_{j})}z}$ decreases exponentially and the term in the absolute value in (4.2) is up to a polynomial rate of growth. From (4.1),
$e^{\lambda_{k}z_{i}}\left(\sum\limits_{t=0}^{n}a_{t,k}(z_{i})\dfrac{f^{(t)}(z_{i})}{f(z_{i})}\right)=o\left(e^{\omega_{j}z_{i}}\left(\sum\limits_{t=0}^{n}a_{t,j}(z_{i})\dfrac{f^{(t)}(z_{i})}{f(z_{i})}\right)\right),$
which leads that for large $|z_{i}|$,
$e^{\omega_{j}z_{i}}\left(\sum\limits_{t=0}^{n}a_{t,j}(z_{i})\dfrac{f^{(t)}(z_{i})}{f(z_{i})}\right)(1+o(1))=0,$
a contradiction to the assumption (4.4).
Therefore, for $z\in S_{j}(\varepsilon),|z|\notin E$, (4.3) can be transformed into an equation
$f^{(n)}+P_{n-1,j}f^{(n-1)}+\cdots+P_{1,j}f^{\prime}+\left(P_{0,j}+o\left(\frac{1}{z^{N_{1}+\deg(a_{0,j})}}\right)\right)f=0,$ (4.5)
where $P_{i,j}=a_{i,j}/a_{n,j},i=0,\ldots,n-1$. Briefly, denote $Q_{0,j}$ as the last coefficient in equation (4.5). Furthermore, it is deduced that every $P_{i,j}$ can be expanded as
$P_{i,j}(z)=z^{d_{i,j}}\sum_{v=0}^{\infty}\frac{b_{i,j,v}}{z^{v}},\quad|z|>|z_{0}|,$
for $d_{i,j}\in\mathbb{N},b_{i,j,v}\in\mathbb{C},i=1,\ldots,n-1,$ $j=0,\ldots,N_{0}^{\prime},$ enough large $|z_{0}|$, $z\in S_{j}(\varepsilon),$ and $|z|\notin E$. $Q_{0}$ is consequently in the asymptotic form of
$Q_{0,j}(z)\sim z^{d_{0,j}}\sum_{v=0}^{\infty}\frac{b_{0,j,v}}{z^{v}},\quad|z|>|z_{0}|,$
for $d_{0,j}\in\mathbb{N},b_{0,j,v}\in\mathbb{C},$ $j=0,\ldots,N_{0}^{\prime},$ enough large $|z_{0}|$, $z\in S_{j}(\varepsilon),$ and $|z|\notin E$.
It is obtained that (4.5) is in the form of (2.9). Using Lemma 2.4, there exists a polynomial $G_{\theta}$ in $z^{1/p_{j}}$ and a a set of zero relative measure $E_{\theta}$ such that
$\log|f(z)|=\Re G_{\theta}\left(z^{\frac{1}{p_{j}}}\right)+O(\log|z|)$ (4.6)
for $|z|\notin E\cup E_{\theta},z\in S_{j}(\varepsilon)$ as $z\rightarrow\infty$ in $\theta\leq\arg z\leq\theta+h,h>0$, and outside of a logarithmic semi-strip $0\leq\arg z-\theta<C\frac{\log^{+}|z|}{|z|^{1/p_{j}}}$. The last semi-strip occurs only if $\arg z=\theta$ is a Stokes ray in $S_{j}$, otherwise, we set $C=0$. Thus, it is natural to be inferred $f(z)$ is of completely regular growth on every ray $\theta\in S_{j}(\varepsilon)\backslash\{\arg z=\xi_{j1},\ldots,\arg z=\xi_{j_{m_{j}}}\},j=0,\ldots,N_{0},$ where $\{\arg z=\xi_{j1},\ldots,\arg z=\xi_{j_{m_{j}}}\}$ are the Stokes rays of equation (4.5) in $S_{j}(\varepsilon),m_{j}\in\mathbb{N}.$
Except for Stokes rays, we assert that $f(z)$ is of completely regular growth on every ray $\theta=\arg z$ except for the critical rays $\arg z=\eta_{j},j=0,\ldots,N_{0}.$ Since
$S_{j}(\varepsilon)=\{z|\eta_{j-1}+\varepsilon\leq\arg z\leq\eta_{j}-\varepsilon\},$
if there exists a $\eta_{j}-\varepsilon<\theta_{0}<\eta_{j}$ such that $f(z)$ is not of completely regular growth on the ray $\arg z=\theta_{0},$ we can always find a $\varepsilon_{1}<\eta_{j}-\theta_{0},$ with $f$ of completely regular growth on the ray in
$S_{j}(\varepsilon_{1})=\{z|\eta_{j-1}+\varepsilon_{1}\leq\arg z\leq\eta_{j}-\varepsilon_{1}\},$
a contradiction.
That means the set of arguments of completely regular growth
$\mathbb{M}=(0,2\pi]\backslash\{\eta_{0},\ldots,\eta_{N_{0}},\xi_{01},\ldots,\xi_{0{m_{0}}},\ldots,\xi_{N_{0},1},\ldots,\xi_{N_{0}{{m_{N_{0}}}}}\}.$
Then, by Lemma 2.1, $f(z)$ is of completely regular growth on the set of rays with arguments $\overline{\mathbb{M}}=(0,2\pi]$. $\Box$
4.2 Proof of Theorem 3.6
Proof. Change (3.6) to the normalized form (3.8). Dividing $f$ of both sides of (3.8), we afford
$\sum_{j=1}^{k}\exp\left(q_{j}z^{s}\right)\left\{\sum_{m=0}^{n}\hat{G}_{m,j}(z)\frac{f^{(m)}(z)}{f(z)}\right\}=0,$ (4.7)
for every $a\in\mathbb{R}.$ Denote $H_{j}$ by
$H_{j}(z)=\sum_{m=0}^{n}\hat{G}_{m,j}(z)\frac{f^{(m)}(z)}{f(z)}=0,$
and then (4.7) becomes
$\sum_{j=1}^{k}H_{j}(z)\exp(q_{j}z^{s})=0.$ (4.8)
The symbols $r,k,q,H,\omega,co(\cdot),S(\varepsilon),T(\varepsilon)$ here below refer to the definition in Section 3.2. Also, set $L_{p}$ as the side between the vertex point $\overline{\omega}_{p}$ and $\overline{\omega}_{{p-1}}$ of the convex hull $co(W)$, and $\Phi_{p}$ and is the parabolic strip of a sufficiently large opening along the critical ray $\arg z=\eta_{p}$ as an axis of symmetry. There is a sequence $\{z_{\mu}\}_{\mu\in\mathbb{N}}$ with $\lim_{\mu\to\infty}z_{\mu}=\infty,$ such that every $z_{\mu}^{s}$ lies in $S_{j}(\varepsilon)$. As a result of
$S_{j}(\varepsilon)\cap(\Phi_{p}\cup\Phi_{{p-1}})=\emptyset$
for $|z_{\mu}^{s}|\rightarrow\infty$, $\{z_{\mu}^{s}\}$ is contained in the area between $\Phi_{p}$ and $\Phi_{{p-1}}$.
Furthermore, assuming there is a $A\in\mathbb{R}$ such that for a sufficiently large sequence $|z_{\mu}|=r_{\mu}$,
$|H_{p}(z_{\mu})|>ce^{Ar_{\mu}^{s-1}}$ (4.9)
is true with an appropriate $c>0$. Therefore, for $t,p=1,\ldots,k,$ we find
$\kappa_{t}(z_{\mu}):=\left|\frac{H_{t}(z_{\mu})e^{q_{t}z_{\mu}^{s}}}{H_{p}(z_{\mu})e^{q_{p}z_{\mu}^{s}}}\right|=\left|\frac{H_{t}(z_{\mu})}{H_{p}(z_{\mu})}\right|e^{\Re{((q_{t}-q_{p})z_{\mu}^{s}})}<\frac{|H_{t}(z_{\mu})|}{c}e^{\Re{((q_{t}-q_{p})z_{\mu}^{s}})-Ar_{\mu}^{s-1}},t\not=p.$
Further, three circumstances will show up.
  1. Case 1.
    If $\overline{q}_{t}$ does not lie on $L_{p}$ or $L_{p+1}.$ Then it follows from Lemma 3.1:
    $\Re((q_{t}-q_{p})z_{\mu}^{s})\leq-|z_{\mu}^{s}||q_{t}-q_{p}||\sin\varepsilon|,$
    with $z_{\mu}^{s}\in S_{p}(\varepsilon)$ a sufficiently small $\varepsilon>0.$ Then it follows
    $\kappa_{t}(z_{\mu})\leq\exp(-|z_{\mu}^{s}||q_{t}-q_{p}||\sin\varepsilon|-(A-B)r_{\mu}^{s-1})=o(1),$
    for a $B\in\mathbb{R}$, except for a set $E$ in $r=|z|$ of finite linear measure, since $f^{(m)}/f$ is of polynomial growth except for some set $E$ in $r=|z|$ of finite linear measure, see Gundersen [4, Corollary 3], such that
    $\frac{1}{c}|H_{t}(z_{\mu})|=\frac{1}{c}\left|\sum_{m=0}^{n}\hat{G}_{m,j}(z_{\mu})\frac{f^{(m)}(z_{\mu})}{f(z_{\mu})}\right|\leq e^{Br_{\mu}^{s-1}},$ (4.10)
    with $|z_{\mu}|\notin E,z_{\mu}\rightarrow\infty,t\neq p$.
  2. Case 2.
    If $\overline{q}_{t}$ is on $L_{p},t\neq p$. Without loss of generality, we assume the axis of symmetry of $\Phi_{p}$ is the positive axis. Then $q_{t}-q_{p}$ is imaginary and thus
    $q_{t}-q_{p}=-i|q_{t}-q_{p}|.$
    For the reason that $\{z_{\mu}\}$ are not in the parabolic strips, it is followed by the definition of parabolic strips that
    $\kappa_{t}(z_{\mu})<$ $\frac{|H_{t}(z_{\mu})|}{c}e^{-{(|q_{t}-q_{p}|\Im(z_{\mu}^{s}}))-Ar_{\mu}^{s-1}}$
    $\leq$ $\frac{|H_{t}(z_{\mu})|}{c}e^{\left(-{|q_{t}-q_{p}|r_{\mu}^{s\alpha}}\left(1-A_{1}r_{\mu}^{s\left(1-\alpha-\frac{1}{s}\right)}\right)\right)}.$
    Now, $\alpha$ can be chosen from the interval $(1-1/s,1)$ and
    $\kappa_{t}(z_{\mu})=o(1).$ (4.11)
  3. Case 3.
    If $\overline{q}_{t}$ lies on $s_{p+1},t\neq p.$ As in Case 2, we have $\kappa_{t}(z_{\mu})=o(1).$
Together with (4.8),(4.10),(4.11),
$\sum_{j=1}^{k}H_{j}(z_{\mu})e^{q_{t_{s}}z_{\mu}^{s}}=(1+o(1))H_{p}(z_{\mu})e^{q_{p_{s}}z_{\mu}^{s}}$ (4.12)
for every $a\in\mathbb{R},z_{\mu}\in S_{p}(\varepsilon),|z_{\mu}|\notin E.$ Then for (4.9), we have
$\frac{\left|\sum\limits_{j=1}^{k}H_{j}(z_{\mu})e^{q_{j}z^{s}_{\mu}}\right|}{\left|e^{q_{p}z_{\mu}^{s}}\right|}\geq c(1+o(1))e^{Ar_{\mu}^{s-1}},$ (4.13)
where $c>0,z_{\mu}\in S_{p}(\varepsilon),|z_{\mu}|\notin E.$ Then it follows
$\left|\sum_{j=1}^{k}H_{j}(z_{\mu})e^{q_{j}z_{m}u^{s}}\right|\geq c(1+o(1))e^{Ar_{\mu}^{s-1}-|q_{p}z_{\mu}^{s}|},\enspace|z_{\mu}|\notin E,$
but it contradicts (4.8). Consequently, there are only two possibilities:
  1. (I)
    The assumption (4.9) is incorrect. Then for every $a\in\mathbb{R}$, vertex point $\overline{\omega}_{p}$ of $co(W)$, we have
    $H_{p}(z_{\mu})=o(e^{ar_{\mu}^{s-1}}),z_{\mu}\in S_{p}(\varepsilon).$ (4.14)
  2. (II)
    The sequence $z_{\mu}^{s}$ lies in $T_{p}$ for a sufficiently large $|z_{\mu}|$, the axis of symmetry is the critical ray $\arg z^{s}_{\mu}=\eta_{p}$ of the convex hull $co(W)$. Thus, there are $s$ subsequences $\{z_{\mu 1},\ldots,z_{\mu s}\}$ of the sequence $\{z_{\mu}\}(\mu\in\mathbb{N})$ and $s$ branches $T_{p}^{(1)},\ldots,T_{p}^{(s)}$ of $T_{p}$, such that $z_{\mu j}$ lies in subsectors
    $T_{p}^{(j)}=\left\{z:\frac{\eta_{p}+2j\pi-\varepsilon}{s}<\arg z<\frac{\eta_{p}+2j\pi+\varepsilon}{s}\right\},\quad 1\leq j\leq s.$
Now we set $p=p_{1},$ $H_{p_{1},p_{2}}$ be a term of the coefficient differential equation to $\overline{\omega}_{{p_{1},p_{2}}}$, the vertex of the convex hull of $co(W_{p_{1}})$. If $W_{p_{1}}$ vanishes, we can take $p_{1}=1$ such that the coefficient differential equation does not disappear. We may always denote the subsequence of $\{z_{\mu}\}$ as $\{z_{\mu}\}$, consequently arrive to
$H_{p_{1},p_{2}}(z_{\mu})=o\left(e^{ar_{\mu}^{s-2}}\right),z_{\mu}^{s}\in S_{p_{1}}(\varepsilon)\cap S_{p_{1},p_{2}}(\varepsilon),$
where $j=1,\ldots,s;$ $i=1,\ldots,s-1;1\leq p_{2}\leq r_{p_{1}}\leq k_{p_{1}}.$ Repeatedly, we get a relationship
$H_{p_{1},p_{2},\cdots,p_{s}}(z_{\mu})=o\left(e^{ar_{\mu}^{0}}\right)=o(1),a\in\mathbb{R},z_{\mu}^{s}\in\bigcap_{j=1,2,\cdots,s}S_{p_{1},\cdots,p_{j}}(\varepsilon),$
Therefore, there exist the sequence $\{z_{\mu}\}(\mu\in\mathbb{N})$ and $i$ branches $T_{p_{1},\ldots,p_{j}}^{(1)},\ldots,T_{p_{1},\ldots,p_{j}}^{(i)}$ of $T_{p_{1},\ldots,p_{j}}(\varepsilon)$, such that $z_{\mu}$ does not lie in subsectors
$T_{p_{1},\ldots,p_{j}}^{(i)}(\varepsilon)=\left\{z:\frac{\eta_{p_{1},\cdots,p_{j}}+2i\pi-\varepsilon}{s+1-j}<\arg z<\frac{\eta_{p_{1},\cdots,p_{j}}+2i\pi+\varepsilon}{s+1-j}\right\},$
where $1\leq i\leq s+1-j,1\leq j\leq s.$
We note that here $H_{p_{1},p_{2},\cdots,p_{s}}(z_{\mu})=o(z_{\mu}^{-n_{0}})$ for any $n_{0}\in\mathbb{N}$. Otherwise, we will get a contradiction with the same analysis under (4.3). Since $H_{p_{1},p_{2},\cdots,p_{s}}(z)=0$ is a linear differential equation with polynomial coefficients, we can construct an equation which coefficients have asymptotic forms in sectors
$\bigcap_{j=1,2,\cdots,s;\atop i=1,\cdots,s+1-j}S_{p_{1},\cdots,p_{j}}^{(i)}(\varepsilon),$
where
$S_{p_{1},\ldots,p_{j}}^{(i)}(\varepsilon)=\left\{z:\frac{\eta_{p_{1},\cdots,p_{j}-1}+2i\pi+\varepsilon}{s+1-j}\leq\arg z\leq\frac{\eta_{p_{1},\cdots,p_{j}}+2i\pi-\varepsilon}{s+1-j}\right\}.$
By the similar proof of theorem 3.2 and Lemma 2.4, we afford
$\log|f(z_{\mu})|=\Re G_{\theta}(z_{\mu}^{1/N_{\theta}})+O(\log|z_{\mu}|),z_{\mu}\in\bigcap_{j=1,2,\cdots,s}\left(\bigcup_{i=1,\cdots,s+1-j}S_{p_{1},\cdots,p_{j}}^{(i)}(\varepsilon)\right),$ (4.15)
with some polynomial $G_{\theta}$ in $z^{1/N_{\theta}}$, for $N_{\theta}\in\mathbb{N},$ $\theta\in(0,2\pi],$ $|z_{\mu}|\notin E$, as $z_{\mu}\rightarrow\infty,$ in $\theta\leq\arg z_{\mu}\leq\theta+h,h>0,$ outside of a logarithmic semi-strip $0\leq\arg z_{\mu}-\theta<C\frac{\log^{+}|z_{\mu}|}{|z_{\mu}|^{1/p}}$. The last semi-strip occurs only if $\arg z_{\mu}=\theta$ is a Stokes ray, otherwise, we set $C=0$.
In the final, we consider the property of completely regular growth of $f$. (4.15) holds on any ray from the original point except for finitly many possible rays – all stokes rays in
$\bigcap_{j=1,2,\cdots,s}\left(\bigcup_{i=1,\cdots,s+1-j}S_{p_{1},\cdots,p_{j}}^{(i)}(\varepsilon)\right),$
and all axes of symetry in $T^{(i)}_{p_{1},\ldots,p_{j}}$ for $i=1,\ldots,s+1-j;j=1,\ldots,s$ and admissble $p_{1},\ldots,p_{s}$. Then, by Lemma 2.1, $f(z)$ is of completely regular growth on the set of rays with argument $(0,2\pi]$. $\Box$
5 A blick of the boundary of function class c.r.g.
Observing the two main Theorems 3.2 and 3.6, the solutions of completely regular growth have asymptotic expressions
$\sum_{j=0}^{n}e^{P_{j}(z)}z^{c_{j}}Q_{j}(z,\log z)$ (5.1)
in every sector except for at most finite many narrow sectors. Here $P_{j}(z)$ are at most different polynomials in $z^{1/p_{j}}$ for some integer $p_{j}$, $c_{j}$ is a complex number, and $Q_{j}$ are polynomials in $\log z$ whose coefficients have asymptotic forms
$\sum_{k=0}^{\infty}\alpha_{j,k}z^{-\frac{k}{p}}.$
The asymptotic solutions (5.1) are mainly controlled by one exponential polynomial term $\exp{P_{j}(z)}$ in any sector for some $j$. Therefore, we discover that the essence of the space of solutions of equation (3.2) does not have a big difference with its coefficients, the class of exponential polynomials. The following further result is deduced:
Theorem 5.1
Let the coefficients $a_{0},\ldots,a_{n-1}$ of the equation
$f^{(n)}+a_{n-1}f^{(n-1)}+\cdots+a_{1}f^{\prime}+a_{0}f=0$ (5.2)
be functions in the class of solutions of (3.6). All finite order trancendental solutions of (5.2) are of c.r.g..
Proof. The main idea to prove this theorem is similar to Theorem (3.6). Firstly, we know the influential factor of $p$ is the number of sums of critical rays by Lemma 2.3, so the quantity of $p$ for all solutions in angular domains is limited. Owing to it being an integer, we can find the lowest common multiple and also denote it as $p$. Then, we transform (5.2) with $t=z^{1/p}$ to an equation in $t$
${g}^{(n)}+b_{n-1}(t)g^{(n-1)}+\cdots+b_{1}(t)g^{\prime}+b_{0}(t)g=0,$ (5.3)
with $f(z)=g(t),a_{j}(z)=b_{j}(t),j=1,\ldots,n-1.$ So
$b_{j}(t)=\sum_{j=0}^{n}e^{P_{j}(t)}t^{pc_{j}}Q_{j}(t,p\log t),$
where $Q_{j}$ is a polynomial in $\log z$ over the field of formal series $\sum_{k=0}^{\infty}\alpha_{j,k}z^{-{k}}.$ Given a $\theta$, there exists a sector $S:|\arg z-\theta|<h$
$\begin{split}b_{j}(t)&=\sum_{j=0}^{n}e^{P_{j}(t)}t^{pc_{j}}\sum_{m=0}^{d}p^{m}\left(\sum_{k=0}^{\infty}\alpha_{j,k,m}t^{-k}\right)\log^{m}t\\ &=\sum_{j=0}^{n}e^{P_{j}(t)}t^{pc_{j}}\sum_{m=0}^{d}\left(\hat{Q}_{j,m}(t)\log^{m}t+o\left(\frac{{|\log t|}^{m}}{{|t|}^{N}}\right)\right),\end{split}$
where $|t|^{N}\hat{Q}_{j,m}$ are polynomials of degree $N$, for some $N\in\mathbb{N}$. Accordingly,
$b_{j}(t)=\sum_{j=0}^{n}e^{P_{j}(t)+O(\log|t|)}R_{j}(t)(1+o(1)),\quad t\rightarrow\infty.$ (5.4)
Here, $R_{j}(t)$ are rational functions with only pole 0.
Further, if we take $b_{j}$ as the coefficients in equation (3.6), we can also construct coefficient differential equations, by the method of Theorem 3.6, successively construct the fundamental coefficient differential equations in sectors. Only the last recursive process will be taken into consideration. Namely,
$H_{j_{1},j_{2},\cdots,j_{s-1}}(t_{\mu})=\sum_{j_{s}=1}^{k_{j_{1},\ldots,j_{s-1}}}e^{q_{j_{1},j_{2},\cdots,j_{s}}t+O(\log t)}\cdot\sum_{m=0}^{n}\hat{R}_{j_{1},j_{2},\cdots,j_{s},m}(t)(1+o(1))\frac{g^{(m)}}{g}=o\left(e^{ar_{\mu}}\right),$
where $a\in\mathbb{R}$, $q_{j_{1},j_{2},\cdots,j_{s}}$ are some complex numbers, $\hat{R}_{j_{1},j_{2},\cdots,j_{s},m}$ are rational functions with only pole 0, and $r_{\mu}=|t_{\mu}|$. Without generality, we simplify this equation as
$\sum_{l=1}^{K}e^{c_{l}t+O(\log|t|)}\cdot\sum_{m=0}^{n}\hat{R}_{l,j}(t)(1+o(1))\frac{g^{(j)}}{g}=o\left(e^{ar}\right),\quad r=|t|,\enspace K\in\mathbb{N},\enspace\in\mathbb{C}.$ (5.5)
Set
$\tilde{\kappa}_{l,m}(t):=\left|\frac{e^{\lambda_{l}t+O(\log|t|)}\left(\sum\limits_{j=0}^{n}\hat{R}_{l,j}\dfrac{g^{(j)}}{g}\right)}{e^{\omega_{m}t+O(\log|t|)}\left(\sum\limits_{j=0}^{n}\hat{R}_{m,j}\dfrac{g^{(j)}}{g}\right)}\right|=e^{\Re{(\lambda_{l}-\omega_{m})}t+O(\log r)}\left|\frac{\sum\limits_{j=0}^{n}\hat{R}_{l,j}\dfrac{g^{(j)}}{g}}{\sum\limits_{j=0}^{n}\hat{R}_{m,j}\dfrac{g^{(j)}}{g}}\right|.$
From lemma (3.1), there exists $\varepsilon=\arcsin\left(\left(2C\log r\right)/\left(r|\lambda_{l}-\omega_{m}|\right)\right)$ satisfies
$\Re((\lambda_{l}-\omega_{m})t)+C\log r\leq-|t|\cdot|\lambda_{l}-\omega_{m}|\cdot|\sin\varepsilon|+C\log r<0,$
for $t\in S_{m}(\varepsilon)$, every $C>0$, $r\geq r_{0}$. Thus, the remanent proof is the same with the sections in 3.2 and 3.6. The part relevant to denote the angular domains is more complicated, which is omitted here. Besides some areas near the critical rays, and parabolic strips, it is always established the fundamental coefficient differential equations are in the form of
$\gamma_{n}(t)g^{(n)}+\gamma_{n-1}(t)g^{(n-1)}+\cdots+\gamma_{0}(t)g=0,$ (5.6)
where $\gamma_{j}(t)$ has an asymptotic expansion
$\gamma_{j}(t)=t^{n_{j}}\sum_{k=0}^{\infty}\frac{1}{t^{k}},t\rightarrow\infty.$
By the similar proof of theorem 3.2 with Lemma 2.4, it follows
$\log|g|=\Re G+O(\log|t|),$ (5.7)
for every narrow sector, except for some areas constructed by the critical rays, stokes rays and discussed above. Here $G$ is a polynomial in $t^{1/p_{0}},p_{0}\in\mathbb{N^{*}}.$
Considering the indicator function of $f$ by (5.7), $t=z^{p}$, $f(z)=g(t)$ and Lemma 2.1, $f$ is of finite rational order and completely regular growth. $\Box$
We set the Class of finite order transcendental solutions of (3.6) in Theorem 3.6, and (5.2) in Theorem 5.1 as $\mathcal{EP}^{0}$, $\mathcal{EP}^{1}$ individually. On this point of view, if we take the solutions of equation (5.2) as coefficients of a linear differential again, then we will get the same conclusion the finite order transcendental solutions are of c.r.g. in $\mathcal{EP}^{2}$ and the orders are rational. As if we do this progress recursively, even if there may exist new finite order transcendental functions, but every asymptotic form in a narrow angular domain with an open angle $\varepsilon$ is just the same form as
$\log|f|=\Re P(z)+O(\log|z|)$ (5.8)
with $P(z)$ a polynomial in $z^{1/p}$ for some integer $p$. We name the class of functions, which are in $\bigcup_{k=0}^{\infty}\mathcal{EP}^{k}$, by repeatedly putting the class of solutions of former equations to the coefficients of later equations, as of exponential polynomial type. Thus, it is not hard to get this result below
Corollary 5.2
Set the coefficients $a_{0},\ldots,a_{n-1}$ of the equation
$f^{(n)}+a_{n-1}f^{(n-1)}+\cdots+a_{1}f^{\prime}+a_{0}f=0$ (5.9)
be functions of exponential polynomial type. All finite order transcendental solutions of (5.9) are of c.r.g.. In other words, functions of exponential polynomial type are of c.r.g..
From this property, it is natural to explore, what is the maximal probability of the class of functions of exponential polynomial type and if the functional class will be closed. By comparison between the function of exponential polynomial type (5.8), and the definition of c.r.g. in an angular domain in the form of
$\log|f|=h(\theta)r^{\rho}+o(r^{\rho}),$
we find that there may exist commonality. When $\rho$ is a rational number, their difference is just the small quantity of $r^{\rho}$, referring to the equality
$h(\theta)=A(\theta)\cos(\rho\theta+\varphi),$
for $z=re^{i\theta}$ in a narrow angular domain $S$, where the density of zeros in $S$ is zero, and $A,\varphi\in\mathbb{R}$ in the book of Levin[9, p.115]. This property is also very natural for function of exponential polynomial type in $S$ on the rays where zeros not mainly distributed, from the conclusion that the zeros here are just $O(\log r)$ resulting from Steinmetz [12, Cor. 1].
On the other, it leaves some space for us to reflect on:
  1. 1.
    How large is the gap between the two classes of functions, of rational order with completely regular growth, and the one of exponential polynomial type?
  2. 2.
    Rational growth order can not be the boundary condition of Gol’dberg-Ostrovskiǐ’s Problem. In other words, we can, under the combination of functions of exponential polynomial type and Bergweiler’s method, take $H$ in the proof of [1, Theorem 1.1] as any function of exponential polynomial type and rational growth, so that $A$ and $B$ of the linear differential equation
    $f^{\prime\prime}+Af^{\prime}+Bf=0$
    are of c.r.g. and rational order. However, the order and lower order of the constructive solution $f$ are not equal. Does there exist one class of functions of irrational orders, which make the both, coefficients and solutions of c.r.g.?
Acknowledgements
Thanks to my adviser Zhi-Tao Wen for providing some useful materials and Prof. Walter Bergweiler for giving pieces of advice during my exchange studies at Christian-Albrechts University of Kiel.