### v2

# On the Goldberg-Ostrovskii Problem for Linear Differential Equations with Exponential Polynomial Coefficients

Abstract
The Goldberg–Ostrovskii problem asks whether finite-order solutions of a linear differential equation inherit the property of completely regular growth (c.r.g.) from its coefficients. While Bergweiler’s counterexample demonstrated that the answer is negative in general, this paper proves that when the coefficients are restricted to the classical and rich subclass of exponential polynomials, the regularity transmission does hold. Thereby we affirm the conjecture posed by Heittokangas, Ishizaki, Tohge and Wen.
Our results reveal the closed nature of exponential polynomials in the context of regularity transfer from the perspective of equation dynamics, and provide a new perspective for the study of the structure of related function classes.
Keywords: Exponential polynomials; Completely regular growth; Goldberg-Ostrovskii’s problem; Asymptotic analysis
2020 MSC: 30D15, 34M10, 30D35
1 Introduction
Completely regular growth and classic problems on equations
The theory of linear differential equations in the complex domain represents one of the most profound intersections of complex analysis, operator theory, and dynamical systems. At the heart of this theory lies the fundamental problem of understanding the growth and distribution of zeros of entire function solutions. The Ukrainian mathematician B. Ja. Levin revealed a crucial connection through the concept of completely regular growth (c.r.g.), which establishes a deep relationship between growth control and zero distribution on sectors.
The Phragmén-Lindelöf indicator function of an entire function $f(z)$ of finite order $\rho(f)>0$ is given by
$h_{f}(\theta)=\limsup_{r\to\infty}r^{-\rho(f)}\log|f(re^{i\theta})|$
for $\theta\in[0,2\pi)$. We say that $f$ is of completely regular growth (briefly denoted by c.r.g., see [23, pp. 139-140] or [32, pp. 6-8]) if there exists a sequence of Euclidean discs $D(z_{k},r_{k})$ satisfying
$\sum_{|z_{k}|\leq r}r_{k}=o(r)$ (1.1)
such that
$\log|f(re^{i\theta})|=(h_{f}(\theta)+o(1))r^{\rho(f)},\quad re^{i\theta}\not\in\bigcup_{k}D(z_{k},r_{k}),$ (1.2)
as $r\to\infty$ uniformly in $\theta$. A function of c.r.g. is equivalent to the regular distribution of its zeros, see [23, p. 158].
A. A. Goldberg and I. V. Ostrovskii stated a classic problem in the book [15, p. 300]:
(Goldberg-Ostrovskii) Suppose that $f$ is a finite order transcendental solution of f^(n)+a_n-1f^(n-1)++a_1f’+a_0f=0, whose coefficients are entire functions of c.r.g.. Is $f$ also of c.r.g.?
This problem was formulated by V.P. Petrenko in [29, p. 132] but he omitted the condition of c.r.g. of the coefficients $a_{j}(z)$, for $1\leq j\leq n-1$. As the book was published posthumously without the author’s final proofreading, this omission may have been accidental. Crucially, without this c.r.g. condition, the answer to the problem is negative, see [15, p. 300]. Bergweiler [4] constructed a counterexample showing that additional conditions are necessary, thus motivating the search for the precise conditions under which regularity is inherited. This makes us wonder, what should the “missed condition” on coefficients be that makes all the finite order solutions of c.r.g.?
We wish to recognize just what functions are of c.r.g., which is an important consideration on recapturing this problem. Detailed studies of functions of c.r.g. can be found in [23] and [32]. Obviously, non-transcendental functions can not be functions of c.r.g. and $e^{z}$ should be the simplest one without zeros.
The combination of exponential terms and polynomials, such as $1-z+e^{z}$, can be also easily verified from the condition (1.2). Mathematicians such as Polyà [30], Schwengeler [33] and Dickson [6] have found that the functions – exponential sums, see (3.1), has a distribution of zeros that asymptotic to some rays, i.e., just as $1-z+e^{z}$ does. Although they give different estimates of the asymptotic trend, the regularly distributed zero property tells us such functions are equivalently of c.r.g..
A more general function class comparing exponential sums is known to us below. An exponential polynomial is an entire function of the form
$G=P_{1}e^{Q_{1}}+\cdots+P_{l}e^{Q_{l}}$ (1.3)
where $P_{j},Q_{j},j=1,\ldots,l,$ are polynomials, $l\in\mathbb{N}$. It is known that exponential sums are exponential polynomials with all $Q_{j}$ of degree 1. Transcendental exponential polynomial functions are also of c.r.g., see [17, Lemma 1.3] for example.
For the more complicated functions of c.r.g., it was not possible to write the explicit forms, but it was found that the exponential terms played a crucial role. To explain it clearly, we know different exponential terms control the growth of an exponential sum in different sectors. Such as $e^{-z}$ controls the growth of $e^{-z}+e^{z}$ on the left-half plane and $e^{z}$ controls it on the right-half plane. In the area around the positive and negative imaginary axes, there will be some sudden changes in the function values to produce zeros, such a phenomenon is known as the Stokes phenomenon, see Wasow [38, Sec. 15].
Although the asymptotic analysis theory on Stokes pheonomenon does not only deals with entire functions, such as also the example – Hankel functions – which are the linearly independent solutions of Bessel’s differential equation introduced in [38, p. 78], the well-known Airy functions $\mathrm{Ai}(z)$ and $\mathrm{Bi}(z)$, as the solutions of the Airy equation
$f^{\prime\prime}-zf=0,$
has the typical Stokes phenomenon on the negative axis, see [27, Chap. 9]. Petrenko [29, pp. 104-112] and Steinmetz [34] independently proved that the entire solutions of linear differential equations with polynomial coefficients (Steinmetz indicated this result also holds for rational coefficients) are of c.r.g.. More such entire special functions of c.r.g., e.g. generalized Airy functions $A_{n}(z)$ and $B_{n}(z)(n\in\mathbb{Z})$, Bessel functions of the first kind $J_{n}(z)(n\in\mathbb{Z})$, and Kummer functions $M(a,b,z)(a,b\in\mathbb{C})$, can be seen separately in [27, §9.13, §10.2, §13.2].
The fact that so many c.r.g. special functions (or associated forms) fit into the framework of exponential polynomials indicates that this class is a natural candidate for the study of regularity transmission—being both broad yet well-behaved. Any exponential polynomial can still be constructed to be a solution to a linear differential equation whose coefficients are polynomials, see [37]. Heittokangas et al. are expert in exponential polynomials and they indicated the regularity that for certain LDE the class of finite order solutions is always larger (at least not smaller) than the class of coefficients, see the survey [19] on exponential polynomials. Therefore, in [19, p. 33] they proposed a relevant question by strengthening the assumption on G.-O.’s problem.
(H.-I.-T.-W.) If the coefficients $a_{0}(z),\ldots,a_{n-1}(z)$ of
$f^{(n)}+a_{n-1}f^{(n-1)}+\cdots+a_{1}f^{\prime}+a_{0}f=0,$ (1.4)
are exponential polynomials, and if it possesses a transcendental entire solution $f$ of finite order of growth, then is it true that $f$ is of c.r.g.?
Our recent work in [24] provides a partial answer to this line of inquiry, showing that for linear differential equations whose coefficients are exponential sums (a subclass of exponential polynomials) with rational frequencies, all finite-order solutions are themselves exponential sums and consequently of completely regular growth.
This paper is to give a complete answer to this problem. We give an estimation of $\log|f|$ on every small sector except for some areas around a finite number of rays from the origin, resulting $f$ of c.r.g.. See Theorem 3.2, Theorem 3.5, and Theorem 3.9. In summary, this paper completely proves that every finite-order transcendental entire solution of a linear differential equation with exponential polynomial coefficients is of completely regular growth. This conclusively settles the Heittokangas–Ishizaki–Tohge–Wen conjecture.
Organization
Section 2 recalls necessary concepts: functions of c.r.g. on rays and its compactification, the asymptotic theory of ODEs in sectors by Wasow, and Steinmetz’s application to LDEs. Section 3 presents our main results for equations with exponential sum coefficients (Theorem 3.2, Theorem 3.5) and with exponential polynomial coefficients (Theorem 3.9), including a description of the asymptotic sectors and exceptional sets. Section 4 contains the complete proofs, detailing the reduction to polynomial-coefficient equations in sectors, the analysis near critical and Stokes rays, and the final verification of the c.r.g. property via the compactness principle. Section 5 offers concluding remarks and outlines future directions, including the exploration of the solution space structure via operator-theoretic methods. Section 6 (Appendix) provides a necessary modification of Wasow’s sectorial asymptotic theorem to regions with holes, which underpins the proof when dealing with exceptional sets.
2 Basic concepts and lemmas
2.1 Compactification for functions of c.r.g. on the rays
We have given the definition of functions of completely regular growth above. In fact, completely regular growth is not only a global property but a local one. A set which can be covered by a sequence of discs $D(z_{k},r_{k})$ satisfying (1.1) will be called a $C_{0}$-set. The relative measure of a set $E$ of positive numbers is defined to be the limit
$m^{*}(E)=\lim\limits_{r\rightarrow\infty}\frac{mes(E_{r})}{r},$
[…]
In the definition of c.r.g., the exception condition $z\notin C_{0}$ and $|z|\notin E_{0}$ is equivalent by the reason of the correspondence of regular distribution of zeros, see [23, Thm. 1&2, Chap. 1] and [23, Thm.4, p. 158], or [32, Thm. 1.2.1].
Refering to the definitions in [23, p. 139], a function $f$ that is holomorphic in $\{re^{i\theta}:\theta_{1}<\theta<\theta_{2},\theta_{1},\theta_{2}\in\mathbb{R},r>0\}$ and of order $\rho$ will be called a function of completely regular growth on the ray $\arg z=\theta$ for some $\theta\in(\theta_{1},\theta_{2})$ if the limit
$h_{f}(\theta)=\displaystyle\lim_{r\to\infty\atop r\notin E_{\theta}}\frac{\log|f(re^{i\theta})|}{r^{\rho}}$
exists, where $E_{\theta}$ is a $E_{0}$-set depending on $\theta$. Denote $R_{\theta}=\{re^{i\theta},r\notin E_{\theta},r>0\}$ for some angle $\theta$, a subset $\mathbb{M}\subset(\theta_{1},\theta_{2})$ and
$R_{\mathbb{M}}=\{re^{i\theta}:r>0,r\notin\displaystyle\bigcup_{\theta\in\mathbb{M}\subset(\theta_{1},\theta_{2})}E_{\theta}\}\subset\displaystyle\bigcup_{\theta\in\mathbb{M}\subset(\theta_{1},\theta_{2})}R_{\theta}.$
We will say a function of completely regular growth on the set of rays $R_{\mathbb{M}}$ if
$H_{f}(r,\theta)=\frac{\log|f(re^{i\theta})|}{r^{\rho}},~r\notin\displaystyle\bigcup_{\theta\in\mathbb{M}}E_{\theta},$
converges uniformly to $h_{f}(\theta)$ for all $\theta,\theta\in\mathbb{M}.$ Additionally, Levin [23, p. 140] has proved that $f(z)$ is a function of completely regular growth on every ray which is a limiting ray of the set of rays along which $f(z)$ is of completely regular growth.
Lemma 2.1
If a holomorphic function of growth order $\rho$ is of completely regular growth on each ray $R_{\theta}$ for $\theta\in\mathbb{M}$, then it is of completely regular growth on the set of rays $R_{\overline{\mathbb{M}}}$.
It follows from this, in particular, that an entire function of completely regular growth on rays that form an everywhere dense set is a function of completely regular growth on the whole plane.
2.2 The asymptotic solutions of differential equations in a sector
Considering the general differential equation
$f^{(n)}(z)+a_{n-1}(z)f^{(n-1)}(z)+\cdots+a_{1}(z)f^{\prime}(z)+a_{0}(z)f(z)=0$ (2.1)
with entire coefficients $a_{0}(z),\ldots,a_{n-1}(z)$, Petrenko [29, pp. 104–112] has shown that transcendental solutions to linear differential equations (2.1) with polynomial coefficients are of completely regular growth. Steinmetz [34, Theorem 1] has proved a more generalized theorem in the following. The differential equation
$f^{(n)}(z)+R_{n-1}(z)f^{(n-1)}(z)+\cdots+R_{1}(z)f^{\prime}(z)+R_{0}(z)f(z)=0$ (2.2)
allows rational coefficients $R_{0}(z),\ldots,R_{n-1}(z)$. And it is only considered the solutions of (2.2) are meromorphic functions.
Lemma 2.2 (Steinmetz)
Let $f(z)$ be a non-rational meromorphic solution of (2.2) and let $\arg z=\theta$ be an arbitrary direction. Then there is $p\in\mathbb{N}$, $h>0$, and a polynomial $Q$ in $z^{1/p}$ such that
$\log|f(z)|=\Re Q(z^{1/p})+O(\log|z|)$ (2.3)
as $z\rightarrow\infty$ in $\theta\leq\arg z\leq\theta+h,$ possibly outside an exceptional set consisting of
  1. 1.
    (countably many) disks $|z-z_{\mu}|<|z_{\mu}|^{1-\varepsilon},$ where $\varepsilon$ is positive and the counting function of the sequence $z_{\mu}$} is $O(\log r)$;
  2. 2.
    a logarithmic semi-strip $0\leq\arg z-\theta<\dfrac{\log^{+}|z|}{|z|^{\frac{1}{p}}}$.
The latter occurs only for some finite numbers of rays.
This kind of exceptional rays are called Stokes rays of $f(z).$ Concerning the definition of Stokes ray, one can refer to Steinmetz [34], and the general definition in Wasow [38, Sec. 15]. In proving our results, our goal is not to assume anymore that the coefficients of the equation (2.2) are rational functions, but to assume that the coefficients have an asymptotic expansion at $\infty$ in a sector $S$. From this point of view, it is necessary to define the asymptotic symbol $\sim$. We write
$f(z)\sim C_{0}+\frac{C_{1}}{z}+\frac{C_{2}}{z^{2}}+\cdots,$
when
$\lim_{z\to\infty}\left(f(z)-C_{0}-\frac{C_{1}}{z}-\frac{C_{2}}{z^{2}}-\cdots-\frac{C_{m}}{z^{m}}\right)z^{m}=0,$
for every positive integer $m\geq 0$, or using the symbol $O(\cdot)$ to denote,
$f(z)=C_{0}+\frac{C_{1}}{z}+\frac{C_{2}}{z^{2}}+\cdots+\frac{C_{m}}{z^{m}}+O\left(\frac{1}{z^{m+1}}\right).$
A linear system which has a pole at the origin written in the matrix expression is
$x^{h}\frac{d\bm{Y}}{dx}=\bm{B}(x)\bm{Y},\kern 5.0pth\in\mathbb{N},$ (2.4)
where $\bm{Y},\bm{B}$ are $n$-by-$n$ matrixes, $n\in\mathbb{N^{*}}$, and the matrix $\bm{B}(x)$ is holomorphic at $x=0$. When $h>1,$ the singular point is irregular. We want to make a transformation $z=1/x$ taking into the equation (2.4), affording
$z^{-d}\frac{d\bm{Y}}{dz}=\bm{A}(z)\bm{Y},$ (2.5)
with d=h2d=h-2 and $\bm{A}(z)=-\bm{B}(1/x)$. The integer $d+1$ is called the rank of the singular point. Then for $n$-by-$n$ constant matrixes $\bm{A}_{r}$, the matrix $\bm{A}(z)$ could have its expansion with form
$\bm{A}(z)=\sum_{j=0}^{\infty}\bm{A}_{j}z^{-j},\quad|z|>z_{0},$ (2.6)
which is holomorphic at $z=\infty$, or an expansion of this form asymptotic to $\bm{A}(z)$ in some sector $S$, i.e.
$\bm{A}(z)\sim\sum_{j=0}^{\infty}\bm{A}_{j}z^{-j},\kern 5.0ptz\in S.$ (2.7)
We pay more interest to the latter case (2.7) and wonder if there exists an asymptotic expression of the solution $\bm{Y}$ in the same area. The matrix solution in the sector $S$ is summarized in the following theorem, see [38, p.111]:
Lemma 2.3
Let $\bm{A}(z)$ be an $n$-by-$n$ matrix function, which is holomorphic and has the asymptotic form (2.7) in $S$. Then, the differential equation (2.5) possesses a fundamental matrix solution of the form
$\bm{Y}(z)=\hat{\bm{Y}}(z)z^{\bm{G}}e^{\bm{Q}(z)}$ (2.8)
corresponding to every sufficiently narrow subsector of $S$. Here $\hat{\bm{Y}}(z)$ permits an asymptotic series in power of $z^{-1/p},p\in\mathbb{N^{*}}$, in this subsector as $z\rightarrow\infty;$ $\bm{G}$ is a constant matrix, $\bm{Q}(z)$ is a diagonal matrix whose diagonal elements are polynomials in $z^{1/p}$.
Suppose a differential equation
$f^{(n)}(z)+A_{n-1}(z)f^{(n-1)}(z)+\cdots+A_{1}(z)f^{\prime}(z)+A_{0}(z)f(z)=0$ (2.9)
possesses coefficients $A_{k}(z)$ with the asymptotic expansions
$A_{k}(z)\sim z^{d_{k}}\sum_{j=0}^{\infty}A_{kj}z^{-j},\kern 5.0ptz\in S,A_{kj}\in\mathbb{C},d_{k}\in\mathbb{N},k=0,\cdots,n-1.$
If we transform the (2.9) into a linear system, setting $\bm{y}=\{f,\ldots,f^{(n-1)}\}^{T},$ where $T$ represents the transpose of matrix, and $\tilde{\bm{A}}$ as the companion matrix
$\tilde{\bm{A}}=\begin{pmatrix}0&1&&&\\ &0&1&&\\ &&\ddots&\ddots&\\ &&&0&1\\ -A_{0}&-A_{1}&\cdots&-A_{n-2}&-A_{n-1}\end{pmatrix}$,
it follows that (2.9) satifies the linear system
$z^{-d}\frac{d\bm{y}}{dz}=\bm{A}(z)\bm{y},\kern 5.0ptz\in S,$
with $z^{d}\bm{A}=\tilde{\bm{A}}$ and integer $d=\max\limits_{k=0,\ldots,n-1}\{d_{k}\}$. Therefore, we transform equation (2.9) to a linear system (2.5) with $\bm{A}$ in the form of (2.7) after set $\bm{Y}=\{\bm{y}_{1},\ldots,\bm{y}_{n}\}$, $\bm{y}_{i}=\{f_{i},\ldots,f^{(n-1)}_{i}\}^{T},$ where $f_{i}$ are $n$ linearly independent solutions of (2.9), $i=1,\ldots,n$. The elements of $\bm{A}_{j}$ in (2.7) consisted of some of the $A_{kj},0,1$, and $k=0,\ldots,n-1,j\in\mathbb{N}.$
By Lemma 2.3 and fundamental calculations of elements of Matrixes, the $n$ linear independent solutions $f_{i}$ in $z$ in $S$ are
$f_{i}(z)=e^{\varLambda_{i}(z^{\frac{1}{p}})}z^{c_{i}}\varOmega_{i}(z,\log z),\quad 1\leq i\leq n.$ (2.10)
Here $\varLambda_{i}$ is a polynomial in $z^{1/p}$, $p\in\mathbb{N^{*}}$, $c_{i}\in\mathbb{C}$, and $\varOmega_{i}$ is a polynomial in $\log z$ with coefficients $\beta_{i,j}(z)$ possessing asymptotic forms
$\beta_{i,j}(z)\sim\sum_{s=0}^{\infty}\alpha_{i,j,s}z^{-s/p},\quad\alpha_{s}\in\mathbb{C},1\leq i\leq n,0\leq j\leq n-1.$
Repeating the proof of Lemma 2.2 given by Steinmetz [34], we get the following lemma:
Lemma 2.4
Let $f(z)$ be any solution of (2.9) in the sector $S$ and let $\arg z=\theta$ be an arbitrary direction in the interior of $S$. Then there is $p\in\mathbb{N},h>0,$ a polynomial $Q_{\theta}$ in $z^{1/p}$, such that
$\log|f(z)|=\Re Q_{\theta}(z^{1/p})+O(\log|z|)$ (2.11)
as $z\rightarrow\infty$ in $\theta\leq\arg z\leq\theta+h$ possibly outside an exceptional set consisting of
  1. 1.
    (countably many) disks $|z-z_{\mu}|<|z_{\mu}|^{1-\varepsilon},$ where $\varepsilon$ is positive and the counting function of the sequence $z_{\mu}$} is $O(\log r)$;
  2. 2.
    a logarithmic semi-strip $0\leq\arg z-\theta<\dfrac{\log^{+}|z|}{|z|^{\frac{1}{p}}}$.
The latter occurs only if $\arg z=\theta$ is a Stokes ray in $S$.
3 Main results on solutions of c.r.g.
3.1 LDE with exponential sums coefficients
An exponential sum is an entire function of the form
$F_{1}(z)e^{\lambda_{1}z}+\cdots+F_{m}(z)e^{\lambda_{m}z},$ (3.1)
where $\lambda_{1},\ldots,\lambda_{m}\in\mathbb{C}$ are distinct constants called leading coefficients (or frequencies) of the exponential sum (3.1), and the coefficients $F_{1}(z),\ldots,F_{m}(z)$ are polynomials.
Firstly, we consider the linear differential equation
$f^{(n)}(z)+A_{n-1}(z)f^{(n-1)}(z)+\cdots+A_{1}(z)f^{\prime}(z)+A_{0}(z)f(z)=0$ (3.2)
with exponential sums coefficients $A_{i}(z)$ for $i=0,1,\ldots,n-1$ of the form (3.1). The equation (3.2) can be transformed into a normal form
$\sum_{j=1}^{N_{0}}e^{\lambda_{j}z}\left(a_{n,j}(z)f^{(n)}(z)+a_{n-1,j}(z)f^{(n-1)}(z)+\cdots+a_{0,j}(z)f(z)\right)=0,$ (3.3)
where $a_{n,j}(z)\in\{0,1\}$, $a_{n-1,j}(z),\ldots,a_{0,j}(z)$ are polynomials, and $\{\lambda_{j}\}$ is the set of all the mutually different leading coefficients of $A_{0}(z),\ldots,A_{n-1}(z)$.
$a_{n,j}(z)f^{(n)}(z)+a_{n-1,j}(z)f^{(n-1)}(z)+\cdots+a_{0,j}(z)f(z)=0$
are called coefficient differential equations about $\lambda_{j}$ (or at $\overline{\lambda}_{j}$) individually for $j=1,\ldots,N_{0}$. Set $W=\left\{\overline{\lambda}_{j}\right\}$, the convex hull $co(W)$ of $W$ is defined as the intersection of all closed convex sets containing $W$, also to be said as the convex hull to the differential equation (3.2). Suppose the vertexes of $co(W)$ consist a set $\tilde{W}=\{\overline{\omega}_{j}\}(j=1,\ldots,N_{0}^{\prime})$ with elements arranged anticlockwise, $N_{0}^{\prime}\leq N_{0}$, and denote $s_{j}$ the segment between $\overline{\omega}_{j}$ and $\overline{\omega}_{j+1}$. The critical ray $\arg z=\eta_{j}$ is defined as the ray originated at 0 with the direction of the outer normal to $s_{j}$. For small enough $\varepsilon>0,$ we define
$S_{j}(\varepsilon)=\{z|\eta_{j-1}+\varepsilon\leq\arg z\leq\eta_{j}-\varepsilon\},\kern 5.0ptT_{j}(\varepsilon)=\{z|\eta_{j}-\varepsilon<\arg z<\eta_{j}+\varepsilon\}$
and
$\overline{S}_{j}=\{z|\eta_{j-1}\leq\arg z\leq\eta_{j}\}.$
See Figure 1.
[…]
Figure 1: $\tilde{W},\arg z=\eta_{j},S_{j}(\varepsilon),T_{j}(\varepsilon),\overline{S}_{j}$
Lemma 3.1
If $\overline{\lambda}\in{W},$ then for every $z\in S_{j}(\varepsilon)$,
$\Re((\omega_{j}-\lambda)z)\geq|z|\cdot|\omega_{j}-\lambda|\cdot|\sin\varepsilon|.$
If a given exponential point $\overline{\lambda}_{k}\notin s_{j}$, then for a given number $\psi_{\varepsilon}>0$ with $\psi_{\varepsilon}\neq 0$, such that for every $z\in\overline{S}_{j}\cap T_{j}(\theta)$ the inequality
$\Re((\omega_{j}-\lambda_{k})z)\geq|z|\cdot|\omega_{j}-\lambda_{k}|\cdot|\sin\psi_{\varepsilon}|$
is valid.
If a given exponential point $\overline{\lambda}_{k}\notin s_{j}$, then for a given number $\psi_{\varepsilon}>0$ with $\psi_{\varepsilon}\neq 0$, such that for every $z\in\overline{S}_{j+1}\cap T_{j}(\theta)$ the inequality
$\Re((\omega_{j}-\lambda_{k})z)\geq|z|\cdot|\omega_{j+1}-\lambda_{k}|\cdot|\sin\psi_{\varepsilon}|$
is valid.
See Dickson [6, Lemma, 1].
Our conclusions are as follows.
Theorem 3.2
Every finite order trancendental entire solution $f$ of (3.2) satisfies
$\log|f(z)|=\Re G_{j\theta}(z^{1/p_{j}})+O(\log|z|),$
for some polynomial $G_{j\theta}$ in $z^{1/p_{j}}$, $p_{j}\in\mathbb{N^{*}},j=1,\ldots,N_{0}^{\prime}$, outside a $r$-value set $E$ of finite linear measure, in $\theta\leq\arg z\leq\theta+h,\theta\in(0,2\pi]$, sufficiently small $h>0$, and enough large $|z|$, besides two kinds of areas:
  1. 1.
    $z$ in $T_{j}(\varepsilon)$, $\varepsilon>0$;
  2. 2.
    the logarithmic semi-strips
    $0\leq\arg z-\xi_{jk}<\frac{\log^{+}|z|}{{|z|}^{\frac{1}{p_{j}}}},k,p\in\mathbb{N^{*}}$ (3.4)
    in $S_{j}(\varepsilon)$, where $\arg z=\xi_{jk}$ are the Stokes rays to the coefficient differential equations at $\overline{\omega}_{j}$ in $S_{j}$,
$j=1,\ldots,N_{0}^{\prime}$ with $N_{0}^{\prime}$ the number of sides to the convex hull $co(W)$.
Note.
  1. 1.
    The polynomial $G_{j,\theta}$ will change according to the rotation of $\arg z=\theta$ from one side of $S_{j}$ to the other inside $S_{j}$, but this change occurs only when $\arg z=\theta$ is a Stokes ray in $S_{j},j=1,\ldots,N_{0}^{\prime}$.
  2. 2.
    Steinmetz [34] indicated us that the exceptional set in the form of logarithmic semi-strips does not affect c.r.g., since this strip does not contain any ray except the boundary ray, the Stokes ray. Namely, if there exists a $\theta_{0}>0$, and outside the strip (3.4) $f$ satisfies c.r.g. on the set of rays
    $R_{\mathbb{M}}=\{re^{i\theta}:r>0,r\notin\displaystyle\bigcup_{\theta\in\mathbb{M}\subset(\xi_{jk},\xi_{jk}+\theta_{0})}E_{\theta}\}\subset\displaystyle\bigcup_{\theta\in\mathbb{M}\subset(\xi_{jk},\xi_{jk}+\theta_{0})}R_{\theta},$
    where $R_{\theta}=\{re^{i\theta},r\notin E_{\theta},r>0\}$, $E_{\theta}$ is a $E_{0}$-set depending on $\theta$ (see the definitions in Section 2.1), then $f$ is still of c.r.g. on any ray $\theta\in\mathbb{M}\subset\left(\xi_{jk},\xi_{jk}+\theta_{0}\right)$. Then, by the compactification of $\left(\xi_{jk},\xi_{jk}+\theta_{0}\right)$, Lemma 2.1, $f$ is of c.r.g. on $R_{[\xi_{jk},\xi_{jk}+\theta_{0}]}$.
  3. 3.
    To obtain an estimate of $\log|f|$ on $T_{j}(\varepsilon)$, we need to perform a more detailed characterization of the convex hull $co(W)$ of $W=\{\overline{\lambda}_{j}\}$ to (3.3). Since $T_{j}(\varepsilon)$ is a narrow sector between $S_{j}(\varepsilon)$ and $S_{j+1}(\varepsilon)$, and each $S_{j}(\varepsilon)$ is dominated by the vertex $\overline{\omega}_{j}$, it is natural to consider whether $\log|f|$ in $T_{j}(\varepsilon)$ is dominated by both vertices $\overline{\omega}_{j}$ and $\overline{\omega}_{j+1}.$ Here we brief introduce, $\log|f|$ in $T_{j}(\varepsilon)$ might be dominated by all the elements of $W$ on the side $s_{j}$ between $\overline{\omega}_{j}$ and $\overline{\omega}_{j+1}$.
In fact, [7, Lem. 8] inspires us the degeneration rules of the coefficients $A_{m}(z)$, $m=0,\ldots,n-1$, to (3.2) in $T_{j}(\varepsilon)$. From the following conclusion, it can be seen that only the conjugate indicators which are on the boundary $s_{j}$ control the growth of these coefficients $A_{m}(z)$ in $T_{j}(\varepsilon)$.
Corollary 3.3
The normal form (3.3) of (3.2) in $T_{j}(\varepsilon)$ has an asymptotic form of
$\sum_{\overline{\lambda}_{k}\in s_{j}}e^{\lambda_{k}z}\sum_{m=0}^{n}c_{m,k}z^{d_{m,k}}(1+\varepsilon(z))f^{(m)}(z)=0,$ (3.5)
where $c_{m,k}z^{d_{m,k}}$ is the highest-order term of the polynomial $a_{m,k}(z)$ in (3.3). The side $s_{j}$ of $co(W)$ has at least two boundary points $\overline{\omega}_{j}$ and $\overline{\omega}_{j+1}.$
Consider $\overline{\lambda}_{1},\ldots,\overline{\lambda}_{K}\in s_{j}$ with $j$ fixed. Let $\tau_{k,m}=\overline{\lambda}_{k}+id_{m,k}e^{i\eta_{j}}$, where $d_{m,k}$ is the degree of $a_{m,k}(z)$, $\arg z=\eta_{j}$ is the critical ray inside $T_{j}(\varepsilon)$. Let $\Omega_{j}$ be the smallest closed convex set containing $\overline{\omega}_{j},\overline{\omega}_{j+1}$ and all $\tau_{k,m}$. Then, without generality, we assume the vertices of $\Omega_{j}$ is anticloskwise set by $\tilde{\omega}_{1},\ldots,\tilde{\omega}_{K^{\prime}}$, $\overline{\omega}_{j+1},$ and $\overline{\omega}_{j}$, where $\tilde{\omega}_{k}=\overline{\lambda}_{k}+i\max_{m}\{d_{m,k}\}e^{i\eta_{j}}$, $\overline{\omega}_{j}=\overline{\lambda}_{1}$, $\overline{\omega}_{j+1}=\overline{\lambda}_{K^{\prime}}$, and $2\leq K^{\prime}\leq K.$ Set the slope of the segment $\tilde{s}_{k}=[\tilde{\omega}_{k},\tilde{\omega}_{k+1}]$ as
$\mu_{k}=\frac{\max_{m}\{d_{m,k}\}-\max_{m}\{d_{m,k+1}\}}{|\overline{\lambda}_{k}-\overline{\lambda}_{k+1}|},$
and two strips
$\begin{split}U_{1}=&\{z|\Im(ze^{-i\eta_{j}})\geq 0,\Re(ze^{-i\eta_{j}})+\mu_{1}\log|z|\geq\delta\},\\ U_{K^{\prime}}=&\{z|\Im(ze^{-i\eta_{j}})\geq 0,\Re(ze^{-i\eta_{j}})+\mu_{K^{\prime}}\log|z|\leq-\delta\},\end{split}$
for $\delta>0$. Since when $z\to\infty$, $\mathbb{C}\backslash(U_{1}\cup U_{K^{\prime}})$ cannot cover any sector or ray (except for $\arg z=\eta_{j}$), it is sufficient for us to consider the growth of functions in $T_{j}(\varepsilon)\cap U_{1}$ and $T_{j}(\varepsilon)\cap U_{K^{\prime}}$ to get c.r.g. in $T_{j}(\varepsilon)$, outside an exceptional set that does not contain any sector. For example, we consider a trivial case – if $K^{\prime}=2$ and $\mu_{1}=0$ (More specifically, all $d_{m,k}$ are zero for $\overline{\lambda}_{k}\in s_{j}$), then $T_{j}(\varepsilon)\backslash(U_{1}\cup U_{2})$ is a strip parallel to the critical ray $\arg z=\eta_{j}$ and having it as its axis of symmetry, with a width of $2H$. A typical conclusion is given by [6, Lemma 10], applied to the result below.
Corollary 3.4
Given the symbols as defined above. Let $z\in U_{1}.$ If $\tau_{p,m}\notin\tilde{s}_{1}$, there exists some $v>0$, then
$|z^{d_{m,p}}e^{\lambda_{p}z}|<|z|^{-v}\cdot\left|z^{\max_{m}\{d_{m,1}\}}e^{\omega_{j}z}\right|.$
If $\tau_{p,m}\in\tilde{s}_{1}$, then
$|z^{d_{m,p}}e^{\lambda_{p}z}|\leq|z^{\max_{m}\{d_{m,1}\}}e^{\omega_{j}z}|\cdot\exp{(-\delta|\lambda_{p}-\omega_{j}|)}.$
It is similar to the case $z\in U_{K^{\prime}}$ by substituting $d_{m,1}$, $\omega_{j}$, and $\tilde{s}_{1}$ into $d_{m,K^{\prime}}$, $\omega_{j+1}$, and $\tilde{s}_{K^{\prime}-1}$ individually.
By Corollary 3.3 and 3.4, we can deal with the case $z\in T_{j}(\varepsilon)$ in Theorem 3.2.
Theorem 3.5
All finite order transcendental entire solutions of (3.2) are of c.r.g..
Example 3.6
The function $f(z)=e^{-\frac{4}{3}z}(1-7e^{z})$ solves the differential equation
$f^{\prime\prime\prime}+3e^{z}f^{\prime\prime}+\left(-\frac{4}{3}-2e^{z}\right)f^{\prime}-\left(e^{z}-\frac{16}{27}\right)f=0.$
The equation can be transformed into its normal form
$\left(f^{\prime\prime\prime}-\frac{4}{3}f^{\prime}+\frac{16}{27}f\right)+e^{z}\left(3f^{\prime\prime}-2f^{\prime}-f\right)=0.$
Consider the differential polynomials in $g$,
$G_{1}(g)=\frac{g^{\prime\prime\prime}}{g}-\frac{4}{3}\frac{g^{\prime}}{g}+\frac{16}{27},\kern 5.0pt\text{G}_{2}(g)=3\frac{g^{\prime\prime}}{g}-2\frac{g^{\prime}}{g}-1,$
and the equations $G_{1}=0,G_{2}=0.$ When we take $f(z)=e^{-\frac{4}{3}z}(1-7e^{z})$ into $G_{1}$ and $G_{2}$ individually, we find
$G_{1}(f(z))=1-\frac{1}{1-7e^{z}}\rightarrow 0,\kern 5.0ptz\in S_{0}(\varepsilon)=\{z|\pi/2+\varepsilon\leq\arg z\leq 3\pi/2-\varepsilon\},$
and
$G_{2}(f(z))=\frac{7}{1-7e^{z}}\rightarrow 0,\kern 5.0ptz\in S_{1}(\varepsilon)=\{z|-\pi/2+\varepsilon\leq\arg z\leq\pi/2-\varepsilon\}.$
They are both $o\left(\frac{1}{z^{N}}\right)$ for $N\in\mathbb{N}$.
Example 3.7
The differential equation
$4ze^{iz}f^{\prime\prime\prime}+(6e^{iz}+4ize^{iz}-z)f^{\prime\prime}+\left((2i-1)e^{iz}-\frac{1}{2}\right)f^{\prime}-\left(ie^{iz}-\frac{1}{4}\right)f=0$ (3.6)
is given. The normal form of (3.7) is
$e^{iz}\{4zf^{\prime\prime\prime}+(6+4iz)f^{\prime\prime}+(2i-1)f^{\prime}-if\}+\left(-zf^{\prime\prime}-\frac{1}{2}f^{\prime}+\frac{1}{4}f\right)=0.$
The solution $f=\cosh\sqrt{z}$ of (3.6) solves both equations
$4zf^{\prime\prime\prime}+(6+4iz)f^{\prime\prime}+(2i-1)f^{\prime}-if=0,\quad-zf^{\prime\prime}-\frac{1}{2}f^{\prime}+\frac{1}{4}f=0.$
3.2 LDE with exponential polynomials coefficients
For an exponential polynomial (1.3), at least one of the degrees of the polynomials $Q_{j},j=1,\ldots,l,$ is bigger than one, or it is an exponential sum or a polynomial. Set $s=\max\{\deg(Q_{j})\}\geq 1.$ An exponential polynomial can be rewritten in the normalized form
$G(z)=\hat{G_{1}}(z)+\sum_{j=2}^{k}\hat{G}_{j}(z)\exp\left({q_{j}z^{s}}\right)$ (3.7)
where $k\in\mathbb{N},$ the $q_{j}$ are pairwise distinct non-zero constants for $2\leq j\leq k$, and $q_{1}=0;$ the coefficients $\hat{G_{j}}(z)$ are exponential polynomials of growth order $\rho(\hat{G_{j}})\leq s-1$ such that $\hat{G_{j}}(z)\not\equiv 0,$ for $2\leq j\leq{k},$ and ${k}\leq l.$
Now, we step forward to deal with the equation with exponential polynomials coefficients. A differential equation
$f^{(n)}(z)+G_{n-1}(z)f^{(n-1)}(z)+\cdots+G_{1}(z)f^{\prime}(z)+G_{0}(z)f(z)=0$ (3.8)
with the exponential polynomials coefficients
$G_{m}(z)=\sum_{t=1}^{l_{m}}P_{m,t}(z)\exp(Q_{m,t}(z)),\kern 5.0pt0\leq m\leq n-1,\kern 5.0ptl_{m}\in\mathbb{N},$ (3.9)
where $P_{m,t}$ and $Q_{m,t}$ are polynomials, can be transformed into the equation
$\sum_{j=1}^{k}\exp\left(q_{j}z^{s}\right)\left\{\sum_{m=0}^{n}\hat{G}_{m,j}(z)f^{(m)}(z)\right\}=0,$ (3.10)
where $k$ is the maximal number of pairwise different complex number $q_{j},$ appearing in all coefficients of $z^{s}$ in $Q_{m,t}$, $q_{1}=0$ and $s=\max\{\deg(Q_{m,t})\}$. Thereby, the degree of every polynomial $\hat{G}_{m,j}(z)$ is smaller or equal to s1.s-1.
The complex numbers $q_{j}$} are called leading coefficients to the degree $s$. The differential equation
$\sum_{m=0}^{n}\hat{G}_{m,j}(z)f^{(m)}(z)=0$ (3.11)
is called a coefficient differential equation at the conjugated leading coefficient $\overline{q}_{j}$ of (3.8).
To describe logically, we denote the convex hull to equation (3.10) as $co(W)$, conjugated leading coefficients as $\overline{q}_{j_{1}},j_{1}=1,\ldots,k$; denote the $r_{1}$ many vertexes of $co(W)$ as $\overline{\omega}_{j_{1}},1\leq j_{1}\leq r_{1}\leq k,$ in counterclockwise order and critical rays $\arg z=\eta_{j_{1}},$ defined as the ray originated at 0 with the direction of the outer normal to the segment between $\overline{\omega}_{j_{1}}$ and $\overline{\omega}_{j_{1}+1}$; denote sectors $S_{j_{1}}(\varepsilon),T_{j_{1}}(\varepsilon)$ as
$S_{j_{1}}(\varepsilon)=\{z|\eta_{j_{1}-1}+\varepsilon\leq\arg z\leq\eta_{j_{1}}-\varepsilon\},\kern 5.0ptT_{j_{1}}(\varepsilon)=\{z|\eta_{j_{1}}-\varepsilon<\arg z<\eta_{j_{1}}+\varepsilon\},$
$1\leq j_{1}\leq r_{1}\leq k$.
[…]
Figure 2: leading coefficients $\{q\}$ of corresponding exponential polynomial terms
In Figure 2, $n(\cdot)$ is a number controlled by the variables $\cdot$". To write inductively, we set $n(j_{1},\ldots,j_{t-1})=k_{j_{1},\ldots,j_{t-1}}$. We have already separated the dominant exponential polynomial terms $e^{q_{j_{1}}z^{s}}$ in (3.10) and their coefficient differential equation (3.11) at $q_{j_{1}}$. Within each coefficient differential equation (3.11) at $q_{j_{1}}$, we can do this transformation again and sparate the dominant exponential polynomial terms $e^{q_{j_{1},j_{2}}z^{s-1}}$ and their coefficient differential equations at $q_{j_{1},j_{2}}$, where $q_{j_{1},j_{2}}$ are leading coefficients to the degree s1s-1. We denote the convex hull to equation (3.11) as $co(W_{j_{1}}),j_{1}=1,\ldots,k$; denote the $r_{j_{1},j_{2}}$ many conjugated leading coefficients on the vertexes of convex hull $co(W_{j_{1}})$ as $\overline{\omega}_{j_{1},j_{2}},j_{1}=1,\ldots,k,j_{2}=1,\ldots,k_{j_{1}}$ in counterclockwise order, where $k_{j_{1}}$ is the maximal number of pairwise different complex number $q_{j_{1},j_{2}}$; denote the $r_{j_{1},j_{2}}$ many critical rays $\arg z=\eta_{j_{1},j_{2}},$ defined as the ray originated at 0 with the direction of the outer normal to the segment between $\overline{\omega}_{j_{1},j_{2}}$ and $\overline{\omega}_{j_{1},j_{2}+1}$, corresponding sectors as
$S_{j_{1},j_{2}}(\varepsilon)=\{z|\eta_{j_{1},j_{2}-1}+\varepsilon\leq\arg z\leq\eta_{j_{1},j_{2}}-\varepsilon\},\kern 5.0ptT_{j_{1},j_{2}}(\varepsilon)=\{z|\eta_{j_{1},j_{2}}-\varepsilon<\arg z<\eta_{j_{1},j_{2}}+\varepsilon\},$
$1\leq j_{1}\leq r_{1}\leq k,1\leq j_{2}\leq r_{j_{1}}\leq k_{j_{1}}$.
Repeating the division of original equation by the dominant exponential polynomial terms $e^{q_{j_{1}}z^{s}},e^{q_{j_{1},j_{2}}z^{s-1}},\ldots,e^{q_{j_{1},\ldots,j_{s}}z}$ for $s$ times inductively, if we retain the symbols and get a coefficient differential equation with only polynomials coefficients, we call it a fundamental coefficient differential equation at $\overline{\omega}_{j_{1},\ldots,j_{s}},1\leq j_{t}\leq r_{j_{1},\ldots,j_{t-1}}\leq k_{j_{1},\ldots,j_{t-1}},1\leq t\leq s,r_{j_{0}}=r_{1},k_{j_{0}}=k$. $k_{j_{1},\ldots,j_{t}}$ is the maximal number of pairwise different complex number $q_{j_{1},\ldots,j_{t}},$ see Figure 2, the $r_{j_{1},\ldots,j_{t}}$ many vertexes of $co(W_{j_{1},\ldots,j_{t-1}})$ are denoted by $\overline{\omega}_{j_{1},\ldots,j_{t}}$ in counterclockwise order and critical rays $\arg z=\eta_{j_{1},\ldots,j_{t}},$ defined as the ray originated at 0 with the direction of the outer normal to the segment between $\overline{\omega}_{j_{1},\ldots,j_{t}}$ and $\overline{\omega}_{j_{1},\ldots,j_{t}+1}$, corresponding sectors as
$\begin{split}S_{j_{1},\cdots,j_{t}}(\varepsilon)&=\{z|\eta_{j_{1},,\cdots,j_{t}-1}+\varepsilon\leq\arg z\leq\eta_{j_{1},\cdots,j_{t}}-\varepsilon\},\\ \ T_{j_{1},\cdots,j_{t}}(\varepsilon)&=\{z|\eta_{j_{1},\cdots,j_{t}}-\varepsilon<\arg z<\eta_{j_{1},\cdots,j_{t}}+\varepsilon\},\end{split}$
where $1\leq j_{t}\leq r_{j_{1},\ldots,j_{t-1}}\leq k_{j_{1},\ldots,j_{t-1}},1\leq t\leq s.$ See Figure 3.
[…]
Figure 3: ${co(W_{j_{1},\ldots,j_{t-1}})},\arg z=\eta_{j_{1},\ldots,j_{t}},S_{j_{1},\ldots,j_{t}}(\varepsilon),T_{j_{1},\ldots,j_{t}}(\varepsilon)$
Additionally, we need some relations about the distribution of $z$ and $z^{s}$, where $s\in\mathbb{C}$. Let $0\leq\alpha<1$ and
$\Pi:=\{z\in\mathbb{C}||\Im(z)|<{|z|}^{\alpha}\}.$
If $\Phi$ is formed from $\Pi$ by rotating at a fixed angle around the origin, then $\Phi$ is called the parabolic strip of aperture $\alpha$ along the axis that by rotation around the same angle arising from the positive real axis.
The Lemma is given originally by Droletz [7] in his dissertation, which is compiled below.
Lemma 3.8
Let $s\in\mathbb{N},$ and $\Phi$ be a parabolic strip of aperture $\alpha$ with $z^{s}\in\Phi$. Then there are parabolic strips $\Phi^{(1)},\Phi^{(2)},\ldots,\Phi^{(s)}$, of aperture $\alpha$, such that for a sufficiently large $|z|$ the value $z$ lies in the union
$\bigcup_{k=1}^{s}\Phi^{(k)}.$
Moreover, if the argument to the axis of symmetry of $\Phi$ is $s\eta,$ then the argument to the axis of $\Phi^{(k)}$ are $\eta+\frac{2k\pi}{s}$ individually, for $k=1,\ldots,s.$
Now a similar conclusion to Theorem 3.2 is given for exponential polynomials coefficients.
Theorem 3.9
Every finite order trancendental entire solution $f$ of (3.8) satisfies
$\log|f(z)|=\Re G_{\theta}(z^{1/N_{\theta}})+O(\log|z|),$ (3.12)
for some polynomials $G_{\theta}$ in $z^{1/N_{\theta}}$, $N_{\theta}\in\mathbb{N^{*}},$ outside a $r$-value set $E$ of finite linear measure, in $\theta\leq\arg z\leq\theta+h,\theta\in(0,2\pi]$, sufficiently small $h>0$, and enough large $|z|$, besides two kinds of areas:
  1. 1.
    $z$ in
    $T_{j_{1},\ldots,j_{t}}^{(i)}(\varepsilon)=\left\{z:\frac{\eta_{j_{1},\cdots,j_{t}}+2i\pi-\varepsilon}{s+1-t}<\arg z<\frac{\eta_{j_{1},\cdots,j_{j}}+2i\pi+\varepsilon}{s+1-t}\right\},$
    where $1\leq i\leq s+1-t,1\leq t\leq s;$ $\varepsilon>0$, $1\leq j_{t}\leq r_{j_{1},\ldots,j_{t-1}}\leq k_{j_{1},\ldots,j_{t-1}},1\leq t\leq s$; $\eta_{j_{1},\ldots,j_{t}}$ are arguments of critical rays defined above;
  2. 2.
    the logarithmic semi-strips
    $0\leq\arg z-\xi_{j_{1},\cdots,j_{s},M_{{j_{1},\cdots,j_{s}}}}<\frac{\log^{+}|z|}{{|z|}^{\frac{1}{p_{j_{1},\cdots,j_{s}}}}},p_{j_{1},\cdots,j_{s}}\in\mathbb{N^{*}},M_{j_{1},\cdots,j_{s}}\in\mathbb{N}$
    in
    $\mathcal{S}_{j_{1},\cdots,j_{s}}(\varepsilon)=\bigcap_{t=1,2,\cdots,s}\left(\bigcup_{i=1,\cdots,s+1-t}S_{j_{1},\cdots,j_{t}}^{(i)}(\varepsilon)\right),$
    where $\arg z=\xi_{j_{1},\cdots,j_{s},M_{{j_{1},\cdots,j_{s}}}}$ are the Stokes rays to the coefficient differential equations at $\overline{\omega}_{j_{1},\ldots,j_{s}}$ in $\mathcal{S}_{j_{1},\cdots,j_{s}}(\varepsilon).$
Thus, all finite order transcendental entire solutions are of completely regular growth.
Note. The polynomial $G_{\theta}$ is piecewise constant and changes when $\arg z=\theta$ is a Stokes ray, but these changes occur only for finite many times.
Example 3.10
The function $f(z)=e^{z^{2}}$ solves the equation
$\begin{split}(2z-e^{z}+\cos(z^{2}))f^{\prime\prime}+(e^{z}+e^{2z}-2-4z^{2}-2z\sin(z^{2})-2z\cos(z^{2}))f^{\prime}+&\\ (2e^{z}+4z^{2}e^{z}-2ze^{z}-2ze^{2z}+4z^{2}\sin(z^{2})-2\cos(z^{2}))f=0.\end{split}$ (3.13)
The normal form of (3.13) is
$\begin{split}&e^{iz^{2}}\left\{\frac{1}{2}f^{\prime\prime}+(iz-z)f^{\prime}-(2iz^{2}+1)f\right\}+\\ &e^{-iz^{2}}\left\{\frac{1}{2}f^{\prime\prime}-(iz+z)f^{\prime}+(2iz^{2}-1)f\right\}+\\ &e^{2z}\left\{f^{\prime}-2zf\right\}+e^{z}\left\{-f^{\prime\prime}+f^{\prime}+(2-2z+4z^{2})f\right\}+\left\{2zf^{\prime\prime}-(2+4z^{2})f^{\prime}\right\}=0.\end{split}$
Set the differential polynomial
$\begin{split}&G_{1}(f)=\frac{1}{2}\frac{f^{\prime\prime}}{f}+(iz-z)\frac{f^{\prime}}{f}-(2iz^{2}+1),\\ &G_{2}(f)=\frac{1}{2}\frac{f^{\prime\prime}}{f}-(iz+z)\frac{f^{\prime}}{f}+(2iz^{2}-1),\\ &G_{3}(f)=\frac{f^{\prime}}{f}-2z,G_{4}(f)=-\frac{f^{\prime\prime}}{f}+\frac{f^{\prime}}{f}+2-2z+4z^{2},G_{5}(f)=2z\frac{f^{\prime\prime}}{f}-(2+4z^{2})\frac{f^{\prime}}{f},\end{split}$
and we afford $G_{i}(e^{z^{2}})=0,i=1,\ldots,5.$
4 Proofs of main results
4.1 Proof of Theorem 3.2
Proof. Dividing $f$ of both sides of (3.3), we get
$\sum_{j=0}^{N_{0}}e^{\lambda_{j}z}\left(a_{n,j}f^{(n)}/f+a_{n-1,j}(z)f^{(n-1)}/f+\cdots+a_{0,j}(z)\right)=0.$ (4.1)
For $k,j\in 0,\ldots,N_{0}$ we set
$\kappa_{k,j}(z):=\left|\frac{e^{\lambda_{k}z}\left(\sum\limits_{t=0}^{n}a_{t,k}\dfrac{f^{(t)}}{f}\right)}{e^{\omega_{j}z}\left(\sum\limits_{t=0}^{n}a_{t,j}\dfrac{f^{(t)}}{f}\right)}\right|=e^{\Re{(\lambda_{k}-\omega_{j})}z}\left|\frac{\left(\sum\limits_{t=0}^{n}a_{t,k}\dfrac{f^{(t)}}{f}\right)}{\left(\sum\limits_{t=0}^{n}a_{t,j}\dfrac{f^{(t)}}{f}\right)}\right|.$ (4.2)
From Lemma 3.1, for $z\in S_{j}(\varepsilon)$, $\lambda_{k}\neq\omega_{j}$ ($\omega_{j}$ is a vertex of the minimal convex set containing all the $\lambda_{k},k=0,\ldots,N_{0}$), it follows that
$\Re((\lambda_{k}-\omega_{j})z)\leq-|z|\cdot|\lambda_{k}-\omega_{j}|\cdot|\sin\varepsilon|<0.$
By Gundersen [13, Corollary 3], if we set $\rho=\rho(f)$ as growth order of $f$, there exists $\varepsilon_{0}>0$ and a set $E$ with finite linear measure, such that
$\left|\frac{f^{(k)}}{f}\right|<{|z|^{k(\rho+\varepsilon_{0})}},\quad|z|\notin E.$
We denote
$L_{j}(f)=\sum\limits_{t=0}^{n}a_{t,j}\dfrac{f^{(t)}}{f},$
so (4.1) is transformed to
$\sum_{j=0}^{N_{0}}e^{\lambda_{j}z}L_{j}(f)=0.$
Then, we claim that the equality below is true:
$L_{j}(f)=o\left(\frac{1}{z^{N_{1}}}\right),\kern 5.0pt\text{for any}\kern 5.0ptN_{1}\in\mathbb{N},\kern 5.0ptz\in S_{j}(\varepsilon),\kern 5.0pt|z|\notin E.$ (4.3)
Otherwise, there is a sequence $\{z_{i}\},z_{i}\in S_{j}(\varepsilon),|z_{i}|\notin E$, with $\lim\limits_{i\rightarrow\infty}z_{i}=\infty,$ such that
$\left|\sum\limits_{t=0}^{n}a_{t,j}(z_{i})\frac{f^{(t)}(z_{i})}{f(z_{i})}\right|>\frac{C}{{|z_{i}|}^{N_{1}}},$ (4.4)
for any $C,N_{1}\in\mathbb{N}$. It holds that $\kappa_{k,j}(z_{i})=o(1),$ for the reason that $e^{\Re{(\lambda_{k}-\omega_{j})}z}$ decreases exponentially and the term in the absolute value in (4.2) is up to a polynomial rate of growth. From (4.1),
$e^{\lambda_{k}z_{i}}\left(\sum\limits_{t=0}^{n}a_{t,k}(z_{i})\dfrac{f^{(t)}(z_{i})}{f(z_{i})}\right)=o\left(e^{\omega_{j}z_{i}}\left(\sum\limits_{t=0}^{n}a_{t,j}(z_{i})\dfrac{f^{(t)}(z_{i})}{f(z_{i})}\right)\right),$
which leads that for large $|z_{i}|$,
$e^{\omega_{j}z_{i}}\left(\sum\limits_{t=0}^{n}a_{t,j}(z_{i})\dfrac{f^{(t)}(z_{i})}{f(z_{i})}\right)(1+o(1))=0,$
a contradiction to the assumption (4.4).
Therefore, for $z\in S_{j}(\varepsilon),|z|\notin E$, (4.3) can be transformed into an equation
$f^{(n)}+P_{n-1,j}f^{(n-1)}+\cdots+P_{1,j}f^{\prime}+\left(P_{0,j}+o\left(\frac{1}{z^{N_{1}+\deg(a_{n,j})}}\right)\right)f=0,$ (4.5)
where $P_{i,j}=a_{i,j}/a_{n,j},i=0,\ldots,n-1$. Briefly, denote $Q_{0,j}$ as the last coefficient in equation (4.5). Furthermore, it is deduced that every $P_{i,j}$ can be expanded as
$P_{i,j}(z)=z^{d_{i,j}}\sum_{v=0}^{\infty}\frac{b_{i,j,v}}{z^{v}},\quad|z|>|z_{0}|,$ (4.6)
for $d_{i,j}\in\mathbb{N},b_{i,j,v}\in\mathbb{C},i=1,\ldots,n-1,$ $j=0,\ldots,N_{0}^{\prime},$ enough large $|z_{0}|$, and $z\in S_{j}(\varepsilon)$. $Q_{0}$ is consequently in the asymptotic form of
$Q_{0,j}(z)\sim z^{d_{0,j}}\sum_{v=0}^{\infty}\frac{b_{0,j,v}}{z^{v}},\quad|z|>|z_{0}|,$ (4.7)
for $d_{0,j}\in\mathbb{N},b_{0,j,v}\in\mathbb{C},$ $j=0,\ldots,N_{0}^{\prime},$ enough large $|z_{0}|$, $z\in S_{j}(\varepsilon),$ and $|z|\notin E$.
Next, we introduce the essence of $|z|\in E$. In the proof of Gundersen’s result [13, Corollary 3], the essential idea is to use Cartan’s Lemma [13, Lemma 9] to dig out constructive disks containing the zeros and poles of $f^{(j)}$ (He defaults $f$ as a meromorphic function and here we consider $j=0$ and $f$ entire). And introducing the set of exceptions $E$ with respect to $r$ is exactly the sum of the diameters of these disks. We recall the constructive part of this [13, pp. 100-101], which begins by partitioning the complex plane into a merge of countably many annular strips, i.e., there exists an integer $n_{0}>0$, and for an arbitrary integer $\nu\geq n_{0}$ and a constant $\alpha>1$, there is
$\alpha^{\nu}\leq|z|\leq\alpha^{\nu+1}.$ (4.8)
In each such strip, there will be only a finite number of zeros of $f$ and closed disks $B_{\nu,1},\ldots,B_{\nu,p_{\nu}}$ containing them (note that the center of the circle here is not necessarily a zero of $f$, and a single disk may contain more than one zero). For any $\varepsilon_{0}>0$, by [13, p. 101], when we take the sum of the diameters of these disks not exceeding $4\alpha^{-\varepsilon_{0}\nu}$ the corresponding $E$ is obtained as a finite linear measure by Cartan’s Lemma [13, Lemma 9].
Thus, (4.5) can be modified as
$f^{(n)}+P_{n-1,j}f^{(n-1)}+\cdots+P_{1,j}f^{\prime}+\left(P_{0,j}+o\left(\frac{1}{z^{N_{1}+\deg\{a_{n,j}\}}}\right)\right)f=0,$ (4.9)
which is established for $z\in S_{j}^{*}$, i.e., $z\in S_{j}(\varepsilon)$ and
$z\notin\tilde{B}:=\bigcup_{\nu=n_{0}}^{\infty}\bigcup_{t=1}^{p_{\nu}}B_{\nu,t},~n_{0}>N.$
In a step further, it can be shown that (4.9) satisfies the form of (2.9) if $f$ only possesses finite number of zeros in $S_{j}(\varepsilon)$, which means $\tilde{B}$ disappears when $|z|>|z_{0}|$ for sufficiently large $|z_{0}|$. Using Lemma 2.4, there exists a polynomial $G_{\theta}$ in $z^{1/p_{j}}$, $p_{j}\in\mathbb{N}$, and a set of zero relative measure $E_{\theta}$ such that
$\log|f(z)|=\Re G_{\theta}\left(z^{\frac{1}{p_{j}}}\right)+O(\log|z|)$ (4.10)
for $|z|\notin E_{\theta},z\in S_{j}(\varepsilon)$ as $z\rightarrow\infty$ in $\theta\leq\arg z\leq\theta+h,h>0$, and outside of a logarithmic semi-strip $0\leq\arg z-\theta<C\frac{\log^{+}|z|}{|z|^{1/p_{j}}}$. The last semi-strip occurs only if $\arg z=\theta$ is a Stokes ray in $S_{j}$, otherwise, we set $C=0$. Thus, it is natural to be inferred $f(z)$ is of completely regular growth on every ray $\theta\in S_{j}(\varepsilon)\backslash\{\arg z=\xi_{j1},\ldots,\arg z=\xi_{j_{m_{j}}}\},j=0,\ldots,N_{0},$ where $\{\arg z=\xi_{j1},\ldots,\arg z=\xi_{j_{m_{j}}}\}$ are the Stokes rays of equation (4.5) in $S_{j}(\varepsilon),m_{j}\in\mathbb{N}.$
If $f$ has infinitely many zeros in $S_{j}(\varepsilon)$, the associated exceptional set $\tilde{B}$ consists of countably many closed disks constructed via Cartan’s lemma. Following the notation in (4.8), the complex plane is partitioned into annular strips $\alpha^{\nu}\leq|z|\leq\alpha^{\nu+1}$ with $\alpha>1$ and $\nu\geq n_{0}$. Within each strip, a finite collection of disks $B_{\nu,t}=B(\zeta_{\nu,t},r_{\nu,t})$, $t=1,\dots,p_{\nu}$, covers the zeros of $f$ in that strip. The total linear measure of these disks is finite, i.e., $\sum_{\nu,t}4\alpha^{-\varepsilon_{0}\nu}<\infty$. The classical asymptotic theorems of Wasow and Steinmetz require holomorphy on a full sector, which is not guaranteed on $S_{j}^{*}$ due to the presence of these disks.
To overcome this, we employ the modified Wasow theorem (Theorem 6.5) on a subregion $S_{0}\subset S_{j}^{*}\subset S_{j}(\varepsilon)$ that satisfies the uniform interior condition. The construction of $S_{0}$ involves slightly enlarging each exceptional disk: define $B^{0}_{\nu,t}=B\!\left(\zeta_{\nu,t},r_{\nu,t}+\frac{\alpha_{0}}{|\zeta_{\nu,t}|}\right)$ for a fixed $\alpha_{0}>0$, and set $S_{0}=S_{j}(\varepsilon)\setminus\bigcup_{\nu,t}B^{0}_{\nu,t}$.
The crucial observation is that the total linear measure of the enlarged disks remains finite. Indeed, since the original exceptional set has finite linear measure, we have $\sum_{\nu}4\alpha^{-\varepsilon_{0}\nu}<\infty$. Moreover, because of $\alpha^{\nu}\leq|\zeta_{\nu,t}|\leq\alpha^{\nu+1}$, by [13, eq. (7.2),(7.6),(7.7)], we denote $p_{\nu}$ as the number of cartan disks which intersects with the annulus, and $n(r,1/f)$ as the counting function of zeros of $f$ inside the disk $B(0,r)$, and then we get
$p_{\nu}\leq n\left(\alpha^{\nu+2},\frac{1}{f}\right)\leq\frac{12}{\log\alpha}T(\alpha\alpha^{\nu+2},f)<\frac{12}{\log\alpha}\alpha^{(\nu+3)(\rho+\varepsilon_{0})},$
so the sum of diameter-difference between $B_{\nu,t}$ and $B_{\nu,t}^{\prime}$ is
$\sum_{\nu}\sum_{t=1}^{p_{\nu}}\left(\frac{2\alpha_{0}}{|\zeta_{\nu,t}|}\right)<\sum_{\nu}\sum_{t=1}^{p_{\nu}}\left(2\alpha_{0}\alpha^{-\nu}\right)<C\sum_{\nu}\alpha^{\nu(\rho+\varepsilon_{0}-1)},~C>0.$ (4.11)
Consequently, when $\rho(f)<1,$
$\sum_{\nu}\left(4\alpha^{-\varepsilon_{0}\nu}+\frac{\alpha_{0}}{|\zeta_{\nu,t}|}\right)<\infty,$
so the enlarged exceptional set $\bigcup_{\nu,t}B^{\prime}_{\nu,t}$ also has finite linear measure. Therefore, the subregion $S_{0}$ has full measure in $S_{j}(\varepsilon)$ and satisfies the uniform interior condition: for every $z\in S_{0}$, the disk $B(z,\alpha/|z|)$ is contained in $S_{j}(\varepsilon)\setminus\bigcup_{\nu,t}B_{\nu,t}=S^{*}$.
Since the radius $\alpha/|z|$ of interior condition is sharp in [38, Theorem 8.8] by the nature of Cauchy integral formula, we can not reduce it to $O(1/|z|^{N})$ for some integer $N$, to allow a uniform termwise-differentiable asymptotic form of $f^{\prime}(z)$ on such $S_{0}$ outside “much smaller” disks, in Corollary 6.3.
However, when $\rho(f)\geq 1$ and $n(r,1/f)$ grows extremely fast as r growing, the sum (4.11) may not be coved by a finite linear measure. In this case, we can deal with the exceptional disks on the $\xi$-plane under the tranformation $\xi=z^{n_{0}}$, $n_{0}=\lfloor\rho+1\rfloor$.
Let $g(\xi)=f(z)=f(\xi^{1/n_{0}})$. Without generality, we fix $j=1$ and abberviate the above symbols with the subscript $j$. We choose a single-valued analytic branch of $\xi^{1/n_{0}}$ corresponding to the sector $S(\varepsilon)$, and denote by $S^{\prime}(\varepsilon)$ the image of $S(\varepsilon)$ under this mapping. The exceptional set $\tilde{B}$ is mapped to $\tilde{B}^{\prime}$,
$\tilde{B}^{\prime}:=\bigcup_{\nu=n_{0}}^{\infty}\bigcup_{t=1}^{p_{\nu}}B_{\nu,t}^{\prime}=\bigcup_{\nu=n_{0}}^{\infty}\bigcup_{t=1}^{p_{\nu}}B(\xi_{\nu,t},r_{\nu,t}^{\prime}),~n_{0}>N,$
and we set ${S^{*}}^{\prime}=S^{\prime}(\varepsilon)\setminus\tilde{B}^{\prime}$. Substituting $z=\xi^{1/n_{0}}$ into the asymptotic differential equation (4.9) and using the chain rule, we obtain a new linear differential equation for $g(\xi)$:
$g^{(n)}(\xi)+\sum_{k=1}^{n-1}Q_{k}(\xi)g^{(k)}(\xi)+\Bigl(Q_{0}(\xi)+o\bigl(\xi^{-(N_{1}+\deg\{a_{n}\})/n_{0}}\bigr)\Bigr)g(\xi)=0,$ (4.12)
where the coefficients $Q_{k}(\xi)$ are holomorphic in $S^{*}$$^{\prime}$ and admit asymptotic expansions of the form
$Q_{k}(\xi)\sim\xi^{\delta_{k}}\sum_{\nu=0}^{\infty}\frac{c_{k,\nu}}{\xi^{\nu/n_{0}}},\qquad\xi\to\infty,\;\xi\in{S^{*}}^{\prime},$ (4.13)
with rational exponents $\delta_{k}$ and constants $c_{k,\nu}\in\mathbb{C}$. In parallelled, we can construct $S_{0}^{\prime}$ involves slightly enlarging each exceptional disk: define ${B^{*}_{\nu,t}}=B\!\left(\xi_{\nu,t},r_{\nu,t}^{\prime}+\frac{\alpha_{0}}{|\xi_{\nu,t}|}\right)$ for a fixed $\alpha_{0}>0$, and set $S_{0}^{\prime}=S^{\prime}(\varepsilon)\setminus\bigcup_{\nu,t}B^{*}_{\nu,t}$. It is easy to check the uniform interior condition: there exists a constant $\alpha^{\prime}>0$ such that for every $\xi\in S_{0}^{\prime}$ the disk $B(\xi,\alpha^{\prime}/|\xi|)$ is contained in $S^{*}$$^{\prime}$. Therefore, the modified version of Wasow’s asymtotic theorem (Theorem 6.5) can be similarly used on the $\xi$-plane.
We will not repeat the process to get a similar estimate of $\log|g(\xi)|$ outside some exceptional set via Steinmetz’s method, compared to (4.10). Instead, our main objective is to demonstrate that through transformation $\xi=z^{n_{0}}$, the expansion of the diameters of these disks is measurable. Since the pre-image of each $B^{*}_{\nu,t}$} is $B\!\left(\zeta_{\nu,t},r_{\nu,t}+\frac{\alpha_{1}}{|\zeta_{\nu,t}|^{n_{0}}}\right)\supset B_{\nu,t}$, the sum of diameter-difference in (4.11) becomes
$\sum_{\nu}\sum_{t=1}^{p_{\nu}}\left(\frac{2\alpha_{1}}{|\zeta_{\nu,t}|}\right)<\sum_{\nu}\sum_{t=1}^{p_{\nu}}\left(2\alpha_{1}\alpha^{-n_{0}\nu}\right)<C\sum_{\nu}\alpha^{\nu(\rho-n_{0}+\varepsilon_{0})},~C>0.$
In conclusion, no matter how large is $\rho(f)$, we can treat the cartan disks in Gundersen’s estimate and following operations as a finite linear measure.
Thus, we may apply Theorem 6.5 to the linear differential equation on $S^{*}$, obtaining the asymptotic representation for $f$ as in (4.10) for $z\in S_{0}$ (hence for all $z\in S_{j}(\varepsilon)$ outside a set of zero relative linear measure). On such punctured sector $\B~S_{j}(\varepsilon)\backslash\tilde${B}, Steinmetz’s analysis framework (proof of Lemma 2.2) can still be generalized. We can then only replace the exceptional condition $|z|\notin E_{\theta}$ with
$|z|\notin E_{\theta}~\text{as well as}~z\notin\tilde{B}$,
and maintain that (4.10) holds. Since the definition of c.r.g. with exceptional set $C_{0}$-set and $E_{0}$-set are essentially equivalent, referring to Section 2.1, here we simplify the discussion about them. $\Box$
4.2 Proof of Theorem 3.5
Proof. To deal with the case $z\in T_{j}(\varepsilon)$, without generality, we set $j=1$. Due to Corollary 3.3, the normal form (3.3) possesses asymtptotic form (3.5) in $T_{1}(\varepsilon)$. Dividing $f$ by both sides of (3.5), we get
$\sum_{\overline{\lambda}_{k}\in s_{1}}e^{\lambda_{k}z}\sum_{m=0}^{n}c_{m,k}z^{d_{m,k}}(1+\varepsilon(z))\frac{f^{(m)}}{f}=0.$ (4.14)
First, we consider $z\in T_{1}(\varepsilon)\cap U_{1}$. Contrasting Lemma 3.1 with Corollary 3.4, the summation of (4.14) can be divided by
$\begin{split}&e^{\omega_{1}z}\sum_{m=0}^{n}c_{m,1}z^{d_{m,1}}(1+\varepsilon(z))\frac{f^{(m)}}{f}+\sum_{\tau_{p,m}\notin\tilde{s}_{1}}e^{\lambda_{p}z}\sum_{m=0}^{n}c_{m,p}z^{d_{m,p}}(1+\varepsilon(z))\frac{f^{(m)}}{f}\\ +&\sum_{\tau_{p,m}\in\tilde{s}_{1}}e^{\lambda_{p}z}\sum_{m=0}^{n}c_{m,p}z^{d_{m,p}}(1+\varepsilon(z))\frac{f^{(m)}}{f}=0.\end{split}$ (4.15)
We denote
$L_{k}^{*}(f)=\sum\limits_{m=0}^{n}c_{m,k}z^{d_{m,k}}(1+\varepsilon(z))\dfrac{f^{(m)}}{f},$
so (4.15) is abbreviated to
$e^{\omega_{j}z}L_{1}^{*}(f)+\sum_{\tau_{p,m}\notin\tilde{s}_{1}}e^{\lambda_{p}z}L_{p}^{*}(f)+\sum_{\tau_{p,m}\in\tilde{s}_{1}}e^{\lambda_{p}z}L_{p}^{*}(f)=0.$
Set $\max_{m}\{d_{m,1}\}$ as $d_{M}$. Then, we claim that the equality below is true:
$\frac{L_{1}^{*}(f)}{z^{d_{M}}}=o\left(z^{N_{1}}\right),\kern 5.0pt\text{for some}\kern 5.0ptN_{1}\in\mathbb{N},\kern 5.0ptz\in T_{1}(\varepsilon)\cap U_{1},\kern 5.0pt|z|\notin E,$ (4.16)
where $E$ is a set of finite linear measure.
Otherwise, there is a sequence $\{z_{i}\},z_{i}\in T_{1}(\varepsilon)\cap U_{1},|z_{i}|\notin E$, with $\lim\limits_{i\rightarrow\infty}z_{i}=\infty,$ such that
$\left|\sum\limits_{m=0}^{n}c_{m,1}z_{i}^{d_{m,1}-d_{M}}(1+\varepsilon(z_{i}))\frac{f^{(m)}(z_{i})}{f(z_{i})}\right|>{C}{|z_{i}|}^{N_{1}},$ (4.17)
for any $C,N_{1}\in\mathbb{N}$. [13, Corollary 3] point out that if we set $\rho=\rho(f)$ as growth order of $f$, there exists $\varepsilon_{0}>0$, such that
$\left|\frac{f^{(k)}}{f}\right|<{|z|^{k(\rho+\varepsilon_{0})}},\quad z\notin E.$ (4.18)
It holds for $\tau_{p,m}\in\tilde{s}_{1}$ that
$\kappa_{p}(z_{i}):=\left|\frac{e^{\lambda_{p}z_{i}}\left(\sum\limits_{m=0}^{n}c_{m,p}z_{i}^{d_{m,p}}(1+\varepsilon(z_{i}))\dfrac{f^{(m)}}{f}\right)}{z^{d_{M}}e^{\omega_{1}z_{i}}\left(\sum\limits_{m=0}^{n}c_{m,1}z_{i}^{d_{m,1}-d_{M}}(1+\varepsilon(z_{i}))\dfrac{f^{(m)}}{f}\right)}\right|=o(1),~|z_{j}|\notin E,$
for the reason that $z^{d_{m,p}-d_{M}}e^{\Re{(\lambda_{k}-\omega_{1})}z}$ decreases exponentially by Corollary 3.4, and the remained term in the absolute value is up to a polynomial rate of growth by (4.17) and (4.18). For $\tau_{p,m}\notin\tilde{s}_{1}$, by Corollary 3.4 and [13, Corollary 3],
$\kappa_{p}(z_{i})=o\left(|z_{i}|^{-v+n(\rho+\varepsilon_{0})-N_{1}}\right)=o(1),$
because of the existence of $N_{1}>n(\rho+\varepsilon_{0})-v$. Then for both of $\tau_{p,m}\in\tilde{s}_{1}$ and $\tau_{p,m}\notin\tilde{s}_{1}$,
$e^{\lambda_{p}z_{i}}\left(\sum\limits_{m=0}^{n}c_{m,p}z_{i}^{d_{m,p}}(1+\varepsilon(z_{i}))\dfrac{f^{(m)}}{f}\right)=o\left(e^{\omega_{1}z_{i}}\left(\sum\limits_{m=0}^{n}c_{m,1}z_{i}^{d_{m,1}}(1+\varepsilon(z_{i}))\dfrac{f^{(m)}}{f}\right)\right),$
which leads that for large $|z_{i}|\notin E$,
$z_{i}^{d_{M}}e^{\omega_{1}z_{i}}\left(\sum\limits_{m=0}^{n}c_{m,1}z_{i}^{d_{m,1}-d_{M}}(1+\varepsilon(z_{i}))\dfrac{f^{(m)}}{f}\right)(1+o(1))=0,$
a contradiction to the assumption (4.17).
Therefore, for $z\in T_{1}(\varepsilon)\cap U_{1},|z|\notin E$, (4.16) can be transformed into an equation
$f^{(n)}+P_{n-1}f^{(n-1)}+\cdots+P_{1}f^{\prime}+\left(P_{0}+o\left(z^{N_{1}-d_{n,1}}\right)\right)f=0,$ (4.19)
where $P_{m}=\frac{c_{m,1}z^{d_{m,1}}}{c_{n,1}z^{d_{n,1}}}(1+\varepsilon(z)),m=0,\ldots,n-1$. It is deduced that for $z\in T_{1}(\varepsilon)\cap U_{1},$ $|z|\notin E$, every $P_{m}$ can be asymptotically expanded as
$P_{m}(z)\sim z^{d_{m,1}-d_{n,1}}\sum_{v=0}^{\infty}\frac{b_{m,v}}{z^{v}},\quad|z|>|z_{0}|~\text{large enough},$ (4.20)
for $b_{m,v}\in\mathbb{C},m=1,\ldots,n-1.$ Taking a similar analysis to the proof section 4.1, using the modified version of Lemma 2.4 on a toplogical hull, Theorem 6.5, there exists a polynomial $G_{\theta}$ in $z^{1/p_{1}}$, $p_{1}\in\mathbb{N},$ a set of zero relative measure $E_{\theta}$ such that for $|z|\notin E_{\theta}$,
$z\in\{z|\theta\leq\arg z\leq\theta+h,0<h<\varepsilon_{1}\varepsilon,\varepsilon_{1}>0\}\subset T_{1}(\varepsilon)\cap U_{1}~\text{as}~|z|\rightarrow\infty,$
and outside of a logarithmic semi-strip $0\leq\arg z-\theta<C\frac{\log^{+}|z|}{|z|^{1/p}}$, it holds that
$\log|f(z)|=\Re G_{\theta}\left(z^{\frac{1}{p_{1}}}\right)+O(\log|z|).$ (4.21)
For the case $z\in T_{1}(\varepsilon)\cap U_{K^{\prime}}$, we can take parallel consideration to the case $z\in T_{1}(\varepsilon)\cap U_{1}$. since $j$ was arbitarily chosen as 1 in the above paragraph, we have an estimate of $\log|f|$ in the form of (4.21) for every $j=0,1\ldots,N_{0}$. To denote the general symbols, we write $U_{1}$ and $U_{K^{\prime}}$, the case to $j=1$, as $U_{j,1}$ and $U_{j,K^{\prime}_{j}}$, the case to arbitrary $j$. The area
$T_{j}(\varepsilon)\cap U_{j,1}\bigcup S_{j}(\varepsilon)\bigcup T_{j-1}(\varepsilon)\cap U_{j-1,K^{\prime}_{j-1}},$
which is a continuation of the sector $S_{j}(\varepsilon)$, contains every ray in $\overline{S}_{j}$ (see the definition in Section 3.1) originating from 0 when $|z|\to\infty$, besides the two critical rays $\arg z=\eta_{j-1}$ and $\arg z=\eta_{j}$.
In conclusion, combining (4.10) and (4.21), for $|z|\notin$ a $E_{0}$-set (or $z\notin$ a $C_{0}$-set) and $|z|\to\infty$, we get an estimate of
$\log|f|=\Re G_{\theta}\left(z^{\frac{1}{p_{\theta}}}\right)+O(\log|z|)$
in every narrow sector $\{z|\theta\leq\arg z\leq\theta+h,h>0\}$ outside some possibly existing logarithmic semi-strips $0\leq\arg z-\theta<C\frac{\log^{+}|z|}{|z|^{1/p_{\theta}}},$ when $\arg z=\theta$ is a Stokes ray, and outside every critical ray $\arg z=\eta_{j}$. That means the set of arguments of c.r.g. is
[…]
where $\xi_{j,k}$, $k=1,\ldots,m_{0},$ are the Stokes rays insides $\overline{S}_{j}$. Then, by Lemma 2.1, $f(z)$ is of c.r.g. on the set of rays with arguments $\overline{\mathbb{M}}=(0,2\pi]$. $\Box$
4.3 Proof of Theorem 3.9
Proof. Change (3.8) to the normalized form (3.10). Dividing $f$ of both sides of (3.10), we afford
$\sum_{j=1}^{k}\exp\left(q_{j}z^{s}\right)\left\{\sum_{m=0}^{n}\hat{G}_{m,j}(z)\frac{f^{(m)}(z)}{f(z)}\right\}=0,$ (4.22)
for every $a\in\mathbb{R}.$ Denote $H_{j}$ by
$H_{j}(z)=\sum_{m=0}^{n}\hat{G}_{m,j}(z)\frac{f^{(m)}(z)}{f(z)}=0,$
and then (4.22) becomes
$\sum_{j=1}^{k}H_{j}(z)\exp(q_{j}z^{s})=0.$ (4.23)
The symbols $r,k,q,H,\omega,co(\cdot),S(\varepsilon),T(\varepsilon)$ here below refer to the definition in Section 3.2. Also, set $L_{p}$ as the side between the vertex point $\overline{\omega}_{p}$ and $\overline{\omega}_{{p-1}}$ of the convex hull $co(W)$, and $\Phi_{p}$ and is the parabolic strip of a sufficiently large opening along the critical ray $\arg z=\eta_{p}$ as an axis of symmetry. There is a sequence $\{z_{\mu}\}_{\mu\in\mathbb{N}}$ with $\lim_{\mu\to\infty}z_{\mu}=\infty,$ such that every $z_{\mu}^{s}$ lies in $S_{j}(\varepsilon)$. As a result of
$S_{j}(\varepsilon)\cap(\Phi_{p}\cup\Phi_{{p-1}})=\emptyset$
for $|z_{\mu}^{s}|\rightarrow\infty$, $\{z_{\mu}^{s}\}$ is contained in the area between $\Phi_{p}$ and $\Phi_{{p-1}}$.
Furthermore, we claim there is a $A\in\mathbb{R}$ such that for a sufficiently large sequence $|z_{\mu}|=r_{\mu}$,
$|H_{p}(z_{\mu})|>ce^{Ar_{\mu}^{s-1}}$ (4.24)
is true with an appropriate $c>0$. Therefore, for $t,p=1,\ldots,k,$ we find
$\kappa_{t}(z_{\mu}):=\left|\frac{H_{t}(z_{\mu})e^{q_{t}z_{\mu}^{s}}}{H_{p}(z_{\mu})e^{q_{p}z_{\mu}^{s}}}\right|=\left|\frac{H_{t}(z_{\mu})}{H_{p}(z_{\mu})}\right|e^{\Re{((q_{t}-q_{p})z_{\mu}^{s}})}<\frac{|H_{t}(z_{\mu})|}{c}e^{\Re{((q_{t}-q_{p})z_{\mu}^{s}})-Ar_{\mu}^{s-1}},t\not=p.$
Further, three circumstances will show up.
  1. Case 1.
    If $\overline{q}_{t}$ does not lie on $L_{p}$ or $L_{p+1}.$ Then it follows from Lemma 3.1:
    $\Re((q_{t}-q_{p})z_{\mu}^{s})\leq-|z_{\mu}^{s}||q_{t}-q_{p}||\sin\varepsilon|,$
    with $z_{\mu}^{s}\in S_{p}(\varepsilon)$ a sufficiently small $\varepsilon>0.$ Then it follows
    $\kappa_{t}(z_{\mu})\leq\exp(-|z_{\mu}^{s}||q_{t}-q_{p}||\sin\varepsilon|-(A-B)r_{\mu}^{s-1})=o(1),$
    for a $B\in\mathbb{R}$, except for a set $E$ in $r=|z|$ of finite linear measure, since $f^{(m)}/f$ is of polynomial growth except for some set $E$ in $r=|z|$ of finite linear measure, see Gundersen [13, Corollary 3], such that
    $\frac{1}{c}|H_{t}(z_{\mu})|=\frac{1}{c}\left|\sum_{m=0}^{n}\hat{G}_{m,j}(z_{\mu})\frac{f^{(m)}(z_{\mu})}{f(z_{\mu})}\right|\leq e^{Br_{\mu}^{s-1}},$ (4.25)
    with $|z_{\mu}|\notin E,z_{\mu}\rightarrow\infty,t\neq p$.
  2. Case 2.
    If $\overline{q}_{t}$ is on $L_{p},t\neq p$. Without loss of generality, we assume the axis of symmetry of $\Phi_{p}$ is the positive axis. Then $q_{t}-q_{p}$ is imaginary and thus
    $q_{t}-q_{p}=-i|q_{t}-q_{p}|.$
    For the reason that $\{z_{\mu}\}$ are not in the parabolic strips, it is followed by the definition of parabolic strips that
    $\kappa_{t}(z_{\mu})<$ $\frac{|H_{t}(z_{\mu})|}{c}e^{-{(|q_{t}-q_{p}|\Im(z_{\mu}^{s}}))-Ar_{\mu}^{s-1}}$
    $\leq$ $\frac{|H_{t}(z_{\mu})|}{c}e^{\left(-{|q_{t}-q_{p}|r_{\mu}^{s\alpha}}\left(1-A_{1}r_{\mu}^{s\left(1-\alpha-\frac{1}{s}\right)}\right)\right)}.$
    Now, $\alpha$ can be chosen from the interval (11/s,1)(1-1/s,1) and
    $\kappa_{t}(z_{\mu})=o(1).$ (4.26)
  3. Case 3.
    If $\overline{q}_{t}$ lies on $L_{p+1},t\neq p.$ As in Case 2, we have $\kappa_{t}(z_{\mu})=o(1).$
Together with (4.23),(4.25),(4.26),
$\sum_{j=1}^{k}H_{j}(z_{\mu})e^{q_{t_{s}}z_{\mu}^{s}}=(1+o(1))H_{p}(z_{\mu})e^{q_{p_{s}}z_{\mu}^{s}}$ (4.27)
for every $a\in\mathbb{R},z_{\mu}\in S_{p}(\varepsilon),|z_{\mu}|\notin E.$ Then for (4.24), we have
$\frac{\left|\sum\limits_{j=1}^{k}H_{j}(z_{\mu})e^{q_{j}z^{s}_{\mu}}\right|}{\left|e^{q_{p}z_{\mu}^{s}}\right|}\geq c(1+o(1))e^{Ar_{\mu}^{s-1}},$ (4.28)
where $c>0,z_{\mu}\in S_{p}(\varepsilon),|z_{\mu}|\notin E.$ Then it follows
$\left|\sum_{j=1}^{k}H_{j}(z_{\mu})e^{q_{j}z_{m}u^{s}}\right|\geq c(1+o(1))e^{Ar_{\mu}^{s-1}-|q_{p}z_{\mu}^{s}|},\kern 5.0pt|z_{\mu}|\notin E,$
but it contradicts (4.23). Consequently, there are only two possibilities:
  1. (I)
    The assumption (4.24) is incorrect. Then for every $a\in\mathbb{R}$, vertex point $\overline{\omega}_{p}$ of $co(W)$, we have
    $H_{p}(z_{\mu})=o(e^{ar_{\mu}^{s-1}}),z_{\mu}\in S_{p}(\varepsilon).$ (4.29)
  2. (II)
    The sequence $z_{\mu}^{s}$ lies in $T_{p}$ for a sufficiently large $|z_{\mu}|$, the axis of symmetry is the critical ray $\arg z^{s}_{\mu}=\eta_{p}$ of the convex hull $co(W)$. Thus, there are $s$ subsequences $\{z_{\mu 1},\ldots,z_{\mu s}\}$ of the sequence $\{z_{\mu}\}(\mu\in\mathbb{N})$ and $s$ branches $T_{p}^{(1)},\ldots,T_{p}^{(s)}$ of $T_{p}$, such that $z_{\mu j}$ lies in subsectors
    $T_{p}^{(j)}=\left\{z:\frac{\eta_{p}+2j\pi-\varepsilon}{s}<\arg z<\frac{\eta_{p}+2j\pi+\varepsilon}{s}\right\},\quad 1\leq j\leq s.$
Now we set $p=p_{1},$ $H_{p_{1},p_{2}}$ be a term of the coefficient differential equation to $\overline{\omega}_{{p_{1},p_{2}}}$, the vertex of the convex hull of $co(W_{p_{1}})$. If $W_{p_{1}}$ vanishes, we can take $p_{1}=1$ such that the coefficient differential equation does not disappear. We may always denote the subsequence of $\{z_{\mu}\}$ as $\{z_{\mu}\}$, consequently arrive to
$H_{p_{1},p_{2}}(z_{\mu})=o\left(e^{ar_{\mu}^{s-2}}\right),z_{\mu}^{s}\in S_{p_{1}}(\varepsilon)\cap S_{p_{1},p_{2}}(\varepsilon),$
where $j=1,\ldots,s;$ $i=1,\ldots,s-1;1\leq p_{2}\leq r_{p_{1}}\leq k_{p_{1}}.$ Repeatedly, we get a relationship
$H_{p_{1},p_{2},\cdots,p_{s}}(z_{\mu})=o\left(e^{ar_{\mu}^{0}}\right)=o(1),a\in\mathbb{R},z_{\mu}^{s}\in\bigcap_{j=1,2,\cdots,s}S_{p_{1},\cdots,p_{j}}(\varepsilon),$
Therefore, there exist the sequence $\{z_{\mu}\}(\mu\in\mathbb{N})$ and $i$ branches $T_{p_{1},\ldots,p_{j}}^{(1)},\ldots,T_{p_{1},\ldots,p_{j}}^{(i)}$ of $T_{p_{1},\ldots,p_{j}}(\varepsilon)$, such that $z_{\mu}$ does not lie in subsectors
$T_{p_{1},\ldots,p_{j}}^{(i)}(\varepsilon)=\left\{z:\frac{\eta_{p_{1},\cdots,p_{j}}+2i\pi-\varepsilon}{s+1-j}<\arg z<\frac{\eta_{p_{1},\cdots,p_{j}}+2i\pi+\varepsilon}{s+1-j}\right\},$
where $1\leq i\leq s+1-j,1\leq j\leq s.$
We note that here $H_{p_{1},p_{2},\cdots,p_{s}}(z_{\mu})=o(z_{\mu}^{-n_{0}})$ for any $n_{0}\in\mathbb{N}$. Otherwise, we will get a contradiction with the same analysis under (4.3). Since $H_{p_{1},p_{2},\cdots,p_{s}}(z)=0$ is a linear differential equation with polynomial coefficients, we can construct an equation which coefficients have asymptotic forms in sectors
$\bigcap_{j=1,2,\cdots,s;\atop i=1,\cdots,s+1-j}S_{p_{1},\cdots,p_{j}}^{(i)}(\varepsilon),$
where
$S_{p_{1},\ldots,p_{j}}^{(i)}(\varepsilon)=\left\{z:\frac{\eta_{p_{1},\cdots,p_{j}-1}+2i\pi+\varepsilon}{s+1-j}\leq\arg z\leq\frac{\eta_{p_{1},\cdots,p_{j}}+2i\pi-\varepsilon}{s+1-j}\right\}.$
By the similar proof of theorem 3.2 and Lemma 2.4 (when Lemma 2.4 lose effect for infinite many zeros of $f$ in each sector, we take the same discussion with (4.10) into account, which is abbreviated in this part), we afford
$\log|f(z_{\mu})|=\Re G_{\theta}(z_{\mu}^{1/N_{\theta}})+O(\log|z_{\mu}|),z_{\mu}\in\bigcap_{j=1,2,\cdots,s}\left(\bigcup_{i=1,\cdots,s+1-j}S_{p_{1},\cdots,p_{j}}^{(i)}(\varepsilon)\right),$ (4.30)
with some polynomial $G_{\theta}$ in $z^{1/N_{\theta}}$, for $N_{\theta}\in\mathbb{N},$ $\theta\in(0,2\pi],$ $|z_{\mu}|\notin E$, as $z_{\mu}\rightarrow\infty,$ in $\theta\leq\arg z_{\mu}\leq\theta+h,h>0,$ outside of a logarithmic semi-strip $0\leq\arg z_{\mu}-\theta<C\frac{\log^{+}|z_{\mu}|}{|z_{\mu}|^{1/p}}$. The last semi-strip occurs only if $\arg z_{\mu}=\theta$ is a Stokes ray, otherwise, we set $C=0$.
In this proof, we will not repeat the analysis of the Cartan discs for the uniform interior expansion. It is always contained within a finite linear measure. See the ending of section 4.1.
In the final, we consider the property of completely regular growth of $f$. (4.30) holds on any ray from the original point except for finitly many possible rays – all stokes rays in
$\bigcap_{j=1,2,\cdots,s}\left(\bigcup_{i=1,\cdots,s+1-j}S_{p_{1},\cdots,p_{j}}^{(i)}(\varepsilon)\right),$
and all axes of symetry in $T^{(i)}_{p_{1},\ldots,p_{j}}$ for $i=1,\ldots,s+1-j;j=1,\ldots,s$ and admissble $p_{1},\ldots,p_{s}$. Then, by Lemma 2.1, $f(z)$ is of completely regular growth on the set of rays with argument $(0,2\pi]$. $\Box$
5 A dynamical aspect of c.r.g. solutions of LDEs
5.1 A nested concept – exponential polynomial type
Observing the two main Theorems 3.2 and 3.9, the solutions of completely regular growth have asymptotic expressions
$\sum_{j=0}^{n}e^{P_{j}(z)}z^{c_{j}}Q_{j}(z,\log z)$ (5.1)
in every sector except for at most finite many narrow sectors. Here $P_{j}(z)$ are at most different polynomials in $z^{1/p_{j}}$ for some integer $p_{j}$, $c_{j}$ is a complex number, and $Q_{j}$ are polynomials in $\log z$ whose coefficients have asymptotic forms
$\sum_{k=0}^{\infty}\alpha_{j,k}z^{-\frac{k}{p}}.$
The asymptotic solutions (5.1) are mainly controlled by one exponential polynomial term $\exp{P_{j}(z)}$ in any sector for some $j$. Therefore, we discover that the essence of the space of solutions of equation (3.2) does not have a big difference with its coefficients, the class of exponential polynomials. The following further result is deduced:
Theorem 5.1
Let the coefficients $a_{0},\ldots,a_{n-1}$ of the equation
$f^{(n)}+a_{n-1}f^{(n-1)}+\cdots+a_{1}f^{\prime}+a_{0}f=0$ (5.2)
be functions in the class of solutions of (3.8). All finite order trancendental solutions of (5.2) are of c.r.g..
Proof. The main idea to prove this theorem is similar to Theorem (3.9). Firstly, we know the influential factor of $p$ is the number of sums of critical rays by Lemma 2.3, so the quantity of $p$ for all solutions in angular domains is limited. Owing to it being an integer, we can find the lowest common multiple and also denote it as $p$. Then, we transform (5.2) with $t=z^{1/p}$ to an equation in $t$
${g}^{(n)}+b_{n-1}(t)g^{(n-1)}+\cdots+b_{1}(t)g^{\prime}+b_{0}(t)g=0,$ (5.3)
with $f(z)=g(t),a_{j}(z)=b_{j}(t),j=1,\ldots,n-1.$ So
$b_{j}(t)=\sum_{j=0}^{n}e^{P_{j}(t)}t^{pc_{j}}Q_{j}(t,p\log t),$
where $Q_{j}$ is a polynomial in $\log z$ over the field of formal series $\sum_{k=0}^{\infty}\alpha_{j,k}z^{-{k}}.$ Given a $\theta$, there exists a sector $S:|\arg z-\theta|<h$
$\begin{split}b_{j}(t)&=\sum_{j=0}^{n}e^{P_{j}(t)}t^{pc_{j}}\sum_{m=0}^{d}p^{m}\left(\sum_{k=0}^{\infty}\alpha_{j,k,m}t^{-k}\right)\log^{m}t\\ &=\sum_{j=0}^{n}e^{P_{j}(t)}t^{pc_{j}}\sum_{m=0}^{d}\left(\hat{Q}_{j,m}(t)\log^{m}t+o\left(\frac{{|\log t|}^{m}}{{|t|}^{N}}\right)\right),\end{split}$
where $|t|^{N}\hat{Q}_{j,m}$ are polynomials of degree $N$, for some $N\in\mathbb{N}$. Accordingly,
$b_{j}(t)=\sum_{j=0}^{n}e^{P_{j}(t)+O(\log|t|)}R_{j}(t)(1+o(1)),\quad t\rightarrow\infty.$ (5.4)
Here, $R_{j}(t)$ are rational functions with only pole 0.
Further, if we take $b_{j}$ as the coefficients in equation (3.8), we can also construct coefficient differential equations, by the method of Theorem 3.9, successively construct the fundamental coefficient differential equations in sectors. Only the last recursive process will be taken into consideration. Namely,
$H_{j_{1},j_{2},\cdots,j_{s-1}}(t_{\mu})=\sum_{j_{s}=1}^{k_{j_{1},\ldots,j_{s-1}}}e^{q_{j_{1},j_{2},\cdots,j_{s}}t+O(\log t)}\cdot\sum_{m=0}^{n}\hat{R}_{j_{1},j_{2},\cdots,j_{s},m}(t)(1+o(1))\frac{g^{(m)}}{g}=o\left(e^{ar_{\mu}}\right),$
where $a\in\mathbb{R}$, $q_{j_{1},j_{2},\cdots,j_{s}}$ are some complex numbers, $\hat{R}_{j_{1},j_{2},\cdots,j_{s},m}$ are rational functions with only pole 0, and $r_{\mu}=|t_{\mu}|$. Without generality, we simplify this equation as
$\sum_{l=1}^{K}e^{c_{l}t+O(\log|t|)}\cdot\sum_{m=0}^{n}\hat{R}_{l,j}(t)(1+o(1))\frac{g^{(j)}}{g}=o\left(e^{ar}\right),\quad r=|t|,\kern 5.0ptK\in\mathbb{N},\kern 5.0pt\in\mathbb{C}.$ (5.5)
Set
$\tilde{\kappa}_{l,m}(t):=\left|\frac{e^{\lambda_{l}t+O(\log|t|)}\left(\sum\limits_{j=0}^{n}\hat{R}_{l,j}\dfrac{g^{(j)}}{g}\right)}{e^{\omega_{m}t+O(\log|t|)}\left(\sum\limits_{j=0}^{n}\hat{R}_{m,j}\dfrac{g^{(j)}}{g}\right)}\right|=e^{\Re{(\lambda_{l}-\omega_{m})}t+O(\log r)}\left|\frac{\sum\limits_{j=0}^{n}\hat{R}_{l,j}\dfrac{g^{(j)}}{g}}{\sum\limits_{j=0}^{n}\hat{R}_{m,j}\dfrac{g^{(j)}}{g}}\right|.$
From lemma (3.1), there exists $\varepsilon=\arcsin\left(\left(2C\log r\right)/\left(r|\lambda_{l}-\omega_{m}|\right)\right)$ satisfies
$\Re((\lambda_{l}-\omega_{m})t)+C\log r\leq-|t|\cdot|\lambda_{l}-\omega_{m}|\cdot|\sin\varepsilon|+C\log r<0,$
for $t\in S_{m}(\varepsilon)$, every $C>0$, $r\geq r_{0}$. Thus, the remanent proof is the same with the sections in 3.2 and 3.9. The part relevant to denote the angular domains is more complicated, which is omitted here. Besides some areas near the critical rays, and parabolic strips, it is always established the fundamental coefficient differential equations are in the form of
$\gamma_{n}(t)g^{(n)}+\gamma_{n-1}(t)g^{(n-1)}+\cdots+\gamma_{0}(t)g=0,$ (5.6)
where $\gamma_{j}(t)$ has an asymptotic expansion
$\gamma_{j}(t)=t^{n_{j}}\sum_{k=0}^{\infty}\frac{1}{t^{k}},t\rightarrow\infty.$
By the similar proof of Theorem 3.2 with Lemma 2.4 (when Lemma 2.4 lose effect for infinite many zeros of $f$ in each sector, we take the same discussion with (4.10) into account, which is abbreviated in this part), it follows
$\log|g|=\Re G+O(\log|t|),$ (5.7)
for every narrow sector, except for some areas constructed by the critical rays, stokes rays and discussed above. Here $G$ is a polynomial in $t^{1/p_{0}},p_{0}\in\mathbb{N^{*}}.$
Considering the indicator function of $f$ by (5.7), $t=z^{p}$, $f(z)=g(t)$ and Lemma 2.1, $f$ is of finite rational order and completely regular growth. $\Box$
We set the Class of finite order transcendental solutions of (3.8) in Theorem 3.9, and (5.2) in Theorem 5.1 as $\mathcal{EP}^{0}$, $\mathcal{EP}^{1}$ individually. On this point of view, if we take the solutions of equation (5.2) as coefficients of a linear differential again, then we will get the same conclusion the finite order transcendental solutions are of c.r.g. in a class defined as $\mathcal{EP}^{2}$ and the orders are rational. As if we do this progress recursively, even if there may exist new finite order transcendental functions, but every asymptotic form in a narrow sector with an open angle $\varepsilon$ is just the same form as
$\log|f|=\Re P(z)+O(\log|z|)$ (5.8)
with $P(z)$ a polynomial in $z^{1/p}$ for some integer $p$. We name the class of functions, which are in $\bigcup_{k=0}^{\infty}\mathcal{EP}^{k}$, by repeatedly putting the class of solutions of former equations to the coefficients of later equations, as of exponential polynomial type. Thus, it is not hard to get this result below
Corollary 5.2
Set the coefficients $a_{0},\ldots,a_{n-1}$ of the equation
$f^{(n)}+a_{n-1}f^{(n-1)}+\cdots+a_{1}f^{\prime}+a_{0}f=0$ (5.9)
be functions of exponential polynomial type. All finite order transcendental solutions of (5.9) are of c.r.g.. In other words, functions of exponential polynomial type are of c.r.g..
This corollary reveals a fundamental closure property: the class of functions of exponential polynomial type, defined inductively as $\bigcup_{k=0}^{\infty}\mathcal{EP}^{k}$, constitutes a vast and coherent hierarchy in which the property of completely regular growth is hereditary. Starting from the classical exponential polynomials ($\mathcal{EP}^{0}$), each step $\mathcal{EP}^{k}\to\mathcal{EP}^{k+1}$ corresponds to taking all finite-order transcendental solutions of linear differential equations whose coefficients lie in $\mathcal{EP}^{k}$. Our main theorems guarantee that this process never leaves the world of completely regular growth.
While the inductive definition suggests a potentially expanding hierarchy, the strict inclusion $\mathcal{EP}^{k}\subsetneq\mathcal{EP}^{k+1}$ is not established rigorously here. Perturbative constructions, such as those alluded to in later sections, might generate new functions from $\mathcal{EP}^{0}$ to $\mathcal{EP}^{1}$ for an example, hinting that the classes are indeed proper extensions. However, a full verification would require explicit examples and lies beyond the immediate scope. This hierarchical perspective naturally raises deeper questions about the scope and limits of the class of exponential polynomial type, especially in relation to the broader theory of completely regular growth.
5.2 An Operator-Theoretic Perspective and Parameter Space Analysis
The classical question of existence and regularity of finite-order solutions to linear differential equations can be reframed fruitfully within the framework of operator theory on function spaces. This perspective not only clarifies the underlying mechanisms of growth inheritance but also opens pathways to construct new classes of solutions and to analyze the structure of the parameter space in perturbed equations.
The asymptotic theory of linear differential equations naturally leads to integral formulations. Consider a homogeneous linear differential equation
$f^{(n)}+P_{1}f^{(n-1)}+\cdots+P_{n-1}f^{\prime}+P_{n}f=0.$
By introducing suitable approximating functions $Q_{i}$ and defining the non-homogeneous term
$F(f)=(Q_{1}-P_{1})f^{(n-1)}+\cdots+(Q_{n}-P_{n})f,$
the equation can be rewritten as
$Lf:=f^{(n)}+Q_{1}f^{(n-1)}+\cdots+Q_{n}f=F(f).$
Using the method of variation of constants, the general solution takes the form
$f(z)=\sum_{i=1}^{n}c_{i}F_{i}(z)+\sum_{i=1}^{n}F_{i}(z)\int_{g_{i}}^{z}\frac{\Delta_{i}(t)}{\Delta(t)}F(f(t))dt,$
where $\{F_{i}\}$ are linearly independent solutions of $Lf=0$, $\Delta$ is the Wronskian, and $\Delta_{i}$ are corresponding cofactors [36, p. 128]. This representation is a Volterra integral equation of the second kind:
$f(z)=f_{\text{hom}}(z)+\int_{0}^{z}K(z,t)f(t)dt,$ (5.10)
where $f_{\text{hom}}$ is a particular solution of the homogeneous equation $Lf=0$ and $K$ is a kernel derived from the coefficients. The associated affine integral operator
$(Tf)(z)=\tilde{f}(z)+\int_{0}^{z}K(z,t)f(t)dt$ (5.11)
becomes the central object of study.
To control growth, we work in weighted Banach spaces of entire functions. For a continuous, non-increasing weight $v:[0,\infty)\to[0,\infty)$ with $\lim_{r\to\infty}r^{m}v(r)=0$ for all $m\in\mathbb{N}$, define
$\mathcal{H}_{v}^{\infty}(\mathbb{C}):=\{f\in\mathcal{H}(\mathbb{C}):\|f\|_{v}:=\sup_{z\in\mathbb{C}}v(|z|)|f(z)|<\infty\}.$
Of particular importance are the exponential weights $v(r)=e^{-ar^{\sigma}}$ ($a>0$, $\sigma>0$), which give spaces
$\mathcal{B}_{\sigma,a}=\left\{f\in\mathcal{H}(\mathbb{C}):\|f\|_{\sigma,a}=\sup_{z\in\mathbb{C}}|f(z)|e^{-a|z|^{\sigma}}<\infty\right\}.$
These spaces are closely related to generalized weighted Bergman spaces of entire functions; see, e.g., [3, 5, 10, 11, 25, 26]. An entire function $f$ belongs to $\mathcal{B}_{\sigma,a}$ if its growth order $\rho(f)$ satisfies $\rho<\sigma$, or $\rho=\sigma$ with type $\tau<a$.
Within this framework, the inheritance of completely regular growth (c.r.g.) for equations with exponential polynomial coefficients can be understood as a consequence of the compression properties of the associated Volterra operator on these weighted spaces. The structure of the coefficients ensures the integral kernel preserves the geometric constraints needed for the operator to act as a contraction, forcing the iterative scheme to converge to a well-behaved (c.r.g.) solution.
The integral operator formulation offers a potential pathway for constructing solutions and analyzing their growth. Consider a perturbed equation of the form
$L(f)=\varepsilon F(f),$ (5.12)
where $L$ and $F$ are differential polynomials with polynomial coefficients and exponential polynomial coefficients seperately, and $\deg L>\deg F$. Under this framework, the operator $T$ in (5.11) can be formally decomposed as $Tf=f_{\text{hom}}+\varepsilon T^{\prime}f$. A natural line of inquiry is to investigate the conditions under which $T^{\prime}$ acts boundedly on a suitably chosen weighted Bergman space $\mathcal{B}_{\sigma,a}$. If such boundedness can be established, then for sufficiently small $|\varepsilon|$, the operator $\varepsilon T^{\prime}$ might become a contraction, leading to a unique fixed point of $T$ via the Banach fixed-point theorem. This fixed point would then correspond to a solution of the differential equation with controlled growth properties.
This perspective, while promising, immediately raises several non-trivial questions and highlights the delicate interplay between the operator’s properties, the choice of function space, and the perturbation parameter:
  1. 1.
    Feasibility of the boundedness condition. Under what precise conditions on the coefficients of $L$ and $F$ is the associated operator $T^{\prime}$ bounded on a space $\mathcal{B}_{\sigma,a}$? The answer is not automatic and depends crucially on the structure of the integral kernel derived from the differential equation.
  2. 2.
    Geometry of the valid parameter set. If a boundedness condition holds, what is the nature of the set of $\varepsilon$ for which the fixed-point argument succeeds? Is it necessarily a disk, or could it exhibit a more complex structure (e.g., sectors or petals) due to asymptotic dependencies on $\arg\varepsilon$, akin to Stokes phenomena [38, Chap. 7]?
  3. 3.
    Global existence beyond small perturbations. Can the fixed-point approach be extended beyond the regime of small $|\varepsilon|$? For large $|\varepsilon|$, the operator may no longer be contractive, and the existence of finite-order solutions becomes a much more subtle problem, potentially connected to the spectral and dynamical properties of $T^{\prime}$.
These questions underscore that the operator-theoretic reframing is not a ready-made theorem but a conceptual lens that transforms classical questions about growth inheritance and solution structure into concrete problems in functional analysis. The remainder of this work explores one concrete manifestation of this perspective, establishing the inheritance of completely regular growth for equations with exponential polynomial coefficients. The broader questions above remain open and provide a rich direction for future research.
The operator perspective also sheds light on the relationship between growth orders and the dynamical properties of $T$. When $|\varepsilon|$ is small, $T$ is a contraction on a closed subspace of $\mathcal{B}_{\sigma,a}$ determined by the initial conditions, guaranteeing a unique finite-order fixed point. As $|\varepsilon|$ increases, the contraction property may break down. The operator can, in principle, exhibit more complex dynamics such as hypercyclicity or chaos (i.e., the existence of a dense orbit and a dense set of periodic points) on the function space [2, 3, 5].
In summary, recasting the HITW problem within the operator-theoretic framework does more than provide an alternative proof; it unveils a deep connection between the asymptotic theory of differential equations, the dynamics of linear operators on function spaces, and the geometry of parameter spaces. This viewpoint not only clarifies the inheritance of completely regular growth but also opens new avenues for constructing exotic solutions and understanding the global landscape of perturbed linear differential equations.
6 Appendix: A modified Wasow Theorem on sectors with exceptional sets
We use the definition of the asymptotic symbol $\sim$” in a point-set $T$ near 0 given by Wasow, see [38, p. 32].
Definition 6.1
Let the function $g(x)$ be defined in a point-set $T$ of the complex $x$-plane having $x=0$ as an accumulation point. The power series $\sum_{j=0}^{\infty}a_{j}x^{j}$ is said to represent $g(x)$ asymptotically, as $x\to 0$ in $T$, if
$x^{-m}\left[g(x)-\sum_{j=0}^{m}a_{j}x^{j}\right]$
tend to zero, for all $m\geq 0$, as $x$ tends to zero in $T$.
Our goal is to weaken the conditions for Wasow’s asymptotic conclusions so that the differentiating-operation can be performed on a point-set $T$, rather than necessarily on a complete sector. However, it is not always true, if we give no limitation on this set. For example, if $T$ is only a real axis, then the function $f(x)=e^{-x}\sin(e^{-1/x})$ cannot be termwised differentiated. See [38, p. 38].
Without generality, we write sectors $T$ as
$T:\{x~|~\theta_{1}\leq\arg x\leq\theta_{2},\theta_{2}>\theta_{1},0<|x|<x_{0},x_{0}>0\}$
and $\tilde{T}$ as
$\tilde{T}:\{x~|~\theta_{1}<\tilde{\theta}_{1}\leq\arg x\leq\tilde{\theta}_{2}<\theta_{2},\tilde{\theta}_{2}>\tilde{\theta}_{1},0<|x|<x_{0},x_{0}>0\},$
so $\tilde{T}\subset T$. We consider a subset $T^{*}$ of $T$ with two properties:
  1. 1.
    $T^{*}\cup\{0\}$ is path connected;
  2. 2.
    $\TT\backslash T^{*}$ only has countably many closed connected components $\{D_{j}\}$.
  3. 3.
    (Uniform interior.) There exists a $T_{0}$ such that $T_{0}\subset T^{*}$, and $T\backslash T_{0}$ only has countably many closed connected components $\{E_{j}\}$ with $D_{j}\subset E_{j}$ and $\mathrm{dist}(z,\partial D_{j})\geq\alpha|z|$ for a constant $\alpha>0$ and uniformly for all sufficiently small $z\in T_{0}$.
$T^{*}$ is defined by the reason below.
Suppose $g(x)$ is holomorphic in $T^{*}$. We know that the asymptotic form of $f(x)$ is term-by-term integrable when the integration path is in $T$, see [38, Theorem 8.7].
The set $T^{*}$ generalizes a sector by allowing the removal of countably many closed connected components (the “holes” $D_{j}$). However, to retain the ability to differentiate asymptotic expansions termwise, we must ensure that near each point of $T^{*}$ there exists a disk of radius proportional to the distance to the origin that stays entirely within $T^{*}$. This is guaranteed by the third property, which posits the existence of a subset $T_{0}\subset T^{*}$ such that every point of $T_{0}$ maintains a distance at least $\alpha|x|$ from the boundary of any hole. This uniform interior condition allows us to apply Cauchy estimates exactly as in Wasow’s original proof. Consequently, we obtain the following lemma, which extends [38, Theorem 8.8] to sectors with holes.
Lemma 6.2
If $f(x)$ is holomorphic in a set $T^{*}$ with properties defined above, and if
$f(x)\sim\sum_{j=0}^{\infty}a_{j}x^{j},~~x\in T^{*},$
then
$f^{\prime}(x)\sim\sum_{j=1}^{\infty}ja_{j}x^{j-1}$
in $\tilde{T}\cap T_{0}$, for every proper subsector $\tilde{T}\subset T$.
Similarly, when we consider the asymptotic form of $f(z)$ near $\infty$, by taking $z=1/x$ and $f(z)=g(x)$, we define sectors
$S:\{z~|~\theta_{1}\leq\arg z\leq\theta_{2},\theta_{2}>\theta_{1},|z|>r_{0},r_{0}>0\},$
and for a proper subsector $\tilde{S}$ of $S$ (with $\theta_{1}<\tilde{\theta}_{1}\leq\arg z\leq\tilde{\theta}_{2}<\theta_{2}$ and $|z|>r_{0}$), we have $\tilde{S}\subset S$.
We consider a subset $S^{*}$ of S with the following properties:
  1. 1.
    $S^{*}\cup\{\infty\}$ is path connected;
  2. 2.
    $S\setminus S^{*}$ consists of countably many closed connected components $\{D_{j}^{\prime}\}$ (the “holes”);
  3. 3.
    (Uniform interior.) There exists a subset $S_{0}\subset S^{*}$ such that $S\setminus S_{0}$ consists of countably many closed connected components $\{E_{j}^{\prime}\}$ with $D_{j}^{\prime}\subset E_{j}^{\prime}$ and
    $\mathrm{dist}(z,\partial D_{j}^{\prime})\geq\frac{\alpha}{|z|},\quad\text{for a constant }\alpha>0,$
    uniformly for all sufficiently large $z\in S_{0}$.
Roughly speaking, $S^{*}$ is $S$ with “holes”, and $S_{0}$ is a subregion of $S^{*}$ that maintains a uniform distance from the holes, scaled appropriately near infinity.
Corollary 6.3
If $f(z)$ is holomorphic in a set $S^{*}$ with properties defined above, and if
$f(z)\sim\sum_{j=0}^{\infty}a_{j}z^{-j},\quad z\in S^{*},$
then
$f^{\prime}(z)\sim\sum_{j=1}^{\infty}(-j)a_{j}z^{-j-1}$
in $\tilde{S}\cap S_{0}$, for every proper subsector $\tilde{S}\subset S$.
Definition 6.4
Let $G$ be a subset of the Riemann sphere $\mathbb{C}\cup\{\infty\}$. The topological hull of $G$, denoted by $T(G)$, is defined as the union of $G$ and all bounded components of its complement. In other words, $T(G)$ is the smallest simply connected set (in $\mathbb{C}\cup\{\infty\}$) that contains $G$.
A simple conclusion is that $T(S^{*})=S$.
Considering the linear system
$z^{-d}\frac{d\bm{Y}}{dz}=\bm{A}(z)\bm{Y},$ (6.1)
$\bm{A}(z)\sim\sum_{j=0}^{\infty}\bm{A}_{j}z^{-j},\kern 5.0ptz\in S^{*}.$ (6.2)
We will give a modified result from Wasow’s theorem, and illustrate that the results on $S$ and $S^{*}$ are kept consistent.
Theorem 6.5
Suppose $^{*}, S~\tilde{S}$ and $S_{0}$ be defined above. Let $\bm{A}(z)$ be an $n$-by-$n$ matrix function, which is holomorphic in $S^{*}$ and has the asymptotic form (2.7) in $S^{*}$. Then, for any sufficiently narrow subsector $\tilde{S}\subset S$, the differential equation (2.5) possesses a fundamental matrix solution of the form
$\bm{Y}(z)=\hat{\bm{Y}}(z)z^{\bm{G}}e^{\bm{Q}(z)}$ (6.3)
in the region $\tilde{S}^{*}:=\tilde{S}\cap S_{0}$, where $\hat{\bm{Y}}(z)$ admits an asymptotic series in powers of $z^{-1/p}$ (for some positive integer $p$) as $z\to\infty$ in $\tilde{S}^{*}$, $\bm{G}$ is a constant matrix, and $\bm{Q}(z)$ is a diagonal matrix whose diagonal elements are polynomials in $z^{1/p}$.
Proof. By the conditions of $S^{*}$ we set and Cor. 6.3, we have shown that the asymptotic form of the coefficient matrix $\bm{A}(z)$ of equation (6.2) is term-by-term integrable and term-by-term differentiable. We ensure basic operations on the elements in a matrix in the asymptotic sense. (Other operations can be found in [38, Section 8.1].)
We know that Wasow’s Theorem 2.3 on the description of the sector $S$ exists only for the Main Asymptotic Existence Theorem [38, Theorem 12.1] and for dealing with the effect of the Stokes phenomenon on asymptotic solutions [38, Theorem 15.3]. Here we will not repeat the other steps of the proof of Theorem 2.3, since the only ideas of these steps are classifications of the eigenvalues of the matrix $\bm{A}_{0}$ in (6.2). We first come to illustrate the improvements to the Main Asymptotic Existence Theorem.
We introduce this theorem here below,
Lemma 6.6
Let $S$ be an open sector of the complex $z$-plane with vertex at the origin and a positive central angle not exceeding $\pi/(d+1)$($d$ a nonnegative integer). Let $f(z,\bm{y})$ be an $N$-dimensional vector of $z$ and the $N$-dimensional vector $\bm{y}$ with the following properties.
  1. 1.
    $f(z,\bm{y})$ is a polynomial in the components $y_{j}$ of $\bm{y}$, $j=1,\ldots,N$, with coefficients that are holomorphic in $z$ in the region
    $0<r_{0}\leq|z|<\infty,~~x\in S,~(r_{0}~\text{a constant}).$
  2. 2.
    The coefficients of the polynomial $f(z,\bm{y})$ have asymptotic series in power of $z^{-1}$, as $z\to\infty$ in $S$.
  3. 3.
    If $f_{j}(z,\bm{y})$ denotes the components of $f(z,\bm{y})$ then all eigenvalues $\lambda_{j}$, $j=1,2,\ldots,N$ of the Jacobian matrix
    $\left\{\lim\limits_{z\to\infty\atop z\in S}\left(\left.\frac{\partial f_{j}}{\partial y_{k}}\right|_{\bm{y}=\bm{0}}\right)\right\}$
    are different from zero.
  4. 4.
    The differential equation
    $z^{-d}\bm{y}^{\prime}=f(x,\bm{y})$ (6.4)
    is formally satisfied by a power series of the form
    $\sum_{j=1}^{\infty}\bm{\alpha}_{j}z^{-j}.$
Then there exists, for sufficiently large $z$ in $S$, a solution $\bm{y}=\bm{\phi}(z)$ of (6.4) such that, in every proper subsector $\tilde{S}$ of $S$,
$\bm{\phi}(z)\sim\sum_{j=1}^{\infty}\bm{\alpha}_{j}z^{-j},~~z\to\infty.$
We claim that $S$ in the Lemma 6.6 can be replaced by $S^{*}$, as well as the proper subsector $\tilde{S}$ of $S$ substituted by $\tilde{S}^{*}$ such that the topological hull $T(\tilde{S}^{*})=\tilde{S}$.
We recognize the proof of Lemma 6.6 can be simplified by the assumption that the eigenvalues $\lambda_{j},~j=1,2,\ldots,N$, are distinct. In this part, we will follow the idea of [38, pp. 65-75]. When we set
$\bm{a}(z)=f(z,\bm{0}),~\bm{A}(z)=\left\{\left.\frac{\partial f_{j}}{\partial y_{k}}\right|_{\bm{y}=\bm{0}}\right\},~g(z,\bm{y})=f(z,\bm{y})-a(z)-\bm{A}(z)\bm{y},$ (6.5)
the differential equation (6.4) becomes
$z^{-d}\bm{y}^{\prime}=\bm{a}(z)+\bm{A}(z)\bm{y}+g(x,\bm{y}).$ (6.6)
In (6.2), without loss of generality we may assume that
$\bm{A}_{0}=\mathrm{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{N}),$
assured by a preliminary constant linear transformation of $\bm{y}$ with nonvanishing determinant.
By [38, Theorem 9.3], there exists a vector function $\bm{\phi}(z)$ holomorphic in $S$ for $|z|>r_{0}$ such that
$\bm{\phi}(z)\sim\sum_{j=1}^{\infty}\bm{\alpha}_{j}z^{-j},~z\to\infty,~z\in S.$ (6.7)
It is obvious that (6.7) holds on $S^{*}$, and we have guaranteed the termwise differentiability in (6.7) on $S_{0}\subset S^{*}$, the uniform interior-point set with boundary distance $\alpha/|z|$ for all sufficiently large $z\in S_{0}$. By the transformation
$\bm{u}=\bm{y}-\bm{\phi}(z),$
(6.6) is changed to
$z^{-d}\bm{u}^{\prime}=\bm{b}(z)+\bm{B}(z)\bm{u}+h(z,\bm{u}),$ (6.8)
where $\bm{b}(z)\sim 0$ for $z\to\infty,z\in S_{0}$,
$\lim\limits_{z\to\infty}\bm{B}(z)=\bm{A}_{0},~z\in S_{0},$
and $h(z,\bm{u})$ is a polynomial in the components $u_{j}$ of $\bm{u}$ with coefficients that admit asymptotic power series in $z^{-1}$, as $z\to\infty$ in $S_{0}$, which also has no constant or linear terms in the $u_{j}$, $j=1,\ldots,N$. Therefore, the main work now is to show the differential equation (6.8) admits a solution that is asymptotic to zero, as $z\to\infty$ in $S_{0}$.
Further, equation (6.8) can be modified as
$z^{-d}\bm{u}^{\prime}=\bm{A}_{0}\bm{u}+p(z,\bm{u}),$ (6.9)
where
$\bm{p}(z,\bm{u})=\bm{b}(z)+(\bm{B}(z)-\bm{A}_{0})\bm{u}+\bm{h}(z,\bm{u}).$
Roughly speaking, here the idea is to prove that the function $p(z,\bm{u})$ is much smaller than $\bm{u}$, for large $z$ and small $\bm{u}$, while $\bm{b}(z)$ is asymptotic to zero, $\bm{B}(z)-\bm{A}_{0}$ tends to zero, and $h(z,\bm{u})$ contains no linear terms in $\bm{u}$. If $p(z,\bm{u})$ equals to zero identically, then equation (6.9) has the solution $\bm{u}\equiv\bm{0}$.
We will replace (6.9) by an equivalent integral equation
$\bm{u}(z)=\bm{V}(z)\bm{V}^{-1}(z_{0})\bm{k}+\int_{z_{0}}^{z}\bm{V}(z)\bm{V}^{-1}(t)t^{d}\bm{p}(t,\bm{u}(t))dt.$ (6.10)
Here $\bm{k}$ is a certain constant vector, $z_{0}$ is some fixed point, and $\bm{V}(z)$ is some fundamental matrix solution of the differential equation
$\bm{z}^{-d}\bm{V}^{\prime}=\bm{A}_{0}\bm{V}.$ (6.11)
Substituting a particular solution
$\bm{V}(z)=\exp\left[\frac{z^{d+1}}{d+1}\bm{A}_{0}\right]$
into (6.10) with slightly modifying the path of integration, we have
$\bm{u}(z)=\int_{\bm{\Gamma}^{*}(z)}\exp\left[\frac{z^{d+1}-t^{d+1}}{d+1}\bm{A}_{0}\right]t^{d}\bm{p}(t,\bm{u}(t))dt,$ (6.12)
where $\bm{\Gamma}^{*}(z)$ is a set of $N$ individual paths $\gamma_{j}(z)$ corresponding to the components $u_{k}$ of $\bm{u}$. It is known that $\bm{u}(z)$ in (6.12) is independent of the choice of $z_{0}$ in (6.10). The abbreviation form of (6.12) can be written as a functional equation
$\bm{u}=\mathscr{P}\bm{u},$ (6.13)
which the solution is asymptotic to zero by the classic method of successive approximations. Let $\bm{u}_{j}(z)$ be a sequence of vector functions for $j=0,1,\ldots$. The convergence of the sequence will be established by estimating the difference
$\bm{u}_{j+1}-\bm{u}_{j}=\mathscr{P}\bm{u}_{j}-\mathscr{P}\bm{u}_{j-1},$ (6.14)
where $\bm{u}_{0}\equiv 0$ and $\bm{u}_{j+1}=\mathscr{P}\bm{u}_{j}.$ Then we should find a suitable set of path $\bm{\Gamma}^{*}(z)$.
The specific construction of the integration path can be found in [38, Sec. 14.3]. Here, we provide a brief overview of Wasow’s method, slightly modify some of the paths $\gamma_{j}(z)$, and explain that such modifications are necessary and do not alter the estimates of the continuous approximation method to get the solution of (6.13) asymptotic to zero in $S_{0}$.
We should make the exponential term in (6.12) be bounded by choosing the paths $\gamma_{j}(z)$. By changing the variables,
$\tau=t^{d+1}~\text{and}~\xi=z^{d+1},$ (6.15)
the image of $S_{0}$ and the sector $T(S_{0})$ of the $t$-plane($z$-plane) individually become $\Sigma^{*}$ and a sector $T(\Sigma^{*})$ in the $\tau$-plane($\xi$-plane) with central angle not exceeding $\pi$. We know there exist $N$ lines in the $\tau$-plane, consequently $2N$ rays denoted by $l_{1},\ldots,l_{2N}$ sorted anticlockwise, such that
$\Re(\tau\lambda_{j})=0,$
for $j=1,\ldots,N$. We only consider the case inside the sector
$\Sigma_{j}=\{z~|\arg l_{j-1}<\arg z<\arg l_{j+1}\},$
$j=1,\ldots,2N$, and the analysis in every $\Sigma_{j}$ are same. For the sake of simplicity, We set $\Sigma_{j}$ as $\Sigma_{1}$, so there exists only one $l_{1}\subset\Sigma_{1}$ such that $\Re(\tau\lambda_{k})=0$ for $k=1,\ldots,N$. Let $\Sigma_{1}^{*}=\Sigma^{*}\cap\Sigma_{1}$ and the pre-image of $\Sigma_{1}^{*}$ and $\Sigma_{1}$ be $S_{1}^{*}$ and $S_{1}$ individually. Without generality, we can also treat $T(\Sigma_{1}^{*})$ as a sector and maintain the property $T(\Sigma_{1}^{*})=\Sigma_{1}$, which means that the “holes” in $\Sigma$ have no intersection with the sector $\Sigma_{1}$. Therefore, $T(S_{1}^{*})=S_{1}$.
Next, we arrange all the $\lambda_{j}$ such that $\Re(\tau\lambda_{j})<0$ in $\Sigma_{1}^{*}$ for $j=1,\ldots,j_{1}-1$, $\Re(\tau\lambda_{j_{1}})=0$ in $\Sigma_{1}^{*}$, and $\Re(\tau\lambda_{j})>0$ in $\Sigma_{1}^{*}$ for $j=j_{1}+1,\ldots,N$. Let $\xi_{1}$ be some point on $\Sigma_{1}^{*}$ and $r_{0}$ be a sufficient large constant, such that $|\xi_{1}|>r_{0}^{q+1}$, and denote by $\Sigma_{1}(\xi_{1})$ the closed sector with vertex at $\xi_{1}$ and boundary rays parallel to those of $\Sigma_{1}$. Further, we set
$\Sigma_{1}^{*}(\xi_{1})=\Sigma_{1}(\xi_{1})\cap\Sigma_{1}^{*}.$
Due to the Hausdorff space’s property on $\mathbb{C}$, we can choose $\xi_{1}$ sufficiently close to $\xi$ such that the vector $\vec{\xi_{1}\xi}$ satisfies
$\vec{\xi_{1}\xi}~\bigcap~\Sigma_{1}(\xi_{1})\backslash\Sigma_{1}^{*}(\xi_{1})=\emptyset.$
Thus, for $j=1,\ldots,j_{1}-1$, we choose the common path by the preimage curve of $\vec{\xi_{1}\xi}$ in $S_{1}^{*}$ as integration path $\gamma_{j}(z)$ of $\bm{\Gamma}^{*}(z)$ in (6.12). We notice that the exponential term in (6.12) decreases along $\gamma_{j}(z)$.
For $j\geq j_{1}$, we choose some ray $l_{k}$, along with $\Re(\tau\lambda_{k})>0$ from the origin into $\Sigma_{1}^{*}$. Let $l^{*}_{k}$ be the ray from $\infty$ to the point $\zeta$ parallelling with $l_{k}$. As same as the circumstance above, we can choose $\xi_{2}$ on $l_{k}^{*}$ sufficiently close to $\xi$ such that the vector $\vec{\xi_{2}\xi}$ satisfies
$\vec{\xi_{2}\xi}~\bigcap~\Sigma_{1}(\xi_{1})\backslash\Sigma_{1}^{*}(\xi_{1})=\emptyset.$
Figure 4: $\Sigma_{1}^{*}$ and $\Sigma_{1}^{*}(\xi_{1})$ in $\tau$-plane
Thus, for $j=j_{1},\ldots,N$, we choose the common path by the preimage curve of $\vec{\xi_{2}\xi}$ in $S_{1}^{*}$ as integration path $\gamma_{j}(z)$ of $\bm{\Gamma}^{*}(z)$ in (6.12). We find that the exponential term in (6.12) also decreases along $\gamma_{j}(z)$.
The pre-image $S_{1}(z_{1})$ of $\Sigma_{1}(\xi_{1})$ is a region in the $z$-plane bounded by two curves meeting at $z_{1}=\xi_{1}^{1/(d+1)}$ and asymptotical to the boundary rays of $S_{1}$. Therefore, $S_{1}(z_{1})$ contains all points at sufficiently large distance from the origin of any closed subsector of $S_{1}$. Contrast [38, Figure 14.1], we given a visualized construction in Figure 4 and Figure 5.
Figure 5: $S_{1}^{*}$ and $S_{1}^{*}(z_{1})$ in $t$-plane
We constructed the same integration path $\gamma_{j}$ for $j<j_{1}$ as Wasow gave. For the other integral paths, $\gamma_{j}$ for $j_{1}\leq j\leq N$, we only changed the path from $\infty$ to a fixed point to be part of this path. This local integration path can be found at any inner point of $S_{1}^{*}(z_{1})$, consequently at any inner point of $S_{1}^{*}$ when $|z|\to\infty$, further at any inner point of $S_{0}$ when the opening of topological hull $T(S_{0})$ is not larger than $\pi/(d+1)$ and $|z|\to\infty$, by the arbitrary choice of $\Sigma_{1}^{*}$ in $\Sigma^{*}$. Therefore, it is established that
$\Re\left[(z^{d+1}-t^{d+1})\lambda_{j}/(d+1)\right]\leq-\left|z^{d+1}-t^{d+1}\right|\lambda_{0}\mu/(d+1),$ (6.16)
where $\lambda_{0}=\min\limits_{j}|\lambda_{j}|$ and $\mu$ is a positive constant independent of $\lambda_{0},j$ and $\xi_{1},\xi_{2}$.
We use (6.16) on the successive approximation to the estimation of (6.14). We won’t elaborate on this part of the proof because its steps are basically same as those in [38, Sec. 14.4-14.5]. It is deduced that
[…]
is dominated by a convergent geometric series in $S_{0}$. Namely,
$\lim\limits_{k=0}\bm{u}_{k}=\lim\limits_{k\to\infty}\sum_{s=0}^{k-1}(\bm{u}_{s+1}-\bm{u}_{s})$
exists and is holomorphic for $z\in S_{0}$. Note that the only difference is the estimate on the component of
$\bm{u}_{k+1}-\bm{u}_{k}||,~~k=0,1,2,\cdots,$
where the component path $\gamma_{j}$ for $j\geq j_{1}$ is a curve from $z_{2}=\xi_{2}^{1/(d+1)}$ to $z$, which is a part of the curve from $\infty$ to $z$. In fact, combining (6.12) we have given a more accurate estimate under every relation
[…]
Here $\bm{\Gamma}^{*}(z)$ is the set of all curves $\gamma_{j}(z)$ from $z_{1}$ to $z$, for $j<j_{1}$, and curves $\gamma_{j}(z)$ from $z_{2}$ to $z$, for $j_{1}\leq j\leq N$; $\bm{\Gamma}(z)$ is the set of all curves $\gamma_{j}(z)$ from $z_{1}$ to $z$, for $j<j_{1}$, and curves from $\infty$ to $z$ containing $\gamma_{j}(z)$, for $j_{1}\leq j\leq N$.
Now the functional equation (6.13) has a solution $\bm{u}(z)\sim 0$ for $z\in S_{0}$ when the opening of sector $T(S_{0})$ is not larger than $\pi/(d+1)$, see [38, Eq. 14.40]. Then by
$\bm{y}=\bm{u}+\bm{\phi}(z),~z\in S_{0},~\text{and}~~\bm{y}=\sum_{j=1}^{\infty}\bm{\alpha}_{j}z^{-j},~z\in S_{0},$
we obtain that
$\bm{\phi}(z)\sim\sum_{j=1}^{\infty}\bm{\alpha}_{j}z^{-j},~z\to\infty,~z\in S_{0}.$
Combing [38, Theorem 12.1] and the classification of eigenvalues of $\bm{A}_{0}$, see [38, pp. 88-111], the differential equation (2.5) possesses a fundamental matrix solution of the form (6.3) corresponding to every subset $\tilde{S}^{*}$ of $S^{*}$, with the properties of $\tilde{S}^{*}$:
  1. 1.
[…]
  2. 2.
    $\tilde{S}^{*}\cup\{\infty\}$ is path connected;
  3. 3.
    $\tilde{S}^{*}$ has uniform interior.
Here $\hat{\bm{Y}}(z)$ permits an asymptotic series in power of $z^{-1/p},p\in\mathbb{N^{*}}$, in $\tilde{S}^{*}$ as $z\rightarrow\infty;$ $\bm{G}$ is a constant matrix, $\bm{Q}(z)$ is a diagonal matrix whose diagonal elements are polynomials in $z^{1/p}$.
The last step is continue the set $S^{*}$, when the opening of sector $T(S^{*})$ is not larger than $\pi/(d+1)$, to a set $S^{*}$ without any limitation of the angular measure of sector $T(S^{*})$.
Due to [38, Theorem 15.3] and the note there below, when some ray rotates in a sector $S$ around 0 anticlockwise (or clockwise) across a ray $l_{0}$ with $\bm{y}(z)$ having an extreme change of growth rate, where appears Stokes phenomenon (see [38, Sec. 15.3]) separating $S$ into $S_{1}$ and $S_{2}$, the fundamental matrix solution $\bm{Y}_{2}(z)=\bm{Y}_{1}(z)\bm{C}$ with $\bm{Y}_{2}(z)$ for $z\in S_{2}$, $\bm{Y}_{1}(z)$ for $z\in S_{1}$, and $\bm{C}$ a constant matrix with entries 1 on the main diagonal.
[…]
Acknowledgements
Thanks to my advisor, Professor Zhi-Tao Wen, for his expert guidance throughout this research, particularly for his insight on the theoretical framework.
I am grateful to Professor Walter Bergweiler for his valuable suggestions during my exchange program, especially regarding functions of completely regular growth and the mindset of holomorphic Dynamics.
A preliminary version of this work was presented as a poster at the conference ‘Classical Function Theory in Modern Mathematics’ (Edinburgh, 07, 2024).
My appreciation extends to the thesis defense committee, especially Yuqiu Zhao (SYSU) and Lun Zhang (FDU), for their rigorous evaluation and constructive feedback on the asymptotic analysis methodology on related topics.
I sincerely thank the anonymous referees for their thorough and insightful reviews, which have greatly improved the quality of this work.
Funding
The author acknowledges the financial support from the Shantou University PhD Student Overseas Exchange Program (2023).
Note on references
We annotate literature across different fields:
[CADE] Complex analysis and differential equations on the complex plane
[FAD] Functional analysis and dynamics
[HA] Harmonic analysis
[ISAA] Integrable systems and asymptotic analysis